### v2

# ICONS: Influence Consensus for Vision-Language Data Selection

Abstract
Visual Instruction Tuning typically requires a large amount of vision-language training data. This data often containing redundant information that increases computational costs without proportional performance gains. In this work, we introduce ICONS, a gradient-driven Influence CONsensus approach for vision-language data Selection that selects a compact training dataset for efficient multi-task training. The key element of our approach is cross-task influence consensus, which uses majority voting across task-specific influence matrices to identify samples that are consistently valuable across multiple tasks, allowing us to effectively prioritize data that optimizes for overall performance. Experiments show that models trained on our selected data (20% of LLaVA-665K) achieve 98.6% of the relative performance obtained using the full dataset. Additionally, we release this subset, LLaVA-ICONS-133K, a compact yet highly informative subset of LLaVA-665K visual instruction tuning data, preserving high impact training data for efficient vision-language model development.
Figure 1: Influence Consensus for Vision-Language Data Selection. (Left) Given a large scale visual instruction tuning dataset (LLaVA-665K [23]), our method computes an influence score of each training example on each target task, aiming to select examples with high influence across multiple tasks. Our selected compact 20% subset (LLaVA-ICONS-133K) contains data points that achieve influence consensus across tasks, as indicated by the green check marks showing high influence ranking on corresponding tasks. (Right) The radar plot shows the relative performance on various benchmarks comparing training with LLaVA-665K (the full dataset) and LLaVA-ICONS-133K (our selected subset), where the selected subset achieves comparable performance to the full dataset.
1 Introduction
Visual instruction tuning has become a critical stage in developing Multimodal LLMs [23, 22], enabling these models to effectively interpret and respond to language instructions based on visual content. Current approaches typically rely on large instruction tuning datasets (e.g., LLaVA 1.5 [22] and Cambrian [39], which use 665K and 7M instruction tuning data points, respectively). These large datasets, while effective, pose practical challenges for research iteration and model development: extended development cycles due to lengthy training times [3, 15], increased storage and memory demands for managing large-scale datasets [36, 8], and substantial computational resources required for training [27, 40]. This motivates a fundamental question:
Can we identify a compact subset of training data that preserves model capabilities while enabling faster experimentation?
Prior work has explored various approaches to data selection, including gradient-based methods [42, 6], influence functions [43, 17], and diversity sampling [44, 4]. These methods typically either focus on optimizing for single tasks in isolation, or maximizing source data diversity without considering target task distributions. However, selecting data for visual instruction tuning requires careful consideration of performance across diverse downstream tasks. Simply selecting the most influential samples for an individual task risks creating a dataset that performs well on that specific benchmark but fails to develop broader capabilities. Rather than focusing on task-specific performance, we aim to identify universally valuable training samples by aggregating gradient-based influence measures across diverse tasks using a voting mechanism, effectively selecting data that contributes to broad model capabilities.
We introduce ICONS (Influence CONsensus vision-language data Selection), a method that builds upon the gradient-based influence estimation approach from LESS [42]. Given access to validation data for each target task, our method: (1) computes first-order gradient influence scores to measure how each training sample impacts task-specific performance, and (2) uses influence consensus through majority voting to identify training samples that show consistent positive value across multiple tasks. This consensus-based mechanism identifies universally valuable training examples: while some samples might be highly influential for individual tasks, we prioritize those that demonstrate broad utility across the task spectrum. While the computational cost of influence estimation is expensive, this front-loaded, one-time investment yields a standardized, compact dataset that can significantly accelerate iteration and development of both multimodal data and models.
Using ICONS, we create LLaVA-ICONS-133K, an automatically curated 20% subset of the LLaVA dataset [23]. This compact dataset maintains 98.6% of the original performance across multiple vision-language tasks, providing a 2.8 percentage point improvement over randomly selecting the same-sized subset (95.8%) and eliminating two-thirds of the performance drop from shrinking the training data. Moreover, our ICONS outperforms all baselines across different selection ratios, and remarkably achieves above-full-dataset performance, surpassing the original dataset at a 60% selection ratio. Importantly, LLaVA-ICONS-133K shows strong transferability, maintaining 95.5-113.9% relative performance across diverse unseen tasks such as chart understanding and infographic comprehension, suggesting that ICONS identifies fundamentally valuable training examples rather than task-specific patterns. We summarize our key contributions:
  1. 1.
    We propose ICONS, a simple yet effective approach for multi-task vision-language data selection that identifies broadly valuable training samples through majority voting across task-specific gradient influence scores.
  2. 2.
    Our voting-based aggregation consistently outperforms all baselines (§3.2) and, most notably, exceeds 102% of the full dataset performance at 60% selection ratio (§3.3). We further analyze different influence aggregation approaches and show the effectiveness of our consensus-aware selection over direct-merge alternatives (§3.5).
  3. 3.
    We release LLaVA-ICONS-133K, a compact 20% subset of LLaVA-665K, achieving near-full performance (98.6%), transferring well to unseen tasks (§3.4), and serving as a standardized training set for resource-efficient development of multimodal models.
2 Influence Consensus for Vision-Language Data Selection
We propose a consensus-driven, gradient-based data selection framework for visual instruction tuning datasets. We formalize the problem setup in §2.1 and establish gradient-based influence estimation preliminaries in §2.2. Our two-stage data selection framework consists of: first, the specialist stage (§2.3), which computes task-specific influence scores, followed by the generalist stage (§2.4), which builds cross-task consensus through voting-based aggregation.
Figure 2: ICONS. The Specialist stage (left) processes each task individually through three steps: (1) warmup training on a small subset of data, (2) gradient computation for both training and target task validation data, and (3) influence matrix computation to generate per-task influence scores. This process is repeated for each target task. The Generalist stage (right) performs Influence Consensus to aggregate information across tasks, where samples scoring above the $80^{\text{th}}$ percentile threshold for each task receive a vote. The final selection is made by summing votes across tasks and selecting the top 20% most influential samples, creating a compact yet highly effective training dataset that performs well across all tasks.
2.1 Problem Formulation
Given a large-scale visual instruction tuning dataset $\mathcal{D}=\{({\bm{x}}_{i},{\bm{I}}_{i},{\bm{y}}_{i})\}_{i=1}^{N}$ containing $N$ samples, where each data point ${\bm{z}}_{i}=({\bm{x}}_{i},{\bm{I}}_{i},{\bm{y}}_{i})$ includes natural language instruction ${\bm{x}}_{i}$ and an image ${\bm{I}}_{i}$, with corresponding target response ${\bm{y}}_{i}$ [^1] The framework supports multi-turn conversational data, yet we formalize the problem setup for single-turn instruction-tuning for clarity and simplicity., and given access to validation data $\mathcal{V}_{k}$ for each downstream task $T_{k}\in\mathcal{T}=\{T_{1},...,T_{K}\}$, our goal is to select a compact subset $\mathcal{S}\subset\mathcal{D}$ of size $M\ll N$ that maximizes model performance across multiple downstream tasks:
$\mathcal{S}^{*}=\operatorname*{arg\,max}_{\mathcal{S}\subset\mathcal{D},|\mathcal{S}|=m}\sum_{k=1}^{K}\text{Rel}(f_{\mathcal{S}},T_{k}),$ (1)
$\text{Rel}(f_{\mathcal{S}},T_{k})=\frac{\text{Score}(f_{\mathcal{S}},T_{k})}{\text{Score}(f_{\mathcal{D}},T_{k})},$ (2)
where $f_{\mathcal{S}}$ and $f_{\mathcal{D}}$ denote models trained on subset $\mathcal{S}$ and full dataset $\mathcal{D}$, respectively. $\text{Score}(f,T_{k})$ is the task-specific evaluation score achieved by model $f$ on task $T_{k}$. We define the average relative performance across all tasks as Rel. = $\sum_{k=1}^{K}\text{Rel}(f_{\mathcal{S}},T_{k})/K$. Rel. quantifies the subset-trained model’s performance relative to that of the model trained on the entire dataset, with values close to 1 indicating that the subset maintains the performance of full training [18]. Our objective is to find a subset where Rel. $\approx 1$, i.e., the model trained on the selected subset achieves performance comparable to using the full dataset across all tasks.
2.2 Preliminaries
Building on our problem formulation in §2.1, we formalize how to estimate the influence of training samples on downstream task performance. Specifically, since our goal is to maximize $\text{Rel}(f_{\mathcal{S}},T_{k})$ across tasks as defined in Eqn. 2, we need an efficient way to estimate how each training sample contributes to the $\text{Score}(f_{\mathcal{S}},T_{k})$ term in the numerator. Following [34, 42], we estimate how individual training samples $z_{i}\in\mathcal{D}$ affect validation loss by measuring their gradient alignment with reducing validation loss on $\mathcal{V}_{k}$, which directly impacts task-specific evaluation scores. Denote a training data point as ${\bm{z}}$ and a validation data point as ${\bm{z}}^{\prime}$ from validation set $\mathcal{V}_{k}$ for task $T_{k}$ When training with SGD and batch size 1, using data point ${\bm{z}}$ at timestep $t$ leads to a model update ${\bm{\theta}}_{t+1}={\bm{\theta}}_{t}-\eta_{t}\nabla\ell({\bm{z}};{\bm{\theta}}_{t})$, where $\eta_{t}$ is the learning rate. To reduce the computational cost, we use the first-order Taylor expansion to estimate the loss on a given validation data point ${\bm{z}}^{\prime}$ at time step $t+1$ as:
$\ell({\bm{z}}^{\prime};{\bm{\theta}}_{t+1})\approx\ell({\bm{z}}^{\prime};{\bm{\theta}}_{t})+\langle\nabla\ell({\bm{z}}^{\prime};{\bm{\theta}}_{t}),{\bm{\theta}}_{t+1}-{\bm{\theta}}_{t}\rangle.$
The influence of a training data point ${\bm{z}}$ on a validation data point ${\bm{z}}^{\prime}$ is:
$\mathcal{I}_{t}({\bm{z}}\rightarrow{\bm{z}}^{\prime})$ $=\ell({\bm{z}}^{\prime};{\bm{\theta}}_{t+1})-\ell({\bm{z}}^{\prime};{\bm{\theta}}_{t})$
$\approx-\eta_{t}\langle\nabla\ell({\bm{z}}^{\prime};{\bm{\theta}}_{t}),\nabla\ell({\bm{z}};{\bm{\theta}}_{t})\rangle,$
which we refer to as an influence score. For our case of visual instruction tuning [23] which uses single-epoch training, the influence score is:
$\mathcal{I}_{\text{single}}({\bm{z}}\rightarrow{\bm{z}}^{\prime})=\eta\langle\nabla\ell({\bm{z}}^{\prime};{\bm{\theta}}_{t}),\nabla\ell({\bm{z}};{\bm{\theta}}_{t})\rangle,$
where $\eta$ is the learning rate used during training. This can be generalized to multi-epoch training by accumulating the influence scores over all $n$ epochs:
$\mathcal{I}_{\text{total}}({\bm{z}}\rightarrow{\bm{z}}^{\prime})=\sum_{i=1}^{n}\eta_{i}\langle\nabla\ell({\bm{z}}^{\prime};{\bm{\theta}}_{i}),\nabla\ell({\bm{z}};{\bm{\theta}}_{i})\rangle,$
where $\eta_{i}$ is the learning rate used during the $i$th epoch. The gradient-based selection approach selects training samples $\{{\bm{z}}\}$ that maximize the gradient inner product $\langle\nabla\ell({\bm{z}}^{\prime};{\bm{\theta}}_{t}),\nabla\ell({\bm{z}};{\bm{\theta}}_{t})\rangle$ [^2] In practice, we use cosine similarity instead of direct inner products to avoid biasing selection toward shorter sequences, since gradient norms tend to be inversely correlated with sequence length as noted in [42]. through a greedy, first-order approximation, which leads to larger reductions in validation loss for point ${\bm{z}}^{\prime}$. While this approach omits second-order terms compared to influence functions [17], it provides an efficient approximation for ranking the impact of training samples [11, 42].
2.3 Specialist: Individual Task Influence Ranking
To rank the influence of training data for each target task, we compute the influence score of each training data point on a validation set that represents the target task distribution. Following LESS [42], the process involves three steps: (1) training the model on 5% randomly selected data as a lightweight warm-up phase to initialize visual instruction-following capabilities, (2) computing gradients for training and validation data and compressing the gradients via random projection, and (3) computing the influence score to quantify the impact of each training data on validation set.
Step 1: Warm-up Training.  Following LESS [42], we first perform a brief warmup phase using Low-Rank Adaptation (LoRA) [12] on a small random subset $\mathcal{D}_{\mathrm{warmup}}\subset\mathcal{D}$ (5% of the training data): $f_{\text{warmup}}=\text{LoRA}(f_{\text{base}},\mathcal{D}_{\text{warmup}})$. This warmup phase allows the model to develop basic instruction-following capabilities.
Step 2: Gradient Computation.  For each training data ${\bm{z}}_{i}\in\mathcal{D}$ and validation data ${\bm{z}}^{\prime}_{j}\in\mathcal{D}_{\text{val}}^{k}$ from $\mathcal{T}_{k}$, we compute their gradients with respect to $f_{\text{warmup}}$ parameters $\theta_{w}$:
$g_{i}=\nabla_{\theta_{w}}\mathcal{L}(f_{\text{warmup}}({\bm{z}}_{i}),{\bm{y}}_{i}),\quad g^{\prime}_{j}=\nabla_{\theta_{w}}\mathcal{L}(f_{\text{warmup}}({\bm{z}}^{\prime}_{j}),{\bm{y}}^{\prime}_{j})$
where $y_{i}$ and $y^{\prime}_{j}$ are the targets for $z_{i}$ and $z^{\prime}_{j}$, respectively. In order to reduce computational and storage overhead, we apply random projection to the gradient feature: $\tilde{g}_{i}=Rg_{i}$ and $\tilde{g}^{\prime}_{j}=R^{\prime}g^{\prime}_{j}$, where $R,R^{\prime}\in\mathbb{R}^{d^{\prime}\times d}$ is a random projection matrix with $d^{\prime}\ll d$ that preserves inner products with high probability [14]. We further normalize the projected gradient features, $\tilde{g}_{i}=\frac{\tilde{g}_{i}}{\|\tilde{g}_{i}\|_{2}},\tilde{g}^{\prime}_{j}=\frac{\tilde{g}^{\prime}_{j}}{\|\tilde{g}^{\prime}_{j}\|_{2}}$ to prevent bias from sequence length differences [42].
Step 3: Influence Matrix Computation.  We compute the influence matrix $I\in\mathbb{R}^{|\mathcal{D}|\times|\mathcal{D}_{\text{val}}^{k}|}$ where each entry $I_{ij}=\langle\tilde{g}_{i},\tilde{g}^{\prime}_{j}\rangle$ is the influence of training data ${\bm{z}}_{i}$ on validation data ${\bm{z}}^{\prime}_{j}$, and the average influence of training data ${\bm{z}}_{i}$ on the target task $k$ is then computed as:
$\bar{I}_{k}({\bm{z}}_{i})=\frac{1}{|\mathcal{D}_{\text{val}}^{k}|}\sum_{j=1}^{|\mathcal{D}_{\text{val}}^{k}|}I_{ij}.$ (3)
This influence estimation process provides a task-specific ranking for the training set $\mathcal{D}$ with respect to task $\mathcal{T}_{k}$, where a higher influence score $\bar{I}_{i}$ suggests a higher influence for the target task. However, recall that our ultimate goal is to select a compact subset $\mathcal{S}\subset\mathcal{D}$ of size $M\ll N$ that maximizes $\sum_{k=1}^{K}\text{Rel}(f_{\mathcal{S}},T_{k})/K$ across all tasks. We address this disconnection between task-specific rankings and overall dataset optimization by proposing a voting-based generalist approach; this approach aggregates influence scores for each target task to identify the most broadly impactful training data.
2.4 Generalist: Cross-Task Influence Consensus
Our consensus-based voting strategy identifies training samples that consistently show a high influence score across various tasks. For each task $k$, we set a task-specific threshold $\tau_{k}$ at the $((1-p)\times 100)$-th percentile of the influence score distribution within that task, where $p$ represents the proportion of top samples to consider ($p=0.2$ in our main experiments). A sample ${\bm{z}}_{i}$ is considered important for task $k$ if its influence score $I_{k}({\bm{z}}_{i})\geq\tau_{k}$. Formally, we define the voting-based influence score for sample $z_{i}$ on task $k$ as:
$I_{\text{vote}}({\bm{z}}_{i})=\sum_{k=1}^{K}\mathbbm{1}[I_{k}({\bm{z}}_{i})\geq\tau_{k}],$ (4)
where $\mathbbm{1}[\cdot]$ denotes the indicator function, and $I_{k}({\bm{z}}_{i})$ represents the influence score of sample ${\bm{z}}_{i}$ for task $k$. The final score for each sample is the total number of votes it receives across all tasks, ranging from 0 (indicating no importance for any task) to $K$ (indicating consistent importance for all tasks). Influence Consensus identifies a unique subset of samples that are consistently influential across multiple tasks, rather than those that might be highly influential in only a few tasks. After computing the aggregated influence scores, we rank the training samples and select the top $p\%$ as final subset for model training.
While directly combining validation data from all target tasks and computing a single influence matrix (Direct-merge Selection, §3.5) might seem intuitive, we adopt this consensus-based approach for three reasons. First, specialist selection provides a practical performance upper bound, as the performance achieved by selecting the top $p$of data specifically for a single task represents the theoretical maximum achievable with that data budget, serving as a critical baseline for evaluating generalist selection trade-offs (§3.5). Second, this approach is scalable and naturally supports continual learning scenarios, as influence matrices for new tasks or benchmarks can be computed independently without recomputing training gradients, thereby reducing computational costs. Third, computing influence matrices independently for different tasks give us better understanding of cross-task influence dynamics. Our analysis (§3.3) reveals that the influence ranking of training data vary significantly across different tasks, with highly influential samples for one task often showing limited impact on others. This further provides insights to design more effective consensus mechanisms for generalist selection while maintaining computational tractability.
Method VQAv2 GQA VizWiz SQA-I TextVQA POPE MME MMBench LLaVA-W Rel. (%)
en cn Bench
Full 79.1 63.0 47.8 68.4 58.2 86.4 1476.9 66.1 58.9 67.9 100
Random 75.7 58.9 44.3 68.5 55.3 84.7 1483.0 62.2 54.8 65.0 95.8
CLIP-Score [35] 73.4 51.4 43.0 65.0 54.7 85.3 1331.6 55.2 52.0 66.2 91.2
EL2N [33] 76.2 58.7 43.7 65.5 53.0 84.3 1439.5 53.2 47.4 64.9 92.0
Perplexity [28] 75.8 57.0 47.8 65.1 52.8 82.6 1341.4 52.0 45.8 68.3 91.6
SemDeDup [1] 74.2 54.5 46.9 65.8 55.5 84.7 1376.9 52.2 48.5 70.0 92.6
D2-Pruning [26] 73.0 58.4 41.9 69.3 51.8 85.7 1391.2 65.7 57.6 63.9 94.8
Self-Sup [38] 74.9 59.5 46.0 67.8 49.3 83.5 1335.9 61.4 53.8 63.3 93.4
Self-Filter [5] 73.7 58.3 53.2 61.4 52.9 83.8 1306.2 48.8 45.3 64.9 90.9
COINCIDE [18] 76.5 59.8 46.8 69.2 55.6 86.1 1495.6 63.1 54.5 67.3 97.4
ICONS (ours) 76.3 60.7 50.1 70.8 55.6 87.5 1485.7 63.1 55.8 66.1 98.6
Table 1: Baseline Comparisons. Performance comparison of different data selection approaches when trained on 20% of the LLaVA-665K dataset. Models are evaluated across diverse multimodal benchmarks, with tasks ranging from visual question answering (VQAv2, GQA) to multilingual understanding (MMBench). The best and second best results for each benchmark are shown in bold and underlined, respectively. Our method ICONS achieves the highest overall relative performance (98.6%), consistently outperforming existing approaches including COINCIDE (97.4%) and D2-Pruning (94.8%), while methods like EL2N, Perplexity, and CLIP-Score show limited effectiveness with relative performance around 91-92%.
3 Experiments
In this section, we first discuss our experiment setup and evaluation benchmarks (§3.1). Then we compare it with the latest state-of-the-art methods (§3.2). Next, we provide detailed analysis of our method’s behavior and effectiveness across different dimensions (§3.3). We further evaluate the transferability of our selection approach across both unseen tasks and model architectures (§3.4). Finally, we conduct ablation studies to understand the contribution (§3.5).
3.1 Evaluation Test-Bed
Dataset & Model.  We evaluate ICONS on visual instruction tuning dataset LLaVA-665K [22]. We use LLaVA-v1.5-7b-lora model [22] with a default size of 7B parameters unless otherwise specified. In all experiments, we train the models for one epoch following the official finetuning hyperparameters, which reduces computational overhead while maintaining selection quality.
Target Tasks.  We evaluate the effectiveness of ICONS with a wide range of multimodal benchmarks (Tab. 2) that test different capabilities of vision-language models. The benchmarks include: 1) Multiple-choice understanding: MMBench [47] and MME [7], which assess the model’s ability to select correct answers from given options; 2) Visual question answering: VQAv2 [9], GQA [13], and VizWiz [10], which test basic visual reasoning; 3) Text understanding in images: TextVQA [37], which evaluates OCR capabilities; 4) Scientific reasoning: ScienceQA [25], which tests knowledge-grounded question answering; 5) Open-ended generation: LLaVA-W Bench [23], which evaluate free-form response generation; 6) Factual consistency: POPE [21], which measures hallucination.
MME POPE SQA-I MMBench
en cn
$|\mathcal{D}_{\text{val}}|$ 986 500 424 1,164 1,164
$|\mathcal{D}_{\text{test}}|$ 2,374 8,910 4,241 1,784 1,784
Task Type Y/N Y/N MCQ MCQ MCQ
VQAv2 GQA VizWiz TextVQA LLaVA-W
$|\mathcal{D}_{\text{val}}|$ 1,000 398 8,000 84 84
$|\mathcal{D}_{\text{test}}|$ 36,807 12,578 8,000 5,000 84
Task Type VQA VQA VQA VQA VQA
Table 2: Statistics of Target Tasks. Our target tasks include diverse benchmarks and answer formats, covering different vision-language capabilities. Task types include Multiple-Choice Questions (MCQ), Visual Question Answering (VQA), and Yes/No Questions (Y/N).
3.2 Comparisons with Baselines
We compare our ICONS against several baseline approaches, including random selection. We evaluate CLIP-Score [35] for measuring image-text alignment, EL2N [33] based on embedding L2 norms, and Perplexity [28] using language model scores. We also compare against SemDeDup [1] for semantic deduplication and D2-Pruning [26] for distribution-aware pruning. Additional baselines include Self-Sup [38] leveraging self-supervised signals, while Self-Filter [5] and COINCIDE [18] are the approaches specifically designed for vision-language data selection. All methods are evaluated using a 20% sampling ratio of the original dataset, with results presented in Tab. 1. We reference the baseline results from COINCIDE [18].
We compare our ICONS against existing data selection approaches in Tab. 1. Our method achieves the best overall performance with 98.6% Rel., outperforming all baselines with LLaVA-ICONS-133K, 20% of the training data. Remarkably, we achieve comparable or better performance than training with the entire LLaVA-665K on several tasks: SQA-I (70.8 vs. 68.4), MME (1485.7 vs. 1476.9), VizWiz (50.1 vs. 47.8) and POPE (87.5 vs. 86.4). Among the baselines, COINCIDE achieves strong performance (97.4% Rel.), though it falls short of ICONS on key tasks. Other methods like EL2N, Perplexity, SemDeDup are less effective, with relative scores around 91-92%, showing limitations in preserving performance.
Task Full Specialist Generalist Delta (%)
VQAv2 [9] 79.1 77.1 76.3 1.04
GQA [13] 63.0 61.1 60.7 0.65
VizWiz [10] 47.8 53.1 50.1 5.65
SQA-I [25] 68.4 69.8 70.8 -1.43
TextVQA [37] 58.2 55.7 55.6 0.18
POPE [21] 86.4 86.6 87.5 -1.04
MME [7] 1476.9 1506.1 1485.7 1.35
MMBench (en) [47] 66.1 66.0 63.1 4.34
MMBench (cn) [47] 58.9 56.4 55.8 1.06
LLaVA-W Bench [23] 67.9 67.1 66.1 1.49
Table 3: Single-task Selection (Specialist) vs. Consensus-aware Multi-task Selection (Generalist). The single-task data selection approach selects 20% of data per task, while our consensus-aware multi-task data selection approach selects a total of 20% data across all tasks. The Delta (%) column represents the relative drop in performance from specialist to generalist.
Figure 3: Data Overlap between Specialists and Generalist Selection. It shows varying degrees of overlap, ranging from 3.27% (VQAv2 [9]) to 24.21% (LLAVA-W Bench [23]).
Figure 4: Performance Comparison at Different Selection Ratios. Here we show the average relative performance comparisons of various data selection methods at different selection ratios. Our ICONS consistently outperforms all baseline approaches across different selection ratios, achieving 99.2% relative performance with 40% of the data, and remarkably exceeding 102% at 60% selection ratio. This shows that our selected subset is not only more efficient but can be even more effective than the full dataset for model training.
3.3 Analysis
From Specialist to Generalist.  In order to understand the effectiveness of the intermediate task-specific influence matrices we obtained from the specialist stage, we further conduct experiments on selecting 20% of data for individual tasks. As shown in Tab. 3, using 20% task-specific data, it achieves comparable or even superior performance to full data training in several cases (e.g., MME from 1476.9 to 1506.1 and SQA-I from 68.4 to 69.8). Then, through our influence consensus mechanism in the generalist stage, we identify consistently influential cross-tasks samples and select a 20% subset. This generalist approach maintains remarkably strong performance, with an average drop of 1.33% across most tasks compared to specialist baselines.
As visualized in Fig. 3, the overlap between specialist and generalist selections varies significantly across tasks, ranging from minimal overlap in simpler tasks like VQAv2 [9] (3.27%) and VizWiz [10] (3.28%) to substantial agreement in more complex tasks like LLAVA-W Bench [23] (24.21%). Notably, some tasks even show performance improvements under the generalist selection, such as SQA-I [25] (improving by 1.43%) and POPE [21] (improving by 1.04%). This empirically validates our consensus-based approach: by first understanding task-specific influence patterns and then building consensus across tasks, we can effectively identify a compact, universal training set that maintains strong performance across diverse vision-language tasks while significantly reducing the required training data.
Different Selection Ratio.  We evaluate our ICONS against baseline methods across different selection ratios, ranging from 5% to 60% of the training data. As shown in Fig. 4, our results reveal several key patterns: First, ICONS shows particularly strong performance in the low-selection regime (5-20%), where identifying the most influential samples is crucial. Another notable observation is that as the selection ratio increases, the performance gap between different methods gradually narrows. This convergence pattern is expected, as larger sample sizes naturally capture more of the dataset’s diversity and information. Despite this convergence trend, ICONS consistently maintains its performance edge across all selection ratios, and remarkably reaching above 100% at 60% selection ratio. One hypothesis is that ICONS can also effectively filter out potentially harmful or noisy training samples that might negatively impact model training, thereby surpassing the full training performance.
(a) Pairwise Overlap Heatmap between Specialist Benchmarks. The values are the percentage of selected overlapping samples across different benchmarks, with high overlap seen between POPE and GQA (60.2%) and between VQAv2 and VizWiz (49.0%). Lower overlap, such as 3.3% between MMBench(en) and GQA, shows more distinct selected sample sets.
(b) Pairwise Overlap Heatmap between Different Aggregation Approaches. The values are the percentage of overlapping selected samples between each pair of aggregation approaches. Our Vote-based influence consensus approach shows moderate overlap (53.8%-70.5%) with other methods, while Mean and Max selections are highly correlated (99.9%).
Figure 5: Pairwise Overlap Heatmaps between Different Aggregation Approaches and Benchmarks. (a) compares the overlap between different specialist benchmarks, while (b) shows overlap between various aggregation methods.
Divergent Cross-Task Influence Patterns.  As shown in Fig. 5(a), the pairwise overlap heatmap reveals significant variation in training data influence rankings across different tasks. High overlap, such as between VQAv2 and VizWiz (49.0%) or POPE and GQA (60.2%), suggests that certain samples are beneficial across similar tasks. However, low overlap, like the 3.3% between MMBench (en) and GQA, highlights that highly influential samples for one task may have limited impact on others. Even closely related tasks, such as MMBench in different languages (English and Chinese), share 67.4% of influential samples. These findings empirically demonstrate significant overlap in influential samples across tasks, motivating the use of influence consensus for effective multi-task data selection.
3.4 Transferability
We evaluate the transferability of our data selection approach across two dimensions: unseen-task generalization and cross-architecture-scale generalization. Unseen-task Generalization.  LLaVA-ICONS-133K demonstrates exceptional generalization on entirely unseen benchmarks that were not considered during our data selection process. As shown in Tab. 4, we test across a diverse spectrum of tasks including MMVet [45], NaturalBench [19], AI2D [16], ChartQA [29], DocVQA [30], InfoVQA [31], RealWorldQA [41] and CMMMU [46]. LLaVA-ICONS-133K achieves 95.5-113.9% (Rel.) compared to full dataset training. Notably, we observe improvements on InfoVQA (103.8%), NaturalBench (105.5%), RealWorldQA (104.4%), and CMMMU (113.9%), suggesting that, in some cases, training on LLaVA-ICONS-133K may even outperform training on the full dataset, despite these tasks not being included in the selection process. Importantly, LLaVA-ICONS-133K significantly outperforms random selection across all benchmarks - for example, improving performance on ChartQA (17.1 vs. 15.1), MMVet (29.7 vs. 27.6), CMMMU (25.2 vs. 21.9) and NaturalBench (12.8 vs. 11.1). This consistent improvement over random selection further validates the effectiveness of ICONS. This suggests that our selection approach successfully captures fundamental visual-language understanding capabilities that transfer well across different task formats and domains, and consistently outperforms random selection by a significant margin.
AI2D ChartQA DocVQA InfoVQA MMVet Naturalbench RealworldQA CMMMU Rel. (%)
Full 55.4 17.5 28.9 26.5 31.1 12.4 52.4 22.1 100.0
Random 50.2 15.1 25.2 24.3 27.6 11.1 49.8 21.9 91.6
LLaVA-ICONS-133K 53.9 17.1 27.9 27.5 29.7 12.8 55.0 25.2 98.7
Per-task Rel. (%) 97.3 97.7 96.5 103.8 95.5 103.2 104.4 114.0 -
Table 4: Unseen-Task Generalization. Performance comparison on unseen benchmarks when trained on the selected subset. The relative performance ratios show that our selection approach successfully captures fundamental visual-language understanding capabilities that transfer well across different task formats and domains. Notably, we observe improvements on InfoVQA [31] (103.8%), MMVet [45] (95.5%), NaturalBench [19] (103.2%), RealWorldQA [41] (104.4%), and CMMMU [46] (114.0%), suggesting that in some cases, training on our LLaVA-ICONS-133K may even outperform using the full dataset.
Cross-Architecture-Scale Generalization.  We further conduct experiments on cross architecture scale generalization to evaluate the transferability of our selected data across different model scales. While our subset was initially selected using LLaVA-1.5-7B as the base model, we investigate whether these same examples remain effective for training larger architectures like LLaVA-1.5-13B. This tests whether our selection criteria identify universally valuable training examples rather than model-specific patterns. Our results in Tab. 5 show cross architecture scale generalization, with the 7B-selected achieving 98.1% relative performance, suggesting that our selected subset captures fundamental visual-language understanding capabilities that scale effectively across model sizes.
VQAv2 GQA Vizwiz SQA-I TextVQA POPE MME MMBench MMBench(cn) LLAVA-W Rel. (%)
Full 80.0 63.3 58.9 71.2 60.2 86.7 1541.7 68.5 61.5 69.5 100.0
Random 77.3 60.7 57.6 69.1 56.8 82.9 1517.2 63.2 56.3 67.5 95.7
7B-selected 78.8 60.4 57.4 70.4 58.3 84.3 1527.5 64.9 59.7 68.2 97.3
13B-selected 78.9 61.2 57.5 71.3 58.4 85.9 1535.2 66.1 59.8 68.8 98.1
Table 5: Cross-Architecture-Scale Generalization. Our LLaVA-ICONS-133K selected via LLaVA-v1.5-7b-lora (7B-selected) shows strong cross-architecture transferability, achieving competitive performance when trained with data selected via LLaVA-v1.5-13b-lora (13B-selected). Specifically, while the 7B-selected achieves 97.3% relative performance, the 13B-selected model reaches 98.1%, showing that our selected subset’s effectiveness scales well to larger architectures. Both significantly outperform random selection (95.7%), with the 13B-selected option showing particular strength in complex reasoning tasks like MMBench and POPE. This consistency across model scales highlights that our data selection approach captures fundamental visual-language understanding patterns that generalize well across different model architecture scales.
3.5 Ablation Studies
Different Aggregation Approaches.  To understand how different methods of combining task-specific influence scores affect the performance, we compare our majority voting approach (Vote) with four aggregation approaches: (1) Mean, which averages scores across tasks, (2) Max, which selects the highest score per sample, (3) Rank, which averages the relative ranking of samples within each task, (4) Norm, which normalize the scores before averaging:
$I_{\text{mean}}({\bm{z}}_{i})=\frac{1}{K}\sum_{k=1}^{K}I_{k}({\bm{z}}_{i})\qquad I_{\text{norm}}({\bm{z}}_{i})=\frac{1}{K}\sum_{k=1}^{K}\frac{I_{k}({\bm{z}}_{i})-\mu_{k}}{\sigma_{k}}$
$I_{\text{max}}({\bm{z}}_{i})=\max_{k=1,\ldots,K}I_{k}({\bm{z}}_{i})\qquad I_{\text{rank}}({\bm{z}}_{i})=\frac{1}{K}\sum_{k=1}^{K}\text{rank}(I_{k}({\bm{z}}_{i}))$
Results in Tab. 6 show that our Vote consistently outperforms alternatives across most tasks, achieving the highest overall relative performance (Rel.) of 98.6%. While methods like Rank perform well for MME (1490.0 vs. 1485.7 with Norm), they underperform overall. To better understand the relationships between these aggregation approaches, we visualize their pairwise selection overlap in Fig. 5(b). Specifically, the Mean and Max approaches show an extremely high overlap of 99.9%, suggesting they tend to select very similar influential samples. In contrast, our Vote-based approach exhibits a more moderate overlap with other methods, ranging from 53.8% to 70.5%, potentially those that are influential across a broader variety of tasks rather than overfitting to specific tasks.
Task Mean Max Rank Norm Vote (ours)
VQAv2 [9] 75.7 75.2 74.9 75.1 76.3
GQA [13] 59.6 59.8 58.6 60.1 60.7
VizWiz [10] 47.9 48.1 40.5 46.4 50.1
SQA-I [25] 65.5 66.2 69.8 69.8 70.8
TextVQA [37] 55.5 55.5 55.2 54.5 55.6
POPE [21] 86.0 85.5 85.6 85.6 87.5
MME [7] 1422.1 1470.7 1490.0 1482.6 1485.7
MMBench (en) [47] 59.0 58.3 59.0 58.9 63.1
MMBench (cn) [47] 51.0 51.8 50.8 52.5 55.8
LLaVA-W Bench [23] 66.2 66.2 66.4 66.3 66.1
Rel.(%) 96.4 96.1 95.9 96.8 98.6
Table 6: Comparison of Different Aggregation Approaches. Performance comparison of different methods for combining task-specific influence scores, using 20% of training data. All scores are relative to full dataset training.
Task Full Direct Consensus Delta (%)
merge aware (ours)
VQAv2 [9] 79.1 76.1 76.3 0.26
GQA [13] 63.0 59.4 60.7 2.19
VizWiz [10] 47.8 46.1 50.1 8.67
SQA-I [25] 68.4 68.7 70.8 3.06
TextVQA [37] 58.2 54.1 55.6 2.77
POPE [21] 86.4 85.1 87.5 2.82
MME [7] 1476.9 1419.2 1485.7 4.69
MMBench (en) [47] 66.1 61.9 63.1 1.94
MMBench (cn) [47] 58.9 50.3 55.8 10.94
LLaVA-W Bench [23] 67.9 65.2 66.1 1.38
Rel. (%) 100 94.7 98.6 -
Table 7: Comparison of Consensus-aware vs. Direct-merge Selection. Performance comparison across different benchmarks using 20% of training data. Direct-merge computes influence directly on combined validation sets, while consensus-aware refers to our specialist-to-generalist approach. The Delta (%) column represents the relative improvement from Direct-merge to Consensus-aware. Consensus-aware outperforms Direct-merge across all benchmarks, while Direct-merge notably underperforms even the random baseline.
Consensus-aware Selection vs. Direct-merge Selection.  To validate our choice of a consensus-aware approach (first computing task-specific influence matrices and then aggregating results) versus a direct-merge approach (directly computing influence using a combined validation set), we compare the performance of both approaches. In the direct-merge approach, we combine validation sets from all tasks into a single pool, denoted as $\{{\bm{z}}^{\prime}\}_{\mathrm{merged}}$, and compute influence scores directly against this combined set: $\mathcal{I}_{\text{direct}}({\bm{z}}\rightarrow\{{\bm{z}}^{\prime}\}_{\mathrm{merged}})$. As described in Tab. 7, our consensus-aware approach consistently outperforms direct-merge selection across all benchmarks, with the largest gains observed on complex tasks like MME [7].
A potential reason could be the direct-merge approach does not account for the different vision-language task difficulties or validation set sizes across the target tasks. By pooling all validation examples, it effectively treats each validation data from all tasks equally, which can lead to the influence scores being dominated by tasks with more samples or with higher influence scores on average. The task-aware voting-based influence consensus mechanism of our direct-merge approach are more robust to these issues. Additionally, our approach is more scalable, as it avoids computing a single, large influence matrix across validation data from all target tasks. Adding new tasks only requires computing influence scores for those specific tasks and updating the voting mechanism, while direct-merge selection would need to recompute the entire large influence matrix which makes it less practical.
4 Related Work
Visual Instruction Tuning.
Multimodal large language models (MLLMs), e.g., Flamingo [2], LLaVA [23], BLIP2 [20], and Cambrian [39], enhance the capabilities of large language models (LLMs) on various multimodal tasks. A key component in advancing MLLMs is visual instruction tuning [23], a training process that enables these models to interpret and follow instructions within a vision-language context, transforming them into versatile multimodal assistants. This tuning process not only improves the models’ instruction-following abilities but also aligns their outputs more closely with user expectations, thus enhancing their utility in practical applications [23].
Data Selection.
Data selection methods can be categorized based on the types of information they utilize for selection. Representation-based approaches [1, 18] leverage neural embeddings to capture data representations. Loss trajectory-based methods [32] prioritize data points that contribute most significantly to reducing generalization error over training. Gradient-based techniques [33, 42, 6] select data based on gradient information. Recent work has explored various approaches to select optimal visual instruction tuning datasets. Concurrent work TIVE [24] employs gradient-based selection to identify representative instances. TIVE assumes that the number of specialist data should be proportional to task difficulty and thus samples specialist data based on an estimation of task difficulty. Our method does not rely on this assumption – we directly select samples that benefit the greatest number of tasks. COINCIDE [18] clusters data based on representations associated with concept-skill compositions. Our work follows targeted instruction tuning selection approach LESS [42] to utilize gradient information to calculate the specialist influence (i.e., the influence on a specific task) and extends it to general scenarios by aggregating information from various tasks and selecting data for multiple downstream tasks via majority voting.
5 Conclusion
In this work, we introduce ICONS, a simple yet effective influence consensus-based approach for visual instruction tuning data selection. Our two-stage specialist-to-generalist method, first computes task-specific influence matrices and then select a subset by prioritizing universally influential samples through a cross-task voting mechanism. Through extensive experiments, we show that ICONS achieving 98.6% of full dataset performance using only 20% of the data and it consistently outperform baseline methods across different selection ratios. Additionally, we release LLaVA-ICONS-133K, a 20% subset of LLaVA-665K dataset, that not only maintains strong performance across diverse vision-language tasks but also shows transferability to unseen tasks. We hope our work inspires further exploration into data-efficient methods for vision-language models across diverse applications.
Limitations.  Our approach primarily faces one practical limitation: computing gradients for large training datasets is computationally expensive (§B.2). This computational overhead could potentially constrain the method’s applicability when working with extremely large-scale datasets. To support broader research community, we release our LLaVA-ICONS-133K dataset to help research iteration and model development under resource-constrained settings.
Acknowledgments.  This material is based upon work supported by the National Science Foundation under Grant No. 2107048 and No.2112562. Any opinions, findings, and conclusions, or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the National Science Foundation. This work is also supported by the Singapore National Research Foundation and the National AI Group in the Singapore Ministry of Digital Development and Information under the AI Visiting Professorship Programme (award number AIVP-2024-001). We thank many people for their helpful discussion and feedback, listed in alphabetical order by last name: Allison Chen, Hamish Ivison, Carlos E. Jimenez, Polina Kirichenko, Jaewoo Lee, Tiffany Ling, Zhiqiu Lin, Shengbang Tong, Ethan Tseng, Justin Wang, Zirui Wang.
Appendix for ICONS: Influence Consensus for Vision-Language Data Selection
In this supplement, we provide additional experiment details and ablation studies (§A) and additional analysis (§B). We detail our algorithmic framework (§C) and discuss future research directions (§D). Finally, we include visualization of the selected data (§E) to better understand data samples with high influence.
Appendix A Additional Experiment Details & Ablations
Projection Dimension.  We primarily set the projection dimension to 5120, reducing features from 338.7M to 5120 dimensions. The choice of 5120 was empirically validated for its trade-off between effective capturing gradient representation and maintaining a manageable parameter space. Our LLaVA-v1.5-7b-lora architecture includes a total of 7.4B parameters, with 338.7M parameters being trainable after LoRA adaptation, accounting for approximately 4.58% of the total parameter count. We further ablate different projection dimensions (1024, 2560, 5120, and 10240), with results provided in Fig. 6.
Warm-up Ratio.  To initiate training, we use a warm-up set including 5% of the total training data. We conducted ablation studies to evaluate the impact of varying warm-up ratios (5%, 10%, 20%, and 100%) on selection performance, as shown in Fig. 7. Our experiments reveal that increasing the warm-up data size does not lead to performance improvements. Surprisingly, models trained with smaller warm-up ratios (5-20%) consistently outperform those trained with the full dataset (100%). Specifically, the 5% warm-up ratio achieves the best performance at 98.6%, while using the complete dataset results in a performance drop to 97.8%. This finding suggests that a small subset of training data is sufficient and even beneficial for model initialization, and potentially gives better signals in the early training stages.
Figure 6: Projection Dimension Ablation. We show the performance of ICONS at different projected dimensions (1024, 2560, 5120, 10240), compared to the random baseline. The performance increases with the projected dimension and reaches a plateau around dimension 5120.
Figure 7: Warm-up Ratio Ablation. The blue line represents ICONS performance across different warm-up ratios (5%, 10%, 20%, and 100%), while the red dashed line shows the random baseline performance. Results show that smaller warm-up ratios (5-20%) achieve better performance compared to using the full dataset (100%).
Appendix B Additional Analysis
B.1 Consistency Analysis
To evaluate the consistency of our approach, we conduct three independent runs of the experiment. As shown in Fig. 8, our method demonstrates high consistency across different runs, achieving 98.6$\pm$1.2Rel., which shows a notable improvement over the random baseline, which achieves 95.8$\pm$2.7%. The lower standard deviation in our results (1.2% vs 2.7%) further indicates that our approach produces more stable and reliable outcomes compared to the random baseline.
Figure 8: Rel. (%) Across Runs. We show the performance across three different runs for random selection and our ICONS. Our ICONS achieves 98.6$\pm$1.2while the random baseline reaches 95.8$\pm$2.7%.
B.2 Computational Complexity
Computing gradient-based influence requires significant computational resources. In the specialist stage, the complexity scales with both the dataset size $|\mathcal{D}|$ and the gradient dimension $d$. This stage consists of three steps. First, the warm-up training has a complexity of $\mathcal{O}(|\mathcal{D}_{\text{warmup}}|)$. Second, the gradient computation stage has a computational complexity of $\mathcal{O}(|\mathcal{D}|+|\mathcal{D}_{\text{val}}|)$ for forward and backward passes, with storage requirements of $\mathcal{O}(|\mathcal{D}|\cdot d+|\mathcal{D}_{\text{val}}|\cdot d)$ for the gradients. Third (and finally), the influence matrix computation requires $\mathcal{O}(|\mathcal{D}|\cdot|\mathcal{D}_{\text{val}}|\cdot d^{\prime})$ compute cost, where $d^{\prime}$ is the reduced dimension after projection. The generalist stage, focusing on influence consensus across tasks, has lower computational requirements. It begins with threshold computation, requiring $\mathcal{O}(K\cdot|\mathcal{D}|\log|\mathcal{D}|)$ operations for sorting across $K$ tasks. The voting process then takes $\mathcal{O}(K\cdot|\mathcal{D}|)$ compute, followed by a final selection step with complexity $\mathcal{O}(|\mathcal{D}|\log|\mathcal{D}|)$ for sorting the aggregated votes. Storage requirements for this stage are minimal, primarily needing space for the final selected subset.
In practice, for LLaVA-665K training data, the warmup training phase requires 0.75 hours using eight L40 GPUs. We parallelize the gradient computation across 100 A6000 GPUs, taking approximately one hour and requiring 103GB of total storage for the gradients. The influence consensus stage is notably efficient, completing in less than a minute on a single L40 GPU. While these computational demands are substantial, they represent front-loaded, one-time costs that can be used across multiple target tasks and model iterations. This makes our method extendable for new tasks, as the expensive training data gradient computations only need to be performed once.
B.3 Visual Dependency Influence Ranking
Recent work [39] has shown that vision-language tasks vary in their reliance on visual information: task like MMBench [47] depends heavily on visual grounding, while others like SQA-I [25] can be handled primarily through language, showing only a 5% drop in performance when visual input is removed [39]. To take visual dependency of training data into consideration, we further explored gradient-based Visual Dependency Score (VDS). For each data point, we calculate the gradient of the model’s auto-regressive cross-entropy loss with both the original image and a Gaussian noise image $I_{\text{noise}}\sim\mathcal{N}(0,1)$, keeping the text input constant. This quantifies how much the visual component contributes to the model performance. We construct an adapted influence matrix: visual influence matrix $\mathcal{I}_{\texttt{VDS}}\in\mathbb{R}^{|\mathcal{D}|\times|\mathcal{D}_{\text{val}}|}$, which quantifies the visual influence of each training sample ${\bm{z}}_{i}$ on each validation sample ${\bm{z}}^{\prime}_{j}$ with respect to the model’s gradient alignment and visual dependency. $\mathcal{I}_{\text{VDS}}$ is computed as:
$\mathcal{I}_{\texttt{VDS},ij}=\langle\nabla_{\theta}\mathcal{L}({\bm{z}}^{\prime}_{i}),\nabla_{\theta}\mathcal{L}({\bm{x}}_{j},I_{j})-\nabla_{\theta}\mathcal{L}({\bm{x}}_{j},I_{\text{noise}})\rangle,$ (5)
where $\nabla_{\theta}\mathcal{L}({\bm{x}}_{j},I_{j})$ and $\nabla_{\theta}\mathcal{L}({\bm{x}}_{j},I_{\text{noise}})$ are the gradients computed with the original and Gaussian noise images, respectively. The visual influence matrix $\mathcal{I}_{\texttt{VDS}}$ provides insights into which training samples have the most influence on the validation samples from a visual perspective. This matrix can be used to further rank and select training data that are most impactful for tasks requiring strong visual grounding, ensuring that the selected subset effectively supports vision-dependent performance.
Our empirical results demonstrate that VDS-based data selection has varying effectiveness across different vision-language tasks (Tab. 8). The approach shows substantial improvements on tasks requiring strong visual understanding, such as open-ended generation (LLaVA-W Bench: +2.72%) and multiple-choice understanding (MMBench-EN: +1.90%, MMBench-CN: +0.90%). However, tasks that primarily rely on textual reasoning show decreased performance, including SQA-I (-1.84%) and TextVQA (-1.44%). These results align with and extend the findings in Cambrian [39], demonstrating that the effectiveness of VDS corresponds to a task’s visual dependency - tasks that maintain performance without visual inputs show limited or negative impact from VDS-based selection, while visually-dependent tasks benefit significantly. This pattern suggests that VDS effectively identifies training samples where visual information plays an important role in model learning.
Task w/o VDS w/ VDS Delta (%)
VQAv2 [9] 76.3 75.8 -0.66
GQA [13] 60.7 60.9 +0.33
VizWiz [10] 50.1 50.3 +0.40
SQA-I [25] 70.8 69.5 -1.84
TextVQA [37] 55.6 54.8 -1.44
POPE [21] 87.5 86.8 -0.80
MME [7] 1485.7 1489.3 +0.24
MMBench (en) [47] 63.1 64.3 +1.90
MMBench (cn) [47] 55.8 56.3 +0.90
LLaVA-W Bench [23] 66.1 67.9 +2.72
Table 8: Impact of Visual Dependency Score (VDS) on Selection Performance. Performance comparison between influence calculation with and without VDS. VDS shows strongest improvements on LLaVA-W Bench (+2.72%), MMBench (en) (+1.90%), and MMBench (cn) (+0.90%), while tasks like SQA-I (-1.84%), TextVQA (-1.44%), and POPE (-0.80%) show significant performance drops.
Appendix C Algorithm Details
We provide detailed pseudocode for our two-stage ICONS framework. Stage 1 (specialist) computes task-specific influence scores through gradient-based analysis with efficient random projections. Stage 2 (generalist) implements our voting-based consensus mechanism to select samples that are influential across multiple tasks.
Algorithm 1 ICONS Stage 1: Specialist (Task-specific Influence Computation)
0:  Training dataset $D=\{(x_{i},I_{i},y_{i})\}_{i=1}^{N}$, target tasks $T=\{T_{1},...,T_{K}\}$
0:  Warm-up ratio $r$ (default 5%)
0:  Task-specific influence scores $\{\bar{I}_{k}\}_{k=1}^{K}$
1:  for each task $T_{k}\in T$ do
2:     // Step 1: Warm-up Training
3:     Sample warm-up set $D_{\text{warmup}}\subset D$ of size $r|D|$
4:     $f_{\text{warmup}}\leftarrow\text{LoRA}(f_{\text{base}},D_{\text{warmup}})$
5:     // Step 2: Gradient Computation
6:     for each training data $z_{i}\in D$ do
7:        $g_{i}\leftarrow\nabla_{\theta_{w}}L(f_{\text{warmup}}(z_{i}),y_{i})$
8:        $\tilde{g}_{i}\leftarrow\text{Normalize}(Rg_{i})$Random projection}
9:     end for
10:     for each validation data $z^{\prime}_{j}\in D_{\text{val}}^{k}$ do
11:        $g^{\prime}_{j}\leftarrow\nabla_{\theta_{w}}L(f_{\text{warmup}}(z^{\prime}_{j}),y^{\prime}_{j})$
12:        $\tilde{g}^{\prime}_{j}\leftarrow\text{Normalize}(Rg^{\prime}_{j})$
13:     end for
14:     // Step 3: Influence Matrix Computation
15:     for each $z_{i}\in D$, $z^{\prime}_{j}\in D_{\text{val}}^{k}$ do
16:        $I_{ij}^{k}\leftarrow\langle\tilde{g}_{i},\tilde{g}^{\prime}_{j}\rangle$
17:     end for
18:     // Compute average influence per training sample
19:     $\bar{I}_{k}(z_{i})\leftarrow\frac{1}{|D_{\text{val}}^{k}|}\sum_{j=1}^{|D_{\text{val}}^{k}|}I_{ij}^{k}$
20:  end for
21:  return  Task-specific influence scores $\{\bar{I}_{k}\}_{k=1}^{K}$
Algorithm 2 ICONS Stage 2: Generalist (Influence Consensus-based Data Selection)
0:  Task-specific influence scores $\{\bar{I}_{k}\}_{k=1}^{K}$
0:  Selection ratio $p$, number of tasks $K$
0:  Selected subset $S\subset D$ of size $m\ll N$
1:  // Compute voting thresholds
2:  for each task $T_{k}\in T$ do
3:     $\tau_{k}\leftarrow(1-p)\text{-th percentile of }\{\bar{I}_{k}(z_{i})\}_{i=1}^{N}$
4:  end for
5:  // Voting process
6:  for each training sample $z_{i}\in D$ do
7:     $I_{\text{vote}}(z_{i})\leftarrow 0$
8:     for each task $T_{k}\in T$ do
9:        $\text{vote}_{k}(z_{i})\leftarrow\mathbbm{1}[\bar{I}_{k}(z_{i})\geq\tau_{k}]$
10:        $I_{\text{vote}}(z_{i})\leftarrow I_{\text{vote}}(z_{i})+\text{vote}_{k}(z_{i})$
11:     end for
12:  end for
13:  // Select top samples based on total votes
14:  $S\leftarrow\text{top-}p\text{ samples by }I_{\text{vote}}$
15:  return  Selected subset $S$
Appendix D Future Work
Our work opens several promising research directions for improving vision-language data selection. While our work focuses specifically on visual instruction tuning data, our influence consensus approach can be naturally extended to other stages of MLLM training, such as alignment stage. The majority voting mechanism may under-represent tasks with unique characteristics or those in the long tail, as it prioritizes samples that broadly benefit multiple tasks to build the main knowledge pool. This can lead to limited support for specialized tasks or the reinforcement of spurious correlations spanning multiple tasks. Future work could explore weighted voting mechanisms, in which tasks are assigned weights based on their relative importance or contribution to overall model performance for more balanced data selection. Additionally, investigating more efficient gradient computation and storage methods would help scale these methods to larger datasets while maintaining strong performance across diverse vision-language tasks.
Appendix E Visualizations
We visualize the most influential top three examples across specialists (Figs. 9, 10, 11, 12, 13, 14, 15, 16, 17 and 18) and the generalist selection (Fig. 19), along with samples from their corresponding tasks. Notably, the selected high-influence examples by specialists show strong task-specific characteristics both structurally and contextually - they mirror the key attributes of their target tasks in terms of question structure, reasoning patterns, and required visual-language understanding capabilities. Furthermore, the visualization of top influential examples reveals distinct patterns in what makes training samples valuable for different vision-language tasks. VQAv2, GQA, and SQA-I specialists favor multi-turn Q&A scenarios that test both visual comprehension and contextual understanding, while TextVQA, POPE, and MME specialists emphasize text recognition, object verification, and spatial relationships respectively. MMBench-EN and MMBench-CN show consistent patterns despite language differences, focusing on clear, unambiguous scenes that translate well. The LLaVA-W Bench specialist prioritizes examples requiring detailed explanations and multi-step reasoning, and the answers are generally longer. The generalist model values diverse scenarios that combine multiple skills simultaneously. Common characteristics that make these examples particularly valuable include multi-turn interactions, clear visual elements, factual and inferential reasoning, cross-modal interaction, and the ability to test multiple capabilities within a single example. This suggests that the most effective training samples are those that combine multiple types of reasoning while maintaining clear, unambiguous ground truth that can be consistently learned across tasks.
Figure 9: VQAv2. Top-left: A sample from VQAv2 [9]. Remaining panels show top three influential samples selected using the specialist influence ranking step.
Figure 10: GQA. Top-left: A sample from GQA [13]. Remaining panels show top three influential samples selected using the specialist influence ranking step.
Figure 11: VizWiz. Top-left: A sample from VizWiz [10]. Remaining panels show top three influential samples selected using the specialist influence ranking step.
Figure 12: SQA. Top-left: A sample from SQA [25]. Remaining panels show top three influential samples selected using the specialist influence ranking step.
Figure 13: TextVQA. Top-left: A sample from TextVQA [37]. Remaining panels show top three influential samples selected using the specialist influence ranking step.
Figure 14: Pope. Top-left: A sample from Pope [21]. Remaining panels show top three influential samples selected using the specialist influence ranking step.
Figure 15: MME. Top-left: A sample from MME [7]. Remaining panels show top three influential samples selected using the specialist influence ranking step.
Figure 16: MMBench (en). Top-left: A sample from MMBench (en) [47]. Remaining panels show top three influential samples selected using the specialist influence ranking step.
Figure 17: MMBench (cn). Top-left: A sample from MMBench (cn) [47]. Remaining panels show top three influential samples selected using the specialist influence ranking step.
Figure 18: LLaVA-W Bench. Top-left: A sample from LLaVA-W Bench [23]. Remaining panels show top three influential samples selected using the specialist influence ranking step.
Figure 19: Generalist. We show top three influential samples selected after the generalist step.

### v3

@@
 Abstract
-Visual Instruction Tuning typically requires a large amount of vision-language training data.
-This data often containing redundant information that increases computational costs without proportional performance gains.
-In this work, we introduce ICONS, a gradient-driven Influence CONsensus approach for vision-language data Selection that selects a compact training dataset for efficient multi-task training.
-The key element of our approach is cross-task influence consensus, which uses majority voting across task-specific influence matrices to identify samples that are consistently valuable across multiple tasks, allowing us to effectively prioritize data that optimizes for overall performance.
-Experiments show that models trained on our selected data (20% of LLaVA-665K) achieve 98.6% of the relative performance obtained using the full dataset.
-Additionally, we release this subset, LLaVA-ICONS-133K, a compact yet highly informative subset of LLaVA-665K visual instruction tuning data, preserving high impact training data for efficient vision-language model development.
+Training vision-language models via instruction tuning often relies on large mixtures of data spanning diverse tasks and domains.
+However, these mixtures frequently include redundant information, increasing computational costs without proportional performance gains—necessitating more effective data selection strategies.
+Existing methods typically rely on task-agnostic heuristics to estimate data importance or focus on optimizing single tasks in isolation, limiting their effectiveness in multitask settings.
+In this work, we introduce ICONS, a gradient-based Influence CONsensus approach for vision-language data Selection.
+Our method leverages first-order training dynamics to estimate the influence of individual training examples on validation performance and aggregates these estimates across tasks via majority voting over task-specific influences.
+This cross-task consensus identifies data points that are consistently valuable across tasks, enabling us to prioritize examples that drive overall performance.
+The voting-based design further mitigates issues such as score calibration and outlier sensitivity, resulting in robust and scalable data selection for diverse multitask mixtures.
+With only 20% of the data from LLaVA-665K and Cambrian-7M, our selected subsets retain 98.6% and 98.8% of the performance achieved with full datasets—and can even surpass full-data training at a 60% selection ratio on LLaVA-665K.
+Our approach also generalizes to unseen tasks and architectures, demonstrating strong transfer.
+We release two compact, high-utility subsets—LLaVA-ICONS-133K and Cambrian-ICONS-1.4M —preserving impactful training examples for efficient and scalable vision-language model development.
+1 Introduction
+Visual instruction tuning is a crucial step in training multimodal language models liu2024improved ; llava , enabling them to follow language instructions grounded in visual content.
+Recent approaches rely on large-scale datasets such as LLaVA-665K liu2024improved and Cambrian-7M tong2024cambrian , which contain 665K and 7M examples, respectively.
+While effective, these datasets introduce significant barriers to iteration and deployment: prolonged training times brown2020language ; kaplan2020scaling , high storage demands rajbhandari2020zero ; gadre2024datacomp , and substantial compute costs maini2024rephrasing ; wu2022sustainable .
+Moreover, not all examples contribute equally to all tasks—naively scaling up diverse data mixtures can introduce redundancy and inefficiency.
+This raises a fundamental question:
+{mdframed}
+[leftmargin=0pt, rightmargin=0pt, innerleftmargin=10pt, innerrightmargin=10pt, skipbelow=0pt]
+Can we identify a compact, multitask-effective subset of training data that preserves model capabilities arcoss tasks while enabling faster experimentation?
+Prior work has explored various data selection strategies, including gradient-based approaches xia2024less ; deng2024influential , influence functions yang2022dataset ; koh2017understanding , and diversity-based sampling yu2024diversify ; bukharin2023data .
+However, many of these methods either optimize for single tasks in isolation or maximize source diversity without aligning to downstream needs.
+In multitask visual instruction tuning, this is particularly limiting: optimizing for one task may hurt generalization, and task-agnostic diversity may dilute impact.
+Rather than selecting
 Figure 1: Influence consensus for vision-language data selection.
-(Left) Given a large scale visual instruction tuning dataset (LLaVA-665K [23]), our method computes an influence score of each training example on each target task, aiming to select examples with high influence across multiple tasks.
-Our selected compact 20% subset (LLaVA-ICONS-133K) contains data points that achieve influence consensus across tasks, as indicated by the green check marks showing high influence ranking on corresponding tasks.
-(Right) The radar plot shows the relative performance on various benchmarks comparing training with LLaVA-665K (the full dataset) and LLaVA-ICONS-133K (our selected subset), where the selected subset achieves comparable performance to the full dataset.
-1 Introduction
-Visual instruction tuning has become a critical stage in developing Multimodal LLMs [23, 22], enabling these models to effectively interpret and respond to language instructions based on visual content.
-Current approaches typically rely on large instruction tuning datasets (e.g., LLaVA 1.5 [22] and Cambrian [39], which use 665K and 7M instruction tuning data points, respectively).
-These large datasets, while effective, pose practical challenges for research iteration and model development: extended development cycles due to lengthy training times [3, 15], increased storage and memory demands for managing large-scale datasets [36, 8], and substantial computational resources required for training [27, 40].
-This motivates a fundamental question:
-Can we identify a compact subset of training data that preserves model capabilities while enabling faster experimentation?
-Prior work has explored various approaches to data selection, including gradient-based methods [42, 6], influence functions [43, 17], and diversity sampling [44, 4].
-These methods typically either focus on optimizing for single tasks in isolation, or maximizing source data diversity without considering target task distributions.
-However, selecting data for visual instruction tuning requires careful consideration of performance across diverse downstream tasks.
-Simply selecting the most influential samples for an individual task risks creating a dataset that performs well on that specific benchmark but fails to develop broader capabilities.
-Rather than focusing on task-specific performance, we aim to identify universally valuable training samples by aggregating gradient-based influence measures across diverse tasks using a voting mechanism, effectively selecting data that contributes to broad model capabilities.
-We introduce ICONS (Influence CONsensus vision-language data Selection), a method that builds upon the gradient-based influence estimation approach from LESS [42].
+(Left) Given a large scale visual instruction tuning dataset (LLaVA-665K), our method uses majority voting across task-specific influence scores to identify training samples that are consistently influential across multiple tasks, forming a compact 20% subset (LLaVA-ICONS-133K) with data points achieving influence consensus.
+(Right) The radar plot compares performance between LLaVA-665K and our selected subset, showing the selected subset achieves comparable results to the full dataset.
+data based on per-task influence, we aim to identify samples that are broadly useful—training examples that consistently contribute across tasks.
+To do this, we aggregate gradient-based influence scores using a simple yet effective majority voting scheme.
+We introduce ICONS (Influence CONsensus vision-language data Selection), a method that builds upon the gradient-based selection approach LESS xia2024less .
 Given access to validation data for each target task, our method: (1) computes first-order gradient influence scores to measure how each training sample impacts task-specific performance, and (2) uses influence consensus through majority voting to identify training samples that show consistent positive value across multiple tasks.
 This consensus-based mechanism identifies universally valuable training examples: while some samples might be highly influential for individual tasks, we prioritize those that demonstrate broad utility across the task spectrum.
-While the computational cost of influence estimation is expensive, this front-loaded, one-time investment yields a standardized, compact dataset that can significantly accelerate iteration and development of both multimodal data and models.
-Using ICONS, we create LLaVA-ICONS-133K, an automatically curated 20% subset of the LLaVA dataset [23].
-This compact dataset maintains 98.6% of the original performance across multiple vision-language tasks, providing a 2.8 percentage point improvement over randomly selecting the same-sized subset (95.8%) and eliminating two-thirds of the performance drop from shrinking the training data.
-Moreover, our ICONS outperforms all baselines across different selection ratios, and remarkably achieves above-full-dataset performance, surpassing the original dataset at a 60% selection ratio.
-Importantly, LLaVA-ICONS-133K shows strong transferability, maintaining 95.5-113.9% relative performance across diverse unseen tasks such as chart understanding and infographic comprehension, suggesting that ICONS identifies fundamentally valuable training examples rather than task-specific patterns.
+While the computational cost of influence estimation is expensive, this front-loaded, one-time investment yields a standardized, compact dataset that can significantly accelerate development of multimodal models, and enables reusable gradient datastores that amortize costs across iterations and deliver long-term savings.
+Using ICONS, we create LLaVA-ICONS-133K and Cambrian-ICONS-1.4M, automatically curated 20% subsets of the LLaVA-665K dataset liu2024improved and Cambrian-7M tong2024cambrian dataset, respectively.
+These compact datasets maintain 98.6% and 98.8% of their original performance across multiple vision-language tasks, providing significant improvements over randomly selecting same-sized subsets (95.8% and 95.4%) and eliminating approximately two-thirds of the performance drop from shrinking the training data.
+Moreover, our ICONS outperforms all baselines across different selection ratios, and remarkably achieves above-full-dataset performance, surpassing the original datasets at a 60% selection ratio for LLaVA-665K.
+Importantly, the selected subset shows strong transferability, e.g., LLaVA-ICONS-133K maintains 95.5-113.9% relative performance across unseen tasks, suggesting that ICONS identifies fundamentally valuable training data.
 We summarize our key contributions:
   1. 1.
-    We propose ICONS, a simple yet effective approach for multi-task vision-language data selection that identifies broadly valuable training samples through majority voting across task-specific gradient influence scores.
+    We propose ICONS, a simple yet effective method for multitask vision-language data selection that identifies broadly valuable training examples via majority voting over task-specific gradient influence scores.
   2. 2.
-    Our voting-based aggregation consistently outperforms all baselines (§3.2) and, most notably, exceeds 102% of the full dataset performance at 60% selection ratio (§3.3).
-We further analyze different influence aggregation approaches and show the effectiveness of our consensus-aware selection over direct-merge alternatives (§3.5).
+    Our consensus-based selection outperforms all baselines (§3.2) and we ablate influence aggregation strategies and show the advantage of voting-based consensus (§3.3).
+We further show that ICONS exceeds 102% of full-dataset performance at a 60% selection ratio on LLaVA-665K3.5).
   3. 3.
-    We release LLaVA-ICONS-133K, a compact 20% subset of LLaVA-665K, achieving near-full performance (98.6%), transferring well to unseen tasks (§3.4), and serving as a standardized training set for resource-efficient development of multimodal models.
+    We release LLaVA-ICONS-133K and Cambrian-ICONS-1.4M, compact 20% subsets of LLaVA-665K and Cambrian-7M respectively, achieving near-full performance (98.6% and 98.8%), transferring well to unseen tasks (§3.4), and serving as standardized training sets for resource-efficient development.
 2 Influence consensus for vision-language data selection
-We propose a consensus-driven, gradient-based data selection framework for visual instruction tuning datasets.
+We propose a consensus-driven, gradient-based data selection framework (Fig.
+2) for visual instruction tuning datasets.
 We formalize the problem setup in §2.1 and establish gradient-based influence estimation preliminaries in §2.2.
-Our two-stage data selection framework consists of: first, the specialist stage (§2.3), which computes task-specific influence scores, followed by the generalist stage (§2.4), which builds cross-task consensus through voting-based aggregation.
+Our two-stage data selection framework consists of: the specialist stage (§2.3), which computes task-specific influence scores, and the generalist stage (§2.4), which builds cross-task consensus through voting-based aggregation.
 Figure 2: ICONS.
@@
 2.1 Problem formulation
-Given a large-scale visual instruction tuning dataset $\mathcal{D}=\{({\bm{x}}_{i},{\bm{I}}_{i},{\bm{y}}_{i})\}_{i=1}^{N}$ containing $N$ samples, where each data point ${\bm{z}}_{i}=({\bm{x}}_{i},{\bm{I}}_{i},{\bm{y}}_{i})$ includes natural language instruction ${\bm{x}}_{i}$ and an image ${\bm{I}}_{i}$, with corresponding target response ${\bm{y}}_{i}$ [^1] The framework supports multi-turn conversational data, yet we formalize the problem setup for single-turn instruction-tuning for clarity and simplicity., and given access to validation data $\mathcal{V}_{k}$ for each downstream task $T_{k}\in\mathcal{T}=\{T_{1},...,T_{K}\}$, our goal is to select a compact subset $\mathcal{S}\subset\mathcal{D}$ of size $M\ll N$ that maximizes model performance across multiple downstream tasks:
-$\mathcal{S}^{*}=\operatorname*{arg\,max}_{\mathcal{S}\subset\mathcal{D},|\mathcal{S}|=m}\sum_{k=1}^{K}\text{Rel}(f_{\mathcal{S}},T_{k}),$ (1)
-$\text{Rel}(f_{\mathcal{S}},T_{k})=\frac{\text{Score}(f_{\mathcal{S}},T_{k})}{\text{Score}(f_{\mathcal{D}},T_{k})},$ (2)
+Given a large-scale visual instruction tuning dataset $\mathcal{D}=\{({\bm{I}}_{i},{\bm{x}}_{i},{\bm{y}}_{i})\}_{i=1}^{N}$ containing $N$ samples, where each data point ${\bm{z}}_{i}=({\bm{I}}_{i},{\bm{x}}_{i},{\bm{y}}_{i})$ includes an image ${\bm{I}}_{i}$, natural language instruction ${\bm{x}}_{i}$, and corresponding target response ${\bm{y}}_{i}$ [^1] The framework supports multi-turn conversational data, yet we formalize the problem setup for single-turn instruction-tuning for clarity and simplicity., and given access to validation data $\mathcal{V}_{k}$ for each downstream task $T_{k}\in\mathcal{T}=\{T_{1},...,T_{K}\}$, our goal is to select a compact subset $\mathcal{S}\subset\mathcal{D}$ of size $M\ll N$ that maximizes model performance across multiple downstream tasks:
+$\mathcal{S}^{*}=\operatorname*{arg\,max}_{\mathcal{S}\subset\mathcal{D},|\mathcal{S}|=M}\sum_{k=1}^{K}\text{Rel}(f_{\mathcal{S}},T_{k}),\quad\text{Rel}(f_{\mathcal{S}},T_{k})=\frac{\text{Score}(f_{\mathcal{S}},T_{k})}{\text{Score}(f_{\mathcal{D}},T_{k})},$ (1)
 where $f_{\mathcal{S}}$ and $f_{\mathcal{D}}$ denote models trained on subset $\mathcal{S}$ and full dataset $\mathcal{D}$, respectively.
@@
 Rel.
-quantifies the subset-trained model’s performance relative to that of the model trained on the entire dataset, with values close to 1 indicating that the subset maintains the performance of full training [18].
-Our objective is to find a subset where Rel.
-$\approx 1$, i.e., the model trained on the selected subset achieves performance comparable to using the full dataset across all tasks.
+quantifies the subset-trained model’s performance relative to that of the model trained on the entire dataset, with values close to 1 indicating that the subset maintains the performance of full training coincide .
+Our objective is to select a subset where Rel.
+$\approx 1$, i.e., the model trained on the subset achieves comparable performance to the one trained with full dataset.
 2.2 Preliminaries
-Building on our problem formulation in §2.1, we formalize how to estimate the influence of training samples on downstream task performance.
-Specifically, since our goal is to maximize $\text{Rel}(f_{\mathcal{S}},T_{k})$ across tasks as defined in Eqn.
-2, we need an efficient way to estimate how each training sample contributes to the $\text{Score}(f_{\mathcal{S}},T_{k})$ term in the numerator.
-Following [34, 42], we estimate how individual training samples $z_{i}\in\mathcal{D}$ affect validation loss by measuring their gradient alignment with reducing validation loss on $\mathcal{V}_{k}$, which directly impacts task-specific evaluation scores.
-Denote a training data point as ${\bm{z}}$ and a validation data point as ${\bm{z}}^{\prime}$ from validation set $\mathcal{V}_{k}$ for task $T_{k}$ When training with SGD and batch size 1, using data point ${\bm{z}}$ at timestep $t$ leads to a model update ${\bm{\theta}}_{t+1}={\bm{\theta}}_{t}-\eta_{t}\nabla\ell({\bm{z}};{\bm{\theta}}_{t})$, where $\eta_{t}$ is the learning rate.
+Building on the problem formulation in §2.1, we formalize how to estimate the influence of training samples on downstream task performance.
+Since our goal is to maximize $\text{Rel}(f_{\mathcal{S}},T_{k})$ across tasks as defined in Eqn.
+1, we need an efficient way to estimate how each training sample contributes to the $\text{Score}(f_{\mathcal{S}},T_{k})$ term in the numerator.
+Denote a training data point as ${\bm{z}}$ and a validation data point as ${\bm{z}}^{\prime}$ from validation set $\mathcal{V}_{k}$ for task $T_{k}$.
+Following pruthi2020estimating ; xia2024less , we estimate how ${\bm{z}}\in\mathcal{D}$ affects validation loss by measuring its gradient alignment with reducing validation loss on $\mathcal{V}_{k}$, which directly impacts task-specific evaluation.
+When training with SGD and batch size 1, using data point ${\bm{z}}$ at timestep $t$ leads to a model update ${\bm{\theta}}_{t+1}={\bm{\theta}}_{t}-\eta_{t}\nabla\ell({\bm{z}};{\bm{\theta}}_{t})$, where $\eta_{t}$ is the learning rate.
 To reduce the computational cost, we use the first-order Taylor expansion to estimate the loss on a given validation data point ${\bm{z}}^{\prime}$ at time step $t+1$ as:
@@
 The influence of a training data point ${\bm{z}}$ on a validation data point ${\bm{z}}^{\prime}$ is:
-$\mathcal{I}_{t}({\bm{z}}\rightarrow{\bm{z}}^{\prime})$ $=\ell({\bm{z}}^{\prime};{\bm{\theta}}_{t+1})-\ell({\bm{z}}^{\prime};{\bm{\theta}}_{t})$
-$\approx-\eta_{t}\langle\nabla\ell({\bm{z}}^{\prime};{\bm{\theta}}_{t}),\nabla\ell({\bm{z}};{\bm{\theta}}_{t})\rangle,$
-which we refer to as an influence score. For our case of visual instruction tuning [23] which uses single-epoch training, the influence score is:
-$\mathcal{I}_{\text{single}}({\bm{z}}\rightarrow{\bm{z}}^{\prime})=\eta\langle\nabla\ell({\bm{z}}^{\prime};{\bm{\theta}}_{t}),\nabla\ell({\bm{z}};{\bm{\theta}}_{t})\rangle,$
-where $\eta$ is the learning rate used during training. This can be generalized to multi-epoch training by accumulating the influence scores over all $n$ epochs:
-$\mathcal{I}_{\text{total}}({\bm{z}}\rightarrow{\bm{z}}^{\prime})=\sum_{i=1}^{n}\eta_{i}\langle\nabla\ell({\bm{z}}^{\prime};{\bm{\theta}}_{i}),\nabla\ell({\bm{z}};{\bm{\theta}}_{i})\rangle,$
-where $\eta_{i}$ is the learning rate used during the $i$th epoch.
-The gradient-based selection approach selects training samples $\{{\bm{z}}\}$ that maximize the gradient inner product $\langle\nabla\ell({\bm{z}}^{\prime};{\bm{\theta}}_{t}),\nabla\ell({\bm{z}};{\bm{\theta}}_{t})\rangle$ [^2] In practice, we use cosine similarity instead of direct inner products to avoid biasing selection toward shorter sequences, since gradient norms tend to be inversely correlated with sequence length as noted in [42].
+$\mathcal{I}_{t}({\bm{z}}\rightarrow{\bm{z}}^{\prime})$ $=\ell({\bm{z}}^{\prime};{\bm{\theta}}_{t+1})-\ell({\bm{z}}^{\prime};{\bm{\theta}}_{t})\approx-\eta_{t}\langle\nabla\ell({\bm{z}}^{\prime};{\bm{\theta}}_{t}),\nabla\ell({\bm{z}};{\bm{\theta}}_{t})\rangle,$
+which we refer to as an influence score.
+The gradient-based selection approach selects training samples $\{{\bm{z}}\}$ that maximize the gradient inner product $\langle\nabla\ell({\bm{z}}^{\prime};{\bm{\theta}}_{t}),\nabla\ell({\bm{z}};{\bm{\theta}}_{t})\rangle$ [^2] In practice, we use cosine similarity instead of direct inner products to avoid biasing selection toward shorter sequences, since gradient norms tend to be inversely correlated with sequence length as noted in xia2024less .
 through a greedy, first-order approximation, which leads to larger reductions in validation loss for point ${\bm{z}}^{\prime}$.
-While this approach omits second-order terms compared to influence functions [17], it provides an efficient approximation for ranking the impact of training samples [11, 42].
+While it omits second-order terms compared to influence functions koh2017understanding , it provides an efficient approximation to rank the impact of training samples xia2024less .
 2.3 Specialist: individual task influence ranking
 To rank the influence of training data for each target task, we compute the influence score of each training data point on a validation set that represents the target task distribution.
-Following LESS [42], the process involves three steps: (1) training the model on 5% randomly selected data as a lightweight warm-up phase to initialize visual instruction-following capabilities, (2) computing gradients for training and validation data and compressing the gradients via random projection, and (3) computing the influence score to quantify the impact of each training data on validation set.
+Following LESS xia2024less , the process involves three steps: (1) training the model on 5% randomly selected data as a lightweight warm-up to initialize visual instruction-following capabilities, (2) computing gradients for training and validation data and compressing the gradients via random projection, and (3) computing the influence score to quantify the impact of each training data on validation set.
 Step 1: Warm-up Training.
-Following LESS [42], we first perform a brief warmup phase using Low-Rank Adaptation (LoRA) [12] on a small random subset $\mathcal{D}_{\mathrm{warmup}}\subset\mathcal{D}$ (5% of the training data): $f_{\text{warmup}}=\text{LoRA}(f_{\text{base}},\mathcal{D}_{\text{warmup}})$.
-This warmup phase allows the model to develop basic instruction-following capabilities.
+Following LESS xia2024less , we first apply LoRA hu2021lora on a small random subset $\mathcal{D}_{\mathrm{warmup}}\subset\mathcal{D}$ (5%) to obtain $f{\text{warmup}}=\text{LoRA}(f_{\text{base}},\mathcal{D}_{\text{warmup}})$.
+This allows the model to develop basic visual instruction-following capabilities.
 Step 2: Gradient computation.
@@
 where $y_{i}$ and $y^{\prime}_{j}$ are the targets for $z_{i}$ and $z^{\prime}_{j}$, respectively.
-In order to reduce computational and storage overhead, we apply random projection to the gradient feature: $\tilde{g}_{i}=Rg_{i}$ and $\tilde{g}^{\prime}_{j}=R^{\prime}g^{\prime}_{j}$, where $R,R^{\prime}\in\mathbb{R}^{d^{\prime}\times d}$ is a random projection matrix with $d^{\prime}\ll d$ that preserves inner products with high probability [14].
-We further normalize the projected gradient features, $\tilde{g}_{i}=\frac{\tilde{g}_{i}}{\|\tilde{g}_{i}\|_{2}},\tilde{g}^{\prime}_{j}=\frac{\tilde{g}^{\prime}_{j}}{\|\tilde{g}^{\prime}_{j}\|_{2}}$ to prevent bias from sequence length differences [42].
+In order to reduce computational and storage overhead, we apply random projection to the gradient feature: $\tilde{g}_{i}=Rg_{i}$ and $\tilde{g}^{\prime}_{j}=Rg^{\prime}_{j}$, where $R\in\mathbb{R}^{d^{\prime}\times d}$ is a random projection matrix with $d^{\prime}\ll d$ that preserves inner products with high probability (johnson1984extensions, ).
+We further normalize the projected gradients, $\tilde{g}_{i}=\frac{\tilde{g}_{i}}{\|\tilde{g}_{i}\|_{2}},\tilde{g}^{\prime}_{j}=\frac{\tilde{g}^{\prime}_{j}}{\|\tilde{g}^{\prime}_{j}\|_{2}}$ to prevent bias from sequence length differences xia2024less .
 Step 3: Influence matrix computation.
-We compute the influence matrix $I\in\mathbb{R}^{|\mathcal{D}|\times|\mathcal{D}_{\text{val}}^{k}|}$ where each entry $I_{ij}=\langle\tilde{g}_{i},\tilde{g}^{\prime}_{j}\rangle$ is the influence of training data ${\bm{z}}_{i}$ on validation data ${\bm{z}}^{\prime}_{j}$, and the average influence of training data ${\bm{z}}_{i}$ on the target task $k$ is then computed as:
+We compute the influence matrix $I\in\mathbb{R}^{|\mathcal{D}|\times|\mathcal{D}_{\text{val}}^{k}|}$ where each entry $I_{ij}=\langle\tilde{g}_{i},\tilde{g}^{\prime}_{j}\rangle$ is the influence of the training data ${\bm{z}}_{i}$ on the validation data ${\bm{z}}^{\prime}_{j}$, and then the average influence of training data ${\bm{z}}_{i}$ on the target task $k$ is calculated as:
 $\bar{I}_{k}({\bm{z}}_{i})=\frac{1}{|\mathcal{D}_{\text{val}}^{k}|}\sum_{j=1}^{|\mathcal{D}_{\text{val}}^{k}|}I_{ij}.$ (2)
-This influence estimation process provides a task-specific ranking for the training set $\mathcal{D}$ with respect to task $\mathcal{T}_{k}$, where a higher influence score $\bar{I}_{i}$ suggests a higher influence for the target task.
-However, recall that our ultimate goal is to select a compact subset $\mathcal{S}\subset\mathcal{D}$ of size $M\ll N$ that maximizes $\sum_{k=1}^{K}\text{Rel}(f_{\mathcal{S}},T_{k})/K$ across all tasks.
-We address this disconnection between task-specific rankings and overall dataset optimization by proposing a voting-based generalist approach; this approach aggregates influence scores for each target task to identify the most broadly impactful training data.
+This influence estimation process provides a task-specific ranking for the training set $\mathcal{D}$ with respect to task $\mathcal{T}_{k}$, where a higher influence score $\bar{I}_{i}$ suggests a higher influence $\mathcal{T}_{k}$.
+We can select a small training subset $\mathcal{S}_{k}$ for a given task $k$ by selecting the training examples ${\bm{z}}_{i}$ with the highest-influence values $\bar{I}_{k}({\bm{z}}_{i})$.
+This simple greedy approach has been shown by LESS to be successful, and thus we use it as our task-specific (“specialist”) baseline.
+However, recall that our goal is to select a single compact subset that maximizes the performance across all tasks.
+We address this disconnection between task-specific rankings and overall optimization by proposing a voting-based generalist approach to identify the most broadly impactful training data.
 2.4 Generalist: cross-task influence consensus
-Our consensus-based voting strategy identifies training samples that consistently show a high influence score across various tasks.
-For each task $k$, we set a task-specific threshold $\tau_{k}$ at the $((1-p)\times 100)$-th percentile of the influence score distribution within that task, where $p$ represents the proportion of top samples to consider ($p=0.2$ in our main experiments).
-A sample ${\bm{z}}_{i}$ is considered important for task $k$ if its influence score $I_{k}({\bm{z}}_{i})\geq\tau_{k}$.
-Formally, we define the voting-based influence score for sample $z_{i}$ on task $k$ as:
-$I_{\text{vote}}({\bm{z}}_{i})=\sum_{k=1}^{K}\mathbbm{1}[I_{k}({\bm{z}}_{i})\geq\tau_{k}],$ (4)
-where $\mathbbm{1}[\cdot]$ denotes the indicator function, and $I_{k}({\bm{z}}_{i})$ represents the influence score of sample ${\bm{z}}_{i}$ for task $k$.
-The final score for each sample is the total number of votes it receives across all tasks, ranging from 0 (indicating no importance for any task) to $K$ (indicating consistent importance for all tasks).
-Influence Consensus identifies a unique subset of samples that are consistently influential across multiple tasks, rather than those that might be highly influential in only a few tasks.
-After computing the aggregated influence scores, we rank the training samples and select the top $p\%$ as final subset for model training.
-While directly combining validation data from all target tasks and computing a single influence matrix (Direct-merge Selection, §3.5) might seem intuitive, we adopt this consensus-based approach for three reasons.
-First, specialist selection provides a practical performance upper bound, as the performance achieved by selecting the top $p$of data specifically for a single task represents the theoretical maximum achievable with that data budget, serving as a critical baseline for evaluating generalist selection trade-offs (§3.5).
-Second, this approach is scalable and naturally supports continual learning scenarios, as influence matrices for new tasks or benchmarks can be computed independently without recomputing training gradients, thereby reducing computational costs.
-Third, computing influence matrices independently for different tasks give us better understanding of cross-task influence dynamics.
-Our analysis (§3.3) reveals that the influence ranking of training data vary significantly across different tasks, with highly influential samples for one task often showing limited impact on others.
-This further provides insights to design more effective consensus mechanisms for generalist selection while maintaining computational tractability.
+Our goal is to identify a training set subset $\mathcal{S}\subset\mathcal{D}$ of size $M\ll N$ such that its performance across all tasks remains high as defined by Eqn.
+1.
+There are multiple ways to tackle this, depending on the assumptions one makes about the task-specific influence scores $\bar{I}_{k}({\bm{z}}_{i})$.
+The simplest approach is to merge together all the different tasks’ validation sets $\mathcal{D}_{\text{val}}^{k}$ (normalizing for their different sizes) and compute the total influence score for a training example ${\bm{z}}_{i}$ as:
+$I_{\text{Merge}}({\bm{z}}_{i})=\sum_{k=1}^{K}\bar{I}_{k}({\bm{z}}_{i}).$ (3)
+A similar aggregation approach is the one suggested in LESS xia2024less :
+$I_{\text{Max}}({\bm{z}}_{i})=\max_{k=1,\dots K}\bar{I}_{k}({\bm{z}}_{i}),$ (4)
+i.e., the influence of the data is measured as its highest influence on any tasks.
+The set of $M$ training examples with the highest aggregated influence scores would be selected for inclusion in the training set $\mathcal{S}_{\text{merge}}$ (correspondingly, $\mathcal{S}_{\text{max}}$).
+Both approaches, however, require that the influence scores $\bar{I}_{k}$ be well-calibrated across the different tasks; as we show in §3.3 this may not necessarily be the case.
+An alternative approach which does not require directly comparing influence scores $\bar{I}_{k}$ across tasks $k$ is to leverage the relative rank of the training examples within each task.
+Concretely, we compute $\text{rank}_{k}({\bm{z}}_{i})$ for each example ${\bm{z}}_{i}$ relative to other examples for task $k$ according to their influence scores (higher influence scores correspond to lower rank).
+We can have a couple of options.
+First, we can select the training subset either using the Round Robin (RR) approach ivison2025large where we iterate over tasks and select the lowest-rank example which has not yet been selected to add to our set $\mathcal{S}_{\text{RR}}$.
+Alternatively, we can select the training subset $\mathcal{S}_{\text{MinRank}}$ such that all the examples within it have a low rank for some task $k$.
+Mathematically, albeit somewhat confusingly, this corresponds to:
+$\mathcal{S}_{\text{MinRank}}=\operatorname*{arg\,max}_{\mathcal{S}\subset\mathcal{D},|\mathcal{S}|=M}\min_{\begin{subarray}{c}\text{task }k\\ \text{example }{\bm{z}}_{i}\notin S\end{subarray}}\text{rank}_{k}({\bm{z}}_{i}),$ (5)
+i.e., all examples that are not included in $\mathcal{S}$ would have high relative ranks $\text{rank}_{k}({\bm{z}}_{i})$ for all tasks $k$.
+However, this approach does not consider the potential interplay between tasks.
+Recall that in Eqn.
+1 we aim to maximize sum of the relative performance across all tasks $k$; thus, if a training example is beneficial for multiple tasks, we may want to include it even at the expense of a lower-ranked example for a different task $k$.
+Thus, we introduce a simple consensus-based voting strategy that identifies training samples that consistently show a high influence score across various tasks.
+Concretely, we leverage the specialist training sets $\mathcal{S}_{k}$ as defined in §2.3 consisting of the $M$ highest-influence training examples for each task.
+We then select a combined training set as follows:
+$S_{\text{ICONS}}=\operatorname*{arg\,max}_{\mathcal{S}\subset\mathcal{D},|\mathcal{S}|=M}\sum_{k=1}^{K}\sum_{{\bm{z}}_{i}\in S}\mathbbm{1}[{\bm{z}}_{i}\in S_{k}]$ (6)
+This simple approach offers a key advantage: it does not rely on calibration of influence scores across tasks, and does not make any a-priori assumptions about the relationship between tasks (e.g., that every task needs to have its highest-scoring examples included in the combined training set).
+Our generalist stage converts each task’s ranked list into a binary vote (“above threshold” or not) and then combines these votes across tasks, eliminating the need for task-specific normalization.
+As a result, the selection remains insensitive to scale differences while still capturing relative importance within each task.
+Meanwhile, a training sample is selected only when several tasks independently rank it as influential, preventing over-representation of single-task outliers and ensuring the cross-task utility.
+3 Experiments
+In this section, we first discuss our experiment setup and evaluation benchmarks (§3.1).
+We then present our main results by comparing ICONS with the state-of-the-art methods (§3.2), followed by analysis of different selection strategies (§3.3).
+We further evaluate the transferability of ICONS3.4).
+Lastly, we provide analyze performance trends under different selection ratios (§3.5).
+3.1 Evaluation test-bed
+Datasets & model.
+We apply ICONS on major visual instruction tuning (VIT) training datasets: LLaVA-665K liu2024improved , Cambrian-7M tong2024cambrian and Vision-Flan-186K visionflan .
+The majority of our analysis and ablation experiments are conducted on LLaVA-665K.
+For our experiments, we use the LLaVA-v1.5 model (liu2024improved, ) checkpoint after Stage 1 (pre-training for feature alignment) as defined in the original LLaVA training pipeline, with a default size of 7B parameters and LLaVA-665K unless otherwise specified.
+This checkpoint333 llava-v1.5-mlp2x-336px-pretrain-vicuna-7b-v1.5, which has no prior exposure to the visual instruction tuning data.
+corresponds to the model after training the projector but before any visual instruction tuning in Stage 2.
+Importantly, this model has not been exposed to the LLaVA-665K VIT dataset prior to the data selection process.
+In all experiments, we train the models for one epoch following the official finetuning hyperparameters using LoRA.
+More details on computation, including hardware specifications and runtime are in Appendix A.
+Target tasks.
+We evaluate ICONS across diverse multimodal benchmarks (Appendix C, Tab.
+5) that test different capabilities of vision-language models: 1) Multiple-choice understanding: MMBench (mmbench, ) and MME (mme, )  [^4] For MME, we focus on its perception section following coincide , which evaluates vision capablities.
+2) Visual question answering: VQAv2 (vqav2, ), GQA (gqa, ), and VizWiz (vizwiz, ); 3) Text understanding in images: TextVQA (textvqa, ); 4) Scientific reasoning: ScienceQA (sqa, ); 5) Open-ended generation: LLaVA-W Bench (llava, ); 6) Factual consistency: POPE (pope, ).
+Baselines.
+We compare our ICONS against several baselines, including random selection, CLIP-Score clip for measuring image-text alignment, EL2N el2n based on embedding L2 norms, and Perplexity perplexity using language model scores.
+We also compare against SemDeDup semdedup for semantic deduplication and D2-Pruning (d2, ) for distribution-aware pruning.
+Additional baselines include Self-Sup (selfsup, ) leveraging self-supervised signals, while Self-Filter (self-filter, ) and COINCIDE (coincide, ) are designed for vision-language data selection.
+We reference LLaVA-665K baseline results from COINCIDE coincide .
+Additionally, we compare with representation-based data selection baseline (RDS) ivison2025large ; xia2024less .
 Method VQAv2 GQA VizWiz SQA-I TextVQA POPE MME MMBench LLaVA-W Rel. (%)
@@
 Random 75.7 58.9 44.3 68.5 55.3 84.7 1483.0 62.2 54.8 65.0 95.8
-CLIP-Score [35] 73.4 51.4 43.0 65.0 54.7 85.3 1331.6 55.2 52.0 66.2 91.2
-EL2N [33] 76.2 58.7 43.7 65.5 53.0 84.3 1439.5 53.2 47.4 64.9 92.0
-Perplexity [28] 75.8 57.0 47.8 65.1 52.8 82.6 1341.4 52.0 45.8 68.3 91.6
-SemDeDup [1] 74.2 54.5 46.9 65.8 55.5 84.7 1376.9 52.2 48.5 70.0 92.6
-D2-Pruning [26] 73.0 58.4 41.9 69.3 51.8 85.7 1391.2 65.7 57.6 63.9 94.8
-Self-Sup [38] 74.9 59.5 46.0 67.8 49.3 83.5 1335.9 61.4 53.8 63.3 93.4
-Self-Filter [5] 73.7 58.3 53.2 61.4 52.9 83.8 1306.2 48.8 45.3 64.9 90.9
-COINCIDE [18] 76.5 59.8 46.8 69.2 55.6 86.1 1495.6 63.1 54.5 67.3 97.4
+CLIP-Score clip 73.4 51.4 43.0 65.0 54.7 85.3 1331.6 55.2 52.0 66.2 91.2
+EL2N el2n 76.2 58.7 43.7 65.5 53.0 84.3 1439.5 53.2 47.4 64.9 92.0
+Perplexity perplexity 75.8 57.0 47.8 65.1 52.8 82.6 1341.4 52.0 45.8 68.3 91.6
+SemDeDup semdedup 74.2 54.5 46.9 65.8 55.5 84.7 1376.9 52.2 48.5 70.0 92.6
+D2-Pruning d2 73.0 58.4 41.9 69.3 51.8 85.7 1391.2 65.7 57.6 63.9 94.8
+Self-Sup selfsup 74.9 59.5 46.0 67.8 49.3 83.5 1335.9 61.4 53.8 63.3 93.4
+Self-Filter self-filter 73.7 58.3 53.2 61.4 52.9 83.8 1306.2 48.8 45.3 64.9 90.9
+COINCIDE coincide 76.5 59.8 46.8 69.2 55.6 86.1 1495.6 63.1 54.5 67.3 97.4
+RDS ivison2025large ; xia2024less 75.1 57.9 48.6 68.0 54.9 86.3 1393.8 61.2 52.7 63.7 95.2
 ICONS (ours) 76.3 60.7 50.1 70.8 55.6 87.5 1485.7 63.1 55.8 66.1 98.6
-Table 1: Baseline Comparisons.
+Table 1: Selection results on LLaVA-665K.
 Performance comparison of different data selection approaches when trained on 20% of the LLaVA-665K dataset.
-Models are evaluated across diverse multimodal benchmarks, with tasks ranging from visual question answering (VQAv2, GQA) to multilingual understanding (MMBench).
 The best and second best results for each benchmark are shown in bold and underlined, respectively.
-Our method ICONS achieves the highest overall relative performance (98.6%), consistently outperforming existing approaches including COINCIDE (97.4%) and D2-Pruning (94.8%), while methods like EL2N, Perplexity, and CLIP-Score show limited effectiveness with relative performance around 91-92%.
-3 Experiments
-In this section, we first discuss our experiment setup and evaluation benchmarks (§3.1).
-Then we compare it with the latest state-of-the-art methods (§3.2).
-Next, we provide detailed analysis of our method’s behavior and effectiveness across different dimensions (§3.3).
-We further evaluate the transferability of our selection approach across both unseen tasks and model architectures (§3.4).
-Finally, we conduct ablation studies to understand the contribution (§3.5).
-3.1 Evaluation Test-Bed
-Dataset & Model.
-We evaluate ICONS on visual instruction tuning dataset LLaVA-665K [22].
-We use LLaVA-v1.5-7b-lora model [22] with a default size of 7B parameters unless otherwise specified.
-In all experiments, we train the models for one epoch following the official finetuning hyperparameters, which reduces computational overhead while maintaining selection quality.
-Target Tasks.
-We evaluate the effectiveness of ICONS with a wide range of multimodal benchmarks (Tab.
-2) that test different capabilities of vision-language models.
-The benchmarks include: 1) Multiple-choice understanding: MMBench [47] and MME [7], which assess the model’s ability to select correct answers from given options; 2) Visual question answering: VQAv2 [9], GQA [13], and VizWiz [10], which test basic visual reasoning; 3) Text understanding in images: TextVQA [37], which evaluates OCR capabilities; 4) Scientific reasoning: ScienceQA [25], which tests knowledge-grounded question answering; 5) Open-ended generation: LLaVA-W Bench [23], which evaluate free-form response generation; 6) Factual consistency: POPE [21], which measures hallucination.
-MME POPE SQA-I MMBench
-en cn
-$|\mathcal{D}_{\text{val}}|$ 986 500 424 1,164 1,164
-$|\mathcal{D}_{\text{test}}|$ 2,374 8,910 4,241 1,784 1,784
-Task Type Y/N Y/N MCQ MCQ MCQ
-VQAv2 GQA VizWiz TextVQA LLaVA-W
-$|\mathcal{D}_{\text{val}}|$ 1,000 398 8,000 84 84
-$|\mathcal{D}_{\text{test}}|$ 36,807 12,578 8,000 5,000 84
-Task Type VQA VQA VQA VQA VQA
-Table 2: Statistics of Target Tasks.
-Our target tasks include diverse benchmarks and answer formats, covering different vision-language capabilities.
-Task types include Multiple-Choice Questions (MCQ), Visual Question Answering (VQA), and Yes/No Questions (Y/N).
-3.2 Comparisons with Baselines
-We compare our ICONS against several baseline approaches, including random selection.
-We evaluate CLIP-Score [35] for measuring image-text alignment, EL2N [33] based on embedding L2 norms, and Perplexity [28] using language model scores.
-We also compare against SemDeDup [1] for semantic deduplication and D2-Pruning [26] for distribution-aware pruning.
-Additional baselines include Self-Sup [38] leveraging self-supervised signals, while Self-Filter [5] and COINCIDE [18] are the approaches specifically designed for vision-language data selection.
-All methods are evaluated using a 20% sampling ratio of the original dataset, with results presented in Tab.
-1.
-We reference the baseline results from COINCIDE [18].
-We compare our ICONS against existing data selection approaches in Tab.
-1.
-Our method achieves the best overall performance with 98.6% Rel., outperforming all baselines with LLaVA-ICONS-133K, 20% of the training data.
-Remarkably, we achieve comparable or better performance than training with the entire LLaVA-665K on several tasks: SQA-I (70.8 vs.
+Our method ICONS achieves the highest overall Rel.
+(98.6%), consistently outperforming existing approaches including COINCIDE coincide (97.4%) and D2-Pruning d2 (94.8%).
+3.2 Main results
+LLaVA-665K selection.
+As shown in Tab.
+1, ICONS achieves the best overall performance with 98.6% Rel.
+on LLaVA-665K, outperforming all baselines with LLaVA-ICONS-133K, 20% of the training data.
+Remarkably, we achieve comparable or better performance than full dataset training on several tasks: SQA-I (70.8 vs.
 68.4), MME (1485.7 vs.
-1476.9), VizWiz (50.1 vs.
-47.8) and POPE (87.5 vs.
+1476.9) and POPE (87.5 vs.
 86.4).
-Among the baselines, COINCIDE achieves strong performance (97.4% Rel.), though it falls short of ICONS on key tasks.
-Other methods like EL2N, Perplexity, SemDeDup are less effective, with relative scores around 91-92%, showing limitations in preserving performance.
-Task Full Specialist Generalist Delta (%)
-VQAv2 [9] 79.1 77.1 76.3 1.04
-GQA [13] 63.0 61.1 60.7 0.65
-VizWiz [10] 47.8 53.1 50.1 5.65
-SQA-I [25] 68.4 69.8 70.8 -1.43
-TextVQA [37] 58.2 55.7 55.6 0.18
-POPE [21] 86.4 86.6 87.5 -1.04
-MME [7] 1476.9 1506.1 1485.7 1.35
-MMBench (en) [47] 66.1 66.0 63.1 4.34
-MMBench (cn) [47] 58.9 56.4 55.8 1.06
-LLaVA-W Bench [23] 67.9 67.1 66.1 1.49
-Table 3: Single-task Selection (Specialist) vs.
-Consensus-aware Multi-task Selection (Generalist).
-The single-task data selection approach selects 20% of data per task, while our consensus-aware multi-task data selection approach selects a total of 20% data across all tasks.
-The Delta (%) column represents the relative drop in performance from specialist to generalist.
-Figure 3: Data Overlap between Specialists and Generalist Selection. It shows varying degrees of overlap, ranging from 3.27% (VQAv2 [9]) to 24.21% (LLAVA-W Bench [23]).
-Figure 4: Performance Comparison at Different Selection Ratios.
-Here we show the average relative performance comparisons of various data selection methods at different selection ratios.
-Our ICONS consistently outperforms all baseline approaches across different selection ratios, achieving 99.2% relative performance with 40% of the data, and remarkably exceeding 102% at 60% selection ratio.
-This shows that our selected subset is not only more efficient but can be even more effective than the full dataset for model training.
-3.3 Analysis
-From Specialist to Generalist.
-In order to understand the effectiveness of the intermediate task-specific influence matrices we obtained from the specialist stage, we further conduct experiments on selecting 20% of data for individual tasks.
-As shown in Tab.
-3, using 20% task-specific data, it achieves comparable or even superior performance to full data training in several cases (e.g., MME from 1476.9 to 1506.1 and SQA-I from 68.4 to 69.8).
-Then, through our influence consensus mechanism in the generalist stage, we identify consistently influential cross-tasks samples and select a 20% subset.
-This generalist approach maintains remarkably strong performance, with an average drop of 1.33% across most tasks compared to specialist baselines.
-As visualized in Fig.
-3, the overlap between specialist and generalist selections varies significantly across tasks, ranging from minimal overlap in simpler tasks like VQAv2 [9] (3.27%) and VizWiz [10] (3.28%) to substantial agreement in more complex tasks like LLAVA-W Bench [23] (24.21%).
-Notably, some tasks even show performance improvements under the generalist selection, such as SQA-I [25] (improving by 1.43%) and POPE [21] (improving by 1.04%).
-This empirically validates our consensus-based approach: by first understanding task-specific influence patterns and then building consensus across tasks, we can effectively identify a compact, universal training set that maintains strong performance across diverse vision-language tasks while significantly reducing the required training data.
-Different Selection Ratio.
-We evaluate our ICONS against baseline methods across different selection ratios, ranging from 5% to 60% of the training data.
+While COINCIDE achieves strong performance (97.4% Rel.), it falls short of ICONS on key tasks.
+Approaches like EL2N, Perplexity, SemDeDup achieve only 91-92% Rel., showing limitations in preserving performance.
+Dataset #Data Method VQAv2 GQA VizWiz SQA-I TextVQA POPE MME MMBench LLaVA-W Rel. (%)
+en cn Bench
+Vision-Flan-186K 186k Full 68.0 49.2 41.7 60.8 50.4 83.4 1,263.2 52.6 45.9 63.3 100.0
+37k Random 64.1 45.8 37.5 58.7 45.3 82.9 1,079.8 46.5 39.6 58.7 91.6
+ICONS (ours) 67.4 48.8 44.1 60.2 49.9 83.0 1,252.5 51.9 45.4 63.1 99.8
+Cambrian-7M 7,068k Full 80.2 62.9 58.4 75.3 60.9 86.5 1,524.6 69.1 58.9 67.6 100.0
+1,414k Random 74.2 57.5 61.9 71.0 57.1 86.4 1,465.7 63.3 49.6 70.4 95.4
+ICONS (ours) 79.6 62.1 60.7 73.9 59.8 86.2 1,503.1 67.8 55.8 67.0 98.8
+Table 2: Selection results on Vision-Flan-186K and Cambrian-7M.
+Performance comparison of different data selection approaches when trained on 20% of the Vision-Flan-186K visionflan and Cambrian-7M tong2024cambrian datasets.
+ICONS achieves strong performance (99.8% and 98.8% Rel.
+respectively) while using only 20% of the training data, significantly outperforming random selection which is one of the strongest baselines, and approaching full performance.
+Cambrian-7M & Vision-Flan-186K selection.
+We further provide results on Vision-Flan-186K and Cambrian-7M in Tab.
+2.
+On Vision-Flan-186K, our method achieves near-full performance (99.8% Rel.) using just 37k samples, significantly outperforming random selection (91.6%) across tasks.
+Similarly, on Cambrian-7M, ICONS maintains a strong performance (98.8% Rel.) with 1,414k samples, while random selection achieves 95.4%.
+These results demonstrate that our approach scales effectively to both small and large datasets, consistently preserving model capabilities while drastically reducing the training data required.
+Comparisons with representation-based data selection.
+We compare our method against RDS (Representation-based Data Selection)ivison2025large ; xia2024less , a strong baseline in language-only instruction tuning.
+RDS computes training-validation similarity using final-layer representation of the last token in each sequence instead of gradients.
+For a fair comparison, we use the same influence matrix formulation (Eqn.2) and apply majority voting to reach influence consensus.
+Our method consistently outperforms RDS across all tasks, particularly those requiring perceptual grounding – e.g., higher scores on GQA (60.7 vs.
+57.9), SQA-I (70.8 vs.
+68.0), and MME (1485.7 vs.
+1393.8).
+While RDS is effective in selecting large-scale text-only data (e.g.
+TULU-2/3 ivison2023camels ; lambert2024t ), its evaluation has largely focused on language-only tasks, where semantic similarity alone is often sufficient.
+In contrast, vision-language tasks demand alignment between modalities, where representation similarity is limited as it only reflects current embedding proximity and gradient-based approaches directly estimate each sample’s contribution to the validation loss.
+Our gradient-based approach directly estimates each sample’s impact on validation loss, capturing cross-modal training dynamics and prioritizing impactful training points.
+We further provide qualitative comparisons in Appendix §I.
+3.3 Analysis of aggregation strategies
+Ablations.
+As introduced in §2.4, we explore multiple strategies for aggregating task-specific influence rankings into a single compact subset.
+We compare our majority voting approach (Vote) with different aggregation approaches for combining task-specific influence scores: 1) score-based methods (Merge, Max) and their normalized variants (Merge-GausNorm, Merge-SumNorm), 2) rank-based methods (Round Robin, MinRank).
+Our voting-based strategy outperforms all alternatives (Tab.
+3), achieving the highest overall Rel.
+(98.6%).
+This shows that building a cross-task consensus via majority voting is a simple yet effective strategy for identifying consistently influential examples across tasks, without assuming calibration or comparability of scores.
+Aggregation VQAv2 GQA VizWiz SQA-I TextVQA POPE MME MMB (en) MMB (cn) LLaVA-W Rel. (%)
+Full 79.1 63.0 47.8 68.4 58.2 86.4 1476.9 66.1 58.9 67.9 100
+Merge 75.7 59.6 47.9 65.5 55.5 86.0 1422.1 59.0 51.0 66.2 96.4
+Max 75.2 59.8 48.1 66.2 55.5 85.5 1470.7 58.3 51.8 66.2 96.1
+Merge-GausNorm 75.1 60.1 46.4 69.8 54.5 85.6 1482.6 58.9 52.5 66.3 96.8
+Merge-SumNorm 75.5 59.1 51.7 68.7 43.5 87.1 1478.3 59.5 50.9 69.8 95.3
+Round Robin 75.4 59.1 48.3 70.6 55.2 86.6 1474.5 61.6 51.5 66.9 96.7
+MinRank 75.2 59.0 49.7 70.4 55.1 86.9 1456.3 61.1 52.4 68.4 97.1
+Vote (ours) 76.3 60.7 50.1 70.8 55.6 87.5 1485.7 63.1 55.8 66.1 98.6
+Table 3: Comparison of aggregation approaches.
+Performance of different influence aggregation methods when selecting 20% of the LLAVA-665K dataset.
+Our proposed aggregation approach (Vote) consistently achieves the best overall performance (98.6% Rel.), outperforming both score-based (Merge, Max), their noramlized variants (Merge-GausNorm, Merge-SumNorm) and rank-based (Round Robin, MinRank) baselines.
+Limitations of score-based aggregation.
+Score-based methods like Merge (Eqn.
+3) and Max (Eqn.
+4) assume calibrated influence scores across tasks, which is rarely the case.
+We observe substantial variation in the distribution of influence scores across tasks with standard deviations spanning from $8.15\times 10^{-3}$ (MME) to $1.26\times 10^{-3}$ (VQAv2), indicating that influence scores for MME are much more spread out, while those for VQAv2 are tightly concentrated.
+Similarly, mean influence scores vary in both magnitude and sign: MME has a relatively high positive mean ($1.68\times 10^{-3}$), while tasks like POPE ($-2.83\times 10^{-4}$) and GQA ($-8.89\times 10^{-5}$) have negative means.
+These divergent patterns show that certain tasks have much wider influence scores distributions, making a sample helpful for one task but neutral or harmful for another.
+Aggregating raw scores thus biases selection toward tasks with higher variance or skewed means.
+To address this calibration issue, we experimented with normalization strategies: Merge-SumNorm, which rescales each task’s influence scores by dividing them by a task-specific normalization factor (i.e., sum), and Merge-GausNorm, which normalizes the scores using task-wise mean and standard deviation before averaging:
+$I_{\text{Merge-SumNorm}}({\bm{z}}_{i})=\sum_{k=1}^{K}\frac{I_{k}({\bm{z}}_{i})}{\sum_{j}I_{k}({\bm{z}}_{j})}\qquad I_{\text{Merge-GausNorm}}({\bm{z}}_{i})=\sum_{k=1}^{K}\frac{I_{k}({\bm{z}}_{i})-\mu_{k}}{\sigma_{k}}$
+However, as shown in Tab. 3, both methods still underperform compared to our voting-based strategy, reinforcing the limitations of relying on score magnitudes directly.
+Limitations of rank-based aggregation.
+Rank-based methods sidesteps the challenge of comparing raw influence scores by focusing on within-task ranking.
+Round Robin selects samples by cycling through each task and picking the highest-scoring remaining sample for that task, ensuring balanced coverage.
+MinRank (Eqn.
+5) selects samples that have the best minimum rank across all tasks, prioritizing examples that perform exceptionally well in at least one task regardless of their performance in others.
+Although these methods ensure balanced coverage across tasks, they can overfit to outlier tasks.
+This is particularly evident with LLaVA-W Bench llava , which is an outlier in its influence ranking: both Round Robin and MinRank achieve relatively high scores on it (e.g.
+MinRank: 68.4), but this comes at the cost of lower performance on all other tasks (Tab.
+3).
+This suggests that purely rank-based selection can trade off some overall efficacy on the mainstream tasks and hurt multi-task balance.
+In contrast, our Vote approach is more robust and avoids this by focusing on multi-task consensus rather than forcing equal representation, yielding better balance and higher overall performance (98.6% Rel.), highlighting the importance of identifying broadly influential examples rather than narrowly optimizing per-task rankings.
+Figure 3: Pairwise overlap heatmap between specialists. The values show overlap percentages between benchmarks’ selected samples.
+Figure 4: Data overlap between specialists and generalist selection.
+Overlap varies significantly, from 3.27% (VQAv2) to 24.21% (LLAVA-W Bench), reflecting varying alignment between task-specific and consensus selections.
+Divergent multi-task influence patterns.
 As shown in Fig.
-4, our results reveal several key patterns: First, ICONS shows particularly strong performance in the low-selection regime (5-20%), where identifying the most influential samples is crucial.
-Another notable observation is that as the selection ratio increases, the performance gap between different methods gradually narrows.
-This convergence pattern is expected, as larger sample sizes naturally capture more of the dataset’s diversity and information.
-Despite this convergence trend, ICONS consistently maintains its performance edge across all selection ratios, and remarkably reaching above 100% at 60% selection ratio.
-One hypothesis is that ICONS can also effectively filter out potentially harmful or noisy training samples that might negatively impact model training, thereby surpassing the full training performance.
-(a) Pairwise Overlap Heatmap between Specialist Benchmarks.
-The values are the percentage of selected overlapping samples across different benchmarks, with high overlap seen between POPE and GQA (60.2%) and between VQAv2 and VizWiz (49.0%).
-Lower overlap, such as 3.3% between MMBench(en) and GQA, shows more distinct selected sample sets.
-(b) Pairwise Overlap Heatmap between Different Aggregation Approaches.
-The values are the percentage of overlapping selected samples between each pair of aggregation approaches.
-Our Vote-based influence consensus approach shows moderate overlap (53.8%-70.5%) with other methods, while Mean and Max selections are highly correlated (99.9%).
-Figure 5: Pairwise Overlap Heatmaps between Different Aggregation Approaches and Benchmarks.
-(a) compares the overlap between different specialist benchmarks, while (b) shows overlap between various aggregation methods.
-Divergent Cross-Task Influence Patterns.
-As shown in Fig.
-5(a), the pairwise overlap heatmap reveals significant variation in training data influence rankings across different tasks.
-High overlap, such as between VQAv2 and VizWiz (49.0%) or POPE and GQA (60.2%), suggests that certain samples are beneficial across similar tasks.
+4, the pairwise overlap heatmap shows notable variation in training data influence across tasks.
+High overlap – e.g., VQAv2 and VizWiz (49.0%) or POPE and GQA (60.2%), suggests that certain samples are beneficial across similar tasks.
 However, low overlap, like the 3.3% between MMBench (en) and GQA, highlights that highly influential samples for one task may have limited impact on others.
 Even closely related tasks, such as MMBench in different languages (English and Chinese), share 67.4% of influential samples.
-These findings empirically demonstrate significant overlap in influential samples across tasks, motivating the use of influence consensus for effective multi-task data selection.
-3.4 Transferability
-We evaluate the transferability of our data selection approach across two dimensions: unseen-task generalization and cross-architecture-scale generalization.
-Unseen-task Generalization.
-LLaVA-ICONS-133K demonstrates exceptional generalization on entirely unseen benchmarks that were not considered during our data selection process.
+To understand task-specific influence matrices from the specialist stage, we select the top 20% samples per task (Specialists).
+Overlap with our generalist subset (Fig.
+4) varies significantly, from minimal in tasks like VQAv2 (3.27%) and VizWiz (3.28%) to substantial agreement in tasks like LLAVA-W Bench llava (24.21%).
+These findings empirically demonstrate significant overlap in influential samples across tasks and validate our approach: by analyzing task-specific gradient-based influence patterns and building consensus across tasks, we can identify a compact subset that captures broadly useful samples across tasks, yielding strong performance with significantly less data.
+3.4 ICONS generalizes to unseen tasks
+LLaVA-ICONS-133K demonstrates exceptional generalization on entirely unseen benchmarks that were not used during data selection.
 As shown in Tab.
-4, we test across a diverse spectrum of tasks including MMVet [45], NaturalBench [19], AI2D [16], ChartQA [29], DocVQA [30], InfoVQA [31], RealWorldQA [41] and CMMMU [46].
-LLaVA-ICONS-133K achieves 95.5-113.9% (Rel.) compared to full dataset training.
-Notably, we observe improvements on InfoVQA (103.8%), NaturalBench (105.5%), RealWorldQA (104.4%), and CMMMU (113.9%), suggesting that, in some cases, training on LLaVA-ICONS-133K may even outperform training on the full dataset, despite these tasks not being included in the selection process.
-Importantly, LLaVA-ICONS-133K significantly outperforms random selection across all benchmarks - for example, improving performance on ChartQA (17.1 vs.
-15.1), MMVet (29.7 vs.
-27.6), CMMMU (25.2 vs.
-21.9) and NaturalBench (12.8 vs.
-11.1).
-This consistent improvement over random selection further validates the effectiveness of ICONS.
-This suggests that our selection approach successfully captures fundamental visual-language understanding capabilities that transfer well across different task formats and domains, and consistently outperforms random selection by a significant margin.
+4, we test across a diverse spectrum of tasks including MMVet mmvet , NaturalBench naturalbench , AI2D ai2d , ChartQA chartqa , DocVQA docvqa , InfoVQA infographicvqa , RealWorldQA xai2024grok and CMMMU cmmmu .
+LLaVA-ICONS-133K achieves 95.5-113.9% (Rel.) compared to full dataset training (InfoVQA: 103.8%, NaturalBench: 105.5%, RealWorldQA: 104.4%, and CMMMU: 113.9%), suggesting that, in some cases, training on LLaVA-ICONS-133K may even outperform training on the full dataset, despite these tasks not being included in the selection process.
+Importantly, LLaVA-ICONS-133K significantly outperforms random selection across all benchmarks.
+This suggests that our selection approach successfully captures fundamental visual-language understanding capabilities that transfer well across different task formats and domains.
+We further provide the cross-architecture generalization results in Appendix §D.
 AI2D ChartQA DocVQA InfoVQA MMVet Naturalbench RealworldQA CMMMU Rel. (%)
@@
 Table 4: Unseen-task generalization.
-Performance comparison on unseen benchmarks when trained on the selected subset.
-The relative performance ratios show that our selection approach successfully captures fundamental visual-language understanding capabilities that transfer well across different task formats and domains.
-Notably, we observe improvements on InfoVQA [31] (103.8%), MMVet [45] (95.5%), NaturalBench [19] (103.2%), RealWorldQA [41] (104.4%), and CMMMU [46] (114.0%), suggesting that in some cases, training on our LLaVA-ICONS-133K may even outperform using the full dataset.
-Cross-Architecture-Scale Generalization.
-We further conduct experiments on cross architecture scale generalization to evaluate the transferability of our selected data across different model scales.
-While our subset was initially selected using LLaVA-1.5-7B as the base model, we investigate whether these same examples remain effective for training larger architectures like LLaVA-1.5-13B.
-This tests whether our selection criteria identify universally valuable training examples rather than model-specific patterns.
-Our results in Tab.
-5 show cross architecture scale generalization, with the 7B-selected achieving 98.1% relative performance, suggesting that our selected subset captures fundamental visual-language understanding capabilities that scale effectively across model sizes.
-VQAv2 GQA Vizwiz SQA-I TextVQA POPE MME MMBench MMBench(cn) LLAVA-W Rel. (%)
-Full 80.0 63.3 58.9 71.2 60.2 86.7 1541.7 68.5 61.5 69.5 100.0
-Random 77.3 60.7 57.6 69.1 56.8 82.9 1517.2 63.2 56.3 67.5 95.7
-7B-selected 78.8 60.4 57.4 70.4 58.3 84.3 1527.5 64.9 59.7 68.2 97.3
-13B-selected 78.9 61.2 57.5 71.3 58.4 85.9 1535.2 66.1 59.8 68.8 98.1
-Table 5: Cross-Architecture-Scale Generalization.
-Our LLaVA-ICONS-133K selected via LLaVA-v1.5-7b-lora (7B-selected) shows strong cross-architecture transferability, achieving competitive performance when trained with data selected via LLaVA-v1.5-13b-lora (13B-selected).
-Specifically, while the 7B-selected achieves 97.3% relative performance, the 13B-selected model reaches 98.1%, showing that our selected subset’s effectiveness scales well to larger architectures.
-Both significantly outperform random selection (95.7%), with the 13B-selected option showing particular strength in complex reasoning tasks like MMBench and POPE.
-This consistency across model scales highlights that our data selection approach captures fundamental visual-language understanding patterns that generalize well across different model architecture scales.
-3.5 Ablation Studies
-Different Aggregation Approaches.
-To understand how different methods of combining task-specific influence scores affect the performance, we compare our majority voting approach (Vote) with four aggregation approaches: (1) Mean, which averages scores across tasks, (2) Max, which selects the highest score per sample, (3) Rank, which averages the relative ranking of samples within each task, (4) Norm, which normalize the scores before averaging:
-$I_{\text{mean}}({\bm{z}}_{i})=\frac{1}{K}\sum_{k=1}^{K}I_{k}({\bm{z}}_{i})\qquad I_{\text{norm}}({\bm{z}}_{i})=\frac{1}{K}\sum_{k=1}^{K}\frac{I_{k}({\bm{z}}_{i})-\mu_{k}}{\sigma_{k}}$
-$I_{\text{max}}({\bm{z}}_{i})=\max_{k=1,\ldots,K}I_{k}({\bm{z}}_{i})\qquad I_{\text{rank}}({\bm{z}}_{i})=\frac{1}{K}\sum_{k=1}^{K}\text{rank}(I_{k}({\bm{z}}_{i}))$
-Results in Tab.
-6 show that our Vote consistently outperforms alternatives across most tasks, achieving the highest overall relative performance (Rel.) of 98.6%.
-While methods like Rank perform well for MME (1490.0 vs.
-1485.7 with Norm), they underperform overall.
-To better understand the relationships between these aggregation approaches, we visualize their pairwise selection overlap in Fig.
-5(b).
-Specifically, the Mean and Max approaches show an extremely high overlap of 99.9%, suggesting they tend to select very similar influential samples.
-In contrast, our Vote-based approach exhibits a more moderate overlap with other methods, ranging from 53.8% to 70.5%, potentially those that are influential across a broader variety of tasks rather than overfitting to specific tasks.
-Task Mean Max Rank Norm Vote (ours)
-VQAv2 [9] 75.7 75.2 74.9 75.1 76.3
-GQA [13] 59.6 59.8 58.6 60.1 60.7
-VizWiz [10] 47.9 48.1 40.5 46.4 50.1
-SQA-I [25] 65.5 66.2 69.8 69.8 70.8
-TextVQA [37] 55.5 55.5 55.2 54.5 55.6
-POPE [21] 86.0 85.5 85.6 85.6 87.5
-MME [7] 1422.1 1470.7 1490.0 1482.6 1485.7
-MMBench (en) [47] 59.0 58.3 59.0 58.9 63.1
-MMBench (cn) [47] 51.0 51.8 50.8 52.5 55.8
-LLaVA-W Bench [23] 66.2 66.2 66.4 66.3 66.1
-Rel.(%) 96.4 96.1 95.9 96.8 98.6
-Table 6: Comparison of Different Aggregation Approaches.
-Performance comparison of different methods for combining task-specific influence scores, using 20% of training data.
-All scores are relative to full dataset training.
-Task Full Direct Consensus Delta (%)
-merge aware (ours)
-VQAv2 [9] 79.1 76.1 76.3 0.26
-GQA [13] 63.0 59.4 60.7 2.19
-VizWiz [10] 47.8 46.1 50.1 8.67
-SQA-I [25] 68.4 68.7 70.8 3.06
-TextVQA [37] 58.2 54.1 55.6 2.77
-POPE [21] 86.4 85.1 87.5 2.82
-MME [7] 1476.9 1419.2 1485.7 4.69
-MMBench (en) [47] 66.1 61.9 63.1 1.94
-MMBench (cn) [47] 58.9 50.3 55.8 10.94
-LLaVA-W Bench [23] 67.9 65.2 66.1 1.38
-Rel. (%) 100 94.7 98.6 -
-Table 7: Comparison of Consensus-aware vs.
-Direct-merge Selection.
-Performance comparison across different benchmarks using 20% of training data.
-Direct-merge computes influence directly on combined validation sets, while consensus-aware refers to our specialist-to-generalist approach.
-The Delta (%) column represents the relative improvement from Direct-merge to Consensus-aware.
-Consensus-aware outperforms Direct-merge across all benchmarks, while Direct-merge notably underperforms even the random baseline.
-Consensus-aware Selection vs.
-Direct-merge Selection.
-To validate our choice of a consensus-aware approach (first computing task-specific influence matrices and then aggregating results) versus a direct-merge approach (directly computing influence using a combined validation set), we compare the performance of both approaches.
-In the direct-merge approach, we combine validation sets from all tasks into a single pool, denoted as $\{{\bm{z}}^{\prime}\}_{\mathrm{merged}}$, and compute influence scores directly against this combined set: $\mathcal{I}_{\text{direct}}({\bm{z}}\rightarrow\{{\bm{z}}^{\prime}\}_{\mathrm{merged}})$.
-As described in Tab.
-7, our consensus-aware approach consistently outperforms direct-merge selection across all benchmarks, with the largest gains observed on complex tasks like MME [7].
-A potential reason could be the direct-merge approach does not account for the different vision-language task difficulties or validation set sizes across the target tasks.
-By pooling all validation examples, it effectively treats each validation data from all tasks equally, which can lead to the influence scores being dominated by tasks with more samples or with higher influence scores on average.
-The task-aware voting-based influence consensus mechanism of our direct-merge approach are more robust to these issues.
-Additionally, our approach is more scalable, as it avoids computing a single, large influence matrix across validation data from all target tasks.
-Adding new tasks only requires computing influence scores for those specific tasks and updating the voting mechanism, while direct-merge selection would need to recompute the entire large influence matrix which makes it less practical.
-4 Related Work
-Visual Instruction Tuning.
-Multimodal large language models (MLLMs), e.g., Flamingo [2], LLaVA [23], BLIP2 [20], and Cambrian [39], enhance the capabilities of large language models (LLMs) on various multimodal tasks.
-A key component in advancing MLLMs is visual instruction tuning [23], a training process that enables these models to interpret and follow instructions within a vision-language context, transforming them into versatile multimodal assistants.
-This tuning process not only improves the models’ instruction-following abilities but also aligns their outputs more closely with user expectations, thus enhancing their utility in practical applications [23].
-Data Selection.
-Data selection methods can be categorized based on the types of information they utilize for selection.
-Representation-based approaches [1, 18] leverage neural embeddings to capture data representations.
-Loss trajectory-based methods [32] prioritize data points that contribute most significantly to reducing generalization error over training.
-Gradient-based techniques [33, 42, 6] select data based on gradient information.
-Recent work has explored various approaches to select optimal visual instruction tuning datasets.
-Concurrent work TIVE [24] employs gradient-based selection to identify representative instances.
-TIVE assumes that the number of specialist data should be proportional to task difficulty and thus samples specialist data based on an estimation of task difficulty.
-Our method does not rely on this assumption – we directly select samples that benefit the greatest number of tasks.
-COINCIDE [18] clusters data based on representations associated with concept-skill compositions.
-Our work follows targeted instruction tuning selection approach LESS [42] to utilize gradient information to calculate the specialist influence (i.e., the influence on a specific task) and extends it to general scenarios by aggregating information from various tasks and selecting data for multiple downstream tasks via majority voting.
+Performance comparison on unseen benchmarks when trained on selected subsets.
+Notably, we observe improvements on InfoVQA (103.8%), RealWorldQA (104.4%), and CMMMU (114.0%), highlighting strong generalization to unseen tasks.
+3.5 ICONS outperforms baselines across ratios and exceeds full-data training at 60%
+Figure 5: Different selection ratios. ICONS consistently outperforms all baselines across different selection ratios and remarkably exceeding 102% at 60% selection ratio.
+To understand how ICONS scales, we evaluate it across different selection ratios, progressively scaling the subset size from 5% to 60% of LLaVA-665K.
+As shown in Fig.
+5, our results reveal several key patterns: First, ICONS shows particularly strong performance in the low-selection regime (5-20%), where identifying the most influential samples is crucial.
+Second, as the selection ratio increases, the performance gap between different methods gradually narrows.
+This convergence pattern is expected, as larger sample sizes naturally capture more of the dataset’s diversity and information.
+Despite this convergence trend, ICONS consistently outperforms all baselines across all selection ratios.
+Remarkably, it even surpasses full dataset performance at the 60% ratio, achieving over 102% relative score.
+One hypothesis is that ICONS can also effectively filter out potentially harmful or noisy training samples that might negatively impact model training, thereby surpassing the full training performance.
 4 Conclusion
 In this work, we introduce ICONS, a simple yet effective influence consensus-based approach for visual instruction tuning data selection.
-Our two-stage specialist-to-generalist method, first computes task-specific influence matrices and then select a subset by prioritizing universally influential samples through a cross-task voting mechanism.
-Through extensive experiments, we show that ICONS achieving 98.6% of full dataset performance using only 20% of the data and it consistently outperform baseline methods across different selection ratios.
-Additionally, we release LLaVA-ICONS-133K, a 20% subset of LLaVA-665K dataset, that not only maintains strong performance across diverse vision-language tasks but also shows transferability to unseen tasks.
+By leveraging gradient-based influence estimation and aggregating task-specific selections through majority voting, our two-stage specialist-to-generalist approach selects training examples that are broadly beneficial across multiple downstream tasks.
+ICONS addresses limitations of prior selection methods by avoiding assumptions about score comparability across tasks and reducing the sensitivity to outlier task rankings, which can bias selection in both score-based and rank-based approaches.
+Beyond data selection, it provides a principled way to reason about data influence in multitask mixtures.
+Through extensive experiments, we show that ICONS builds compact, high-impact datasets without sacrificing performance or generalization, achieving 98.6% of full dataset performance using only 20% of LLaVA-665K and generalizing well to unseen tasks or architectures.
+We release LLaVA-ICONS-133K and Cambrian-ICONS-1.4M, 20% subsets of the LLaVA-665K and Cambrian-7M datasets, maintaining strong performance on diverse tasks and transferring well to unseen ones.
 We hope our work inspires further exploration into data-efficient methods for vision-language models across diverse applications.
-Limitations.
-Our approach primarily faces one practical limitation: computing gradients for large training datasets is computationally expensive (§B.2).
-This computational overhead could potentially constrain the method’s applicability when working with extremely large-scale datasets.
-To support broader research community, we release our LLaVA-ICONS-133K dataset to help research iteration and model development under resource-constrained settings.
-Acknowledgments.
-This material is based upon work supported by the National Science Foundation under Grant No.
-2107048 and No.2112562.
-Any opinions, findings, and conclusions, or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the National Science Foundation.
-This work is also supported by the Singapore National Research Foundation and the National AI Group in the Singapore Ministry of Digital Development and Information under the AI Visiting Professorship Programme (award number AIVP-2024-001).
-We thank many people for their helpful discussion and feedback, listed in alphabetical order by last name: Allison Chen, Hamish Ivison, Carlos E.
-Jimenez, Polina Kirichenko, Jaewoo Lee, Tiffany Ling, Zhiqiu Lin, Shengbang Tong, Ethan Tseng, Justin Wang, Zirui Wang.
-Appendix for ICONS: Influence Consensus for Vision-Language Data Selection
-In this supplement, we provide additional experiment details and ablation studies (§A) and additional analysis (§B).
-We detail our algorithmic framework (§C) and discuss future research directions (§D).
-Finally, we include visualization of the selected data (§E) to better understand data samples with high influence.
-Appendix A Additional Experiment Details & Ablations
-Projection Dimension.
-We primarily set the projection dimension to 5120, reducing features from 338.7M to 5120 dimensions.
-The choice of 5120 was empirically validated for its trade-off between effective capturing gradient representation and maintaining a manageable parameter space.
-Our LLaVA-v1.5-7b-lora architecture includes a total of 7.4B parameters, with 338.7M parameters being trainable after LoRA adaptation, accounting for approximately 4.58% of the total parameter count.
-We further ablate different projection dimensions (1024, 2560, 5120, and 10240), with results provided in Fig.
-6.
-Warm-up Ratio.
-To initiate training, we use a warm-up set including 5% of the total training data.
-We conducted ablation studies to evaluate the impact of varying warm-up ratios (5%, 10%, 20%, and 100%) on selection performance, as shown in Fig.
-7.
-Our experiments reveal that increasing the warm-up data size does not lead to performance improvements.
-Surprisingly, models trained with smaller warm-up ratios (5-20%) consistently outperform those trained with the full dataset (100%).
-Specifically, the 5% warm-up ratio achieves the best performance at 98.6%, while using the complete dataset results in a performance drop to 97.8%.
-This finding suggests that a small subset of training data is sufficient and even beneficial for model initialization, and potentially gives better signals in the early training stages.
-Figure 6: Projection Dimension Ablation.
-We show the performance of ICONS at different projected dimensions (1024, 2560, 5120, 10240), compared to the random baseline.
-The performance increases with the projected dimension and reaches a plateau around dimension 5120.
-Figure 7: Warm-up Ratio Ablation.
-The blue line represents ICONS performance across different warm-up ratios (5%, 10%, 20%, and 100%), while the red dashed line shows the random baseline performance.
-Results show that smaller warm-up ratios (5-20%) achieve better performance compared to using the full dataset (100%).
-Appendix B Additional Analysis
-B.1 Consistency Analysis
-To evaluate the consistency of our approach, we conduct three independent runs of the experiment.
-As shown in Fig.
-8, our method demonstrates high consistency across different runs, achieving 98.6$\pm$1.2Rel., which shows a notable improvement over the random baseline, which achieves 95.8$\pm$2.7%.
-The lower standard deviation in our results (1.2% vs 2.7%) further indicates that our approach produces more stable and reliable outcomes compared to the random baseline.
-Figure 8: Rel. (%) Across Runs. We show the performance across three different runs for random selection and our ICONS. Our ICONS achieves 98.6$\pm$1.2while the random baseline reaches 95.8$\pm$2.7%.
-B.2 Computational Complexity
-Computing gradient-based influence requires significant computational resources.
+Appendix A Computational complexity
+A.1 Complexity analysis
+Computing gradient-based influence requires a non-trivial amount of computational resources.
 In the specialist stage, the complexity scales with both the dataset size $|\mathcal{D}|$ and the gradient dimension $d$.
@@
 The voting process then takes $\mathcal{O}(K\cdot|\mathcal{D}|)$ compute, followed by a final selection step with complexity $\mathcal{O}(|\mathcal{D}|\log|\mathcal{D}|)$ for sorting the aggregated votes.
-Storage requirements for this stage are minimal, primarily needing space for the final selected subset.
+Storage requirements for this stage are minimal, primarily for the final selected subset.
+A.2 Resource requirements
 In practice, for LLaVA-665K training data, the warmup training phase requires 0.75 hours using eight L40 GPUs.
@@
 This makes our method extendable for new tasks, as the expensive training data gradient computations only need to be performed once.
-B.3 Visual Dependency Influence Ranking
-Recent work [39] has shown that vision-language tasks vary in their reliance on visual information: task like MMBench [47] depends heavily on visual grounding, while others like SQA-I [25] can be handled primarily through language, showing only a 5% drop in performance when visual input is removed [39].
-To take visual dependency of training data into consideration, we further explored gradient-based Visual Dependency Score (VDS).
+A.3 Discussion on cost-benefit justification
+Although gradient-based data selection is computationally intensive, we argue that the initial cost is justified by three key considerations.
+First, the computational expense is largely a one-time investment: once gradients are computed, they can be stored in our gradient datastore and reused across multiple model iterations, target tasks, and diverse downstream applications.
+This reusability becomes especially valuable as the number of target datasets grows, because each new target dataset can leverage existing gradient computations, making the selection increasingly efficient at scale.
+Second, our empirical results demonstrate substantial performance benefits.
+Training on a strategically chosen 60% subset of data not only reduces training time but also surpasses the performance obtained by using the full dataset.
+This improvement underscores how directing more compute resources toward a carefully selected subset can yield higher returns on a fixed set of data.
+Lastly, the initial compute-intensive investment in data selection is amortized across future training iterations and future developers.
+By leveraging the curated, higher-quality dataset, they can substantially reduce training costs.
+Appendix B Related work
+B.1 Visual instruction tuning
+Multimodal large language models (MLLMs), e.g., Flamingo flamingo , LLaVA llava , BLIP2 blip-2 , and Cambrian (tong2024cambrian, ), enhance the capabilities of large language models (LLMs) on various multimodal tasks.
+A key component in advancing MLLMs is visual instruction tuning (llava, ), a training process that enables these models to interpret and follow instructions within a vision-language context, transforming them into versatile multimodal assistants.
+This tuning process not only improves the models’ instruction-following abilities but also aligns their outputs more closely with user expectations, thus enhancing their utility in practical applications (llava, ).
+B.2 Data selection
+Data selection methods survey can be categorized based on the types of information they utilize for selection.
+Representation-based approaches (semdedup, ; coincide, ) leverage neural embeddings to capture data representations.
+Loss trajectory-based methods (mindermann2022prioritized, ) prioritize data points that contribute most significantly to reducing generalization error over training.
+Gradient-based techniques (el2n, ; xia2024less, ; deng2024influential, ) select data based on gradient information.
+Recent work has explored various approaches to select optimal visual instruction tuning datasets.
+Concurrent work TIVE (liu2024less, ) employs gradient-based selection to identify representative instances.
+TIVE assumes that the number of specialist data should be proportional to task difficulty and thus samples specialist data based on an estimation of task difficulty.
+Our method does not rely on this assumption – we directly select samples that benefit the greatest number of tasks.
+COINCIDE (coincide, ) clusters data based on representations associated with concept-skill compositions.
+Our work follows targeted instruction tuning selection approach LESS xia2024less to utilize gradient information to calculate the specialist influence (i.e., the influence on a specific task) and extends it to general scenarios by aggregating information from various tasks and selecting data for multiple downstream tasks via majority voting.
+Appendix C Additional experiment details & ablations
+Task MME POPE SQA-I MMBench VQAv2 GQA VizWiz TextVQA LLaVA-W
+en cn
+$|\mathcal{D}_{\text{val}}|$ 986 500 424 1,164 1,164 1,000 398 800 84 84
+$|\mathcal{D}_{\text{test}}|$ 2,374 8,910 4,241 1,784 1,784 36,807 12,578 8,000 5,000 84
+Task Type Y/N Y/N MCQ MCQ MCQ VQA VQA VQA VQA VQA
+Table 5: Statistics of Target Tasks.
+Our target tasks include diverse benchmarks and answer formats, covering different vision-language capabilities.
+Task types include Multiple-Choice Questions (MCQ), Visual Question Answering (VQA), and Yes/No Questions (Y/N).
+C.1 Additional task details
+Here, we provide further details on the target tasks, as summarized in Tab.
+5.
+These tasks cover a wide range of multimodal benchmarks commonly used, including Yes/No questions (Y/N), multiple-choice understanding questions (MCQ) and visual question answering (VQA).
+C.2 Projection dimension
+We primarily set the projection dimension to 5120, reducing features from 338.7M to 5120 dimensions.
+The choice of 5120 was empirically validated for its trade-off between effective capturing gradient representation and maintaining a manageable parameter space.
+Our LLaVA-v1.5-7b-lora architecture includes a total of 7.4B parameters, with 338.7M parameters being trainable after LoRA adaptation, accounting for approximately 4.58% of the total parameter count.
+We further ablate different projection dimensions (1024, 2560, 5120, and 10240), with results provided in Fig.
+6(a).
+C.3 Warm-up Ratio
+To initiate training, we use 5% of the total training data.
+We conducted ablation studies to evaluate the impact of varying warm-up ratios (5%, 10%, 20%, and 100%) on selection performance, as shown in Fig.
+6(b).
+Our experiments reveal that increasing the warm-up data size does not lead to performance improvements.
+Surprisingly, models trained with smaller warm-up ratios (5-20%) consistently outperform those trained with the full dataset (100%).
+Specifically, the 5% warm-up ratio achieves the best performance at 98.6%, while using the complete dataset results in a performance drop to 97.8%.
+This finding suggests that a small subset of training data is sufficient and even beneficial for model initialization, and potentially gives better signals in the early training stages.
+(a) Projection Dimension Ablation.
+We show the performance of ICONS at different projected dimensions (1024, 2560, 5120, 10240), compared to the random baseline.
+The performance increases with the projected dimension and reaches a plateau around dimension 5120.
+(b) Warm-up Ratio Ablation.
+The blue line represents ICONS performance across different warm-up ratios (5%, 10%, 20%, and 100%), while the red dashed line shows the random baseline performance.
+Results show that smaller warm-up ratios (5-20%) achieve better performance compared to using the full dataset (100%).
+Figure 6: Ablation studies on (left) projection dimension and (right) warm-up ratio.
+Appendix D Additional analysis
+In this section, we provide additional analyses to better understand different aspects of our ICONS.
+We begin by analyzing the effectiveness of task-specific selections and their aggregation into a generalist subset (§D.1).
+We further explore whether incorporating visual dependency information into the selection process affects performance across different types of vision-language tasks (§D.2).
+We also evaluate the transferability of our selected subset across different model scales (§D.3).
+Additionally, we evaluate the consistency of our method across multiple runs (§D.4), showing its robustness and reliability.
+Method VQAv2 GQA VizWiz SQA-I TextVQA POPE MME MMBench (en) MMBench (cn) LLaVA-W Bench
+Full 79.1 63.0 47.8 68.4 58.2 86.4 1476.9 66.1 58.9 67.9
+Specialist 77.1 61.1 53.1 69.8 55.7 86.6 1506.1 66.0 56.4 67.1
+Generalist 76.3 60.7 50.1 70.8 55.6 87.5 1485.7 63.1 55.8 66.1
+Delta (%) 1.04 0.65 5.65 -1.43 0.18 -1.04 1.35 4.34 1.06 1.49
+Table 6: Single-task Selection (Specialist) vs.
+Consensus-aware Multi-task Selection (Generalist).
+The single-task data selection approach selects 20% of LLaVA-665K per task, while our consensus-aware multi-task data selection approach selects a total of 20% data across all tasks.
+D.1 From specialist to generalist
+To understand the intermediate task-specific influence matrices we obtained from the specialist stage, we select 20% of data for each individual task.
+The task-specific data (Specialists) achieves comparable or superior performance than full data training (Tab.
+6).
+With influence consensus at the generalist stage, we select a 20% subset with only a 1.33% average drop across tasks compared to specialist baselines.
+This validates our approach: by understanding task-specific influence patterns and building consensus across tasks, we can identify a compact, universal training set that maintains strong performance with significantly less data.
+D.2 Visual dependency influence ranking
+Recent work tong2024cambrian has shown that vision-language tasks vary in their reliance on visual information: tasks like MMBench mmbench depend heavily on visual grounding, while others like SQA-I sqa can be handled primarily through language, showing only a 5% drop in performance when visual input is removed tong2024cambrian .
+To take visual dependency of training data into account, we further explored gradient-based Visual Dependency Score (VDS).
 For each data point, we calculate the gradient of the model’s auto-regressive cross-entropy loss with both the original image and a Gaussian noise image $I_{\text{noise}}\sim\mathcal{N}(0,1)$, keeping the text input constant.
-This quantifies how much the visual component contributes to the model performance.
-We construct an adapted influence matrix: visual influence matrix $\mathcal{I}_{\texttt{VDS}}\in\mathbb{R}^{|\mathcal{D}|\times|\mathcal{D}_{\text{val}}|}$, which quantifies the visual influence of each training sample ${\bm{z}}_{i}$ on each validation sample ${\bm{z}}^{\prime}_{j}$ with respect to the model’s gradient alignment and visual dependency.
+This quantifies how much the visual component contributes to model performance.
+We construct an adapted influence matrix: visual influence matrix $\mathcal{I}_{\texttt{VDS}}\in\mathbb{R}^{|\mathcal{D}|\times|\mathcal{D}_{\text{val}}|}$, which quantifies the visual influence of each training sample ${\bm{z}}_{i}$ on each validation sample ${\bm{z}}^{\prime}_{j}$ w.r.t the model’s gradient alignment and visual dependency.
 $\mathcal{I}_{\text{VDS}}$ is computed as:
@@
 Our empirical results demonstrate that VDS-based data selection has varying effectiveness across different vision-language tasks (Tab.
-8).
+7).
 The approach shows substantial improvements on tasks requiring strong visual understanding, such as open-ended generation (LLaVA-W Bench: +2.72%) and multiple-choice understanding (MMBench-EN: +1.90%, MMBench-CN: +0.90%).
 However, tasks that primarily rely on textual reasoning show decreased performance, including SQA-I (-1.84%) and TextVQA (-1.44%).
-These results align with and extend the findings in Cambrian [39], demonstrating that the effectiveness of VDS corresponds to a task’s visual dependency - tasks that maintain performance without visual inputs show limited or negative impact from VDS-based selection, while visually-dependent tasks benefit significantly.
-This pattern suggests that VDS effectively identifies training samples where visual information plays an important role in model learning.
-Task w/o VDS w/ VDS Delta (%)
-VQAv2 [9] 76.3 75.8 -0.66
-GQA [13] 60.7 60.9 +0.33
-VizWiz [10] 50.1 50.3 +0.40
-SQA-I [25] 70.8 69.5 -1.84
-TextVQA [37] 55.6 54.8 -1.44
-POPE [21] 87.5 86.8 -0.80
-MME [7] 1485.7 1489.3 +0.24
-MMBench (en) [47] 63.1 64.3 +1.90
-MMBench (cn) [47] 55.8 56.3 +0.90
-LLaVA-W Bench [23] 66.1 67.9 +2.72
+These results align with and extend the findings in Cambrian tong2024cambrian , demonstrating that the effectiveness of VDS corresponds to a task’s visual dependency - tasks that maintain performance without visual inputs show limited or negative impact from VDS-based selection, while visually-dependent tasks benefit significantly.
+This pattern suggests that VDS effectively identifies training samples where visual information plays an important role in training.
+VQAv2 GQA VizWiz SQA-I TextVQA POPE MME MMBench (en) MMBench (cn) LLaVA-W Bench
+Full 79.1 63.0 47.8 68.4 58.2 86.4 1476.9 66.1 58.9 67.9
+w/o VDS 76.3 60.7 50.1 70.8 55.6 87.5 1485.7 63.1 55.8 66.1
+w/ VDS 75.8 60.9 50.3 69.5 54.8 86.8 1489.3 64.3 56.3 67.9
+Delta (%) -0.66 +0.33 +0.40 -1.84 -1.44 -0.80 +0.24 +1.90 +0.90 +2.72
 Table 7: Impact of Visual Dependency Score (VDS) on Selection Performance.
-Performance comparison between influence calculation with and without VDS.
-VDS shows strongest improvements on LLaVA-W Bench (+2.72%), MMBench (en) (+1.90%), and MMBench (cn) (+0.90%), while tasks like SQA-I (-1.84%), TextVQA (-1.44%), and POPE (-0.80%) show significant performance drops.
-Appendix C Algorithm Details
+Rows show performance without VDS, with VDS, and the performance change (Delta).
+VDS improves performance on LLaVA-W Bench (+2.72%), MMBench (en) (+1.90%), and MMBench (cn) (+0.90%), but decreases performance on SQA-I (-1.84%), TextVQA (-1.44%), and POPE (-0.80%).
+VQAv2 GQA Vizwiz SQA-I TextVQA POPE MME MMBench (en) MMBench (cn) LLAVA-W Rel. (%)
+Full 80.0 63.3 58.9 71.2 60.2 86.7 1541.7 68.5 61.5 69.5 100.0
+Random 77.3 60.7 57.6 69.1 56.8 82.9 1517.2 63.2 56.3 67.5 95.7
+7B-selected 78.8 60.4 57.4 70.4 58.3 84.3 1527.5 64.9 59.7 68.2 97.3
+13B-selected 78.9 61.2 57.5 71.3 58.4 85.9 1535.2 66.1 59.8 68.8 98.1
+Table 8: Cross-Architecture Generalization.
+Our LLaVA-ICONS-133K selected via LLaVA-v1.5-7B model (7B-selected) shows strong cross-architecture transferability, achieving 97.3% Rel., while the data selected via LLaVA-v1.5-13B model (13B-selected) reaches 98.1%, showing that our selected subset generalizes well to different architectures.
+D.3 Cross-architecture generalization
+We further conduct experiments on cross architecture generalization to evaluate the transferability of our selected data across different model scales.
+While our subset was initially selected using LLaVA-v1.5-7B as the base model, we investigate whether these same examples remain effective for training larger architectures like LLaVA-v1.5-13B.
+This tests whether our selection criteria identify universally valuable training examples rather than model-specific patterns.
+Our results in Tab.
+8 show cross-architecture generalization, with 13B model trained on 7B-selected subset achieving 98.1% Rel..
+Both 7B-selected and 13B-selected subsets outperform random selection (95.7%), with the 13B-selected option showing particular strength in reasoning tasks like MMBench and POPE.
+This suggests that our selected subset captures fundamental visual-language understanding patterns that generalize well across different model architectures.
+Figure 7: Rel. (%) Across Runs. We show the performance across three different runs for random selection and our ICONS.
+D.4 Consistency analysis
+To evaluate the consistency of ICONS, we conduct three independent runs of the experiments.
+As shown in Fig.
+7, our method demonstrates high consistency across different runs, achieving 98.6$\pm$1.2Rel., which shows a notable improvement over random baseline, which achieves 95.8$\pm$2.7%.
+The lower standard deviation (1.2% vs 2.7%) further indicates that our approach produces more stable and reliable outcomes compared to the random baseline.
+Appendix E Algorithm details
 We provide detailed pseudocode for our two-stage ICONS framework.
@@
 Algorithm 1 ICONS Stage 1: Specialist (Task-specific Influence Computation)
-0:  Training dataset $D=\{(x_{i},I_{i},y_{i})\}_{i=1}^{N}$, target tasks $T=\{T_{1},...,T_{K}\}$
-0:  Warm-up ratio $r$ (default 5%)
-0:  Task-specific influence scores $\{\bar{I}_{k}\}_{k=1}^{K}$
-1:  for each task $T_{k}\in T$ do
-2:     // Step 1: Warm-up Training
-3:     Sample warm-up set $D_{\text{warmup}}\subset D$ of size $r|D|$
-4:     $f_{\text{warmup}}\leftarrow\text{LoRA}(f_{\text{base}},D_{\text{warmup}})$
-5:     // Step 2: Gradient Computation
-6:     for each training data $z_{i}\in D$ do
-7:        $g_{i}\leftarrow\nabla_{\theta_{w}}L(f_{\text{warmup}}(z_{i}),y_{i})$
-8:        $\tilde{g}_{i}\leftarrow\text{Normalize}(Rg_{i})$Random projection}
-9:     end for
-10:     for each validation data $z^{\prime}_{j}\in D_{\text{val}}^{k}$ do
-11:        $g^{\prime}_{j}\leftarrow\nabla_{\theta_{w}}L(f_{\text{warmup}}(z^{\prime}_{j}),y^{\prime}_{j})$
-12:        $\tilde{g}^{\prime}_{j}\leftarrow\text{Normalize}(Rg^{\prime}_{j})$
-13:     end for
-14:     // Step 3: Influence Matrix Computation
-15:     for each $z_{i}\in D$, $z^{\prime}_{j}\in D_{\text{val}}^{k}$ do
-16:        $I_{ij}^{k}\leftarrow\langle\tilde{g}_{i},\tilde{g}^{\prime}_{j}\rangle$
-17:     end for
+1:Training dataset $D=\{(x_{i},I_{i},y_{i})\}_{i=1}^{N}$, target tasks $T=\{T_{1},...,T_{K}\}$
+2:Warm-up ratio $r$ (default 5%)
+3:Task-specific influence scores $\{\bar{I}_{k}\}_{k=1}^{K}$
+4:for each task $T_{k}\in T$ do
+5:   // Step 1: Warm-up Training
+6:   Sample warm-up set $D_{\text{warmup}}\subset D$ of size $r|D|$
+7:   $f_{\text{warmup}}\leftarrow\text{LoRA}(f_{\text{base}},D_{\text{warmup}})$
+8:   // Step 2: Gradient Computation
+9:   for each training data $z_{i}\in D$ do
+10:      $g_{i}\leftarrow\nabla_{\theta_{w}}L(f_{\text{warmup}}(z_{i}),y_{i})$
+11:      $\tilde{g}_{i}\leftarrow\text{Normalize}(Rg_{i})$Random projection}
+12:   for each validation data $z^{\prime}_{j}\in D_{\text{val}}^{k}$ do
+13:      $g^{\prime}_{j}\leftarrow\nabla_{\theta_{w}}L(f_{\text{warmup}}(z^{\prime}_{j}),y^{\prime}_{j})$
+14:      $\tilde{g}^{\prime}_{j}\leftarrow\text{Normalize}(Rg^{\prime}_{j})$
+15:   // Step 3: Influence Matrix Computation
+16:   for each $z_{i}\in D$, $z^{\prime}_{j}\in D_{\text{val}}^{k}$ do
+17:      $I_{ij}^{k}\leftarrow\langle\tilde{g}_{i},\tilde{g}^{\prime}_{j}\rangle$
 18:   // Compute average influence per training sample
 19:   $\bar{I}_{k}(z_{i})\leftarrow\frac{1}{|D_{\text{val}}^{k}|}\sum_{j=1}^{|D_{\text{val}}^{k}|}I_{ij}^{k}$
-20:  end for
-21:  return  Task-specific influence scores $\{\bar{I}_{k}\}_{k=1}^{K}$
+20:return Task-specific influence scores $\{\bar{I}_{k}\}_{k=1}^{K}$
 Algorithm 2 ICONS Stage 2: Generalist (Influence Consensus-based Data Selection)
-0:  Task-specific influence scores $\{\bar{I}_{k}\}_{k=1}^{K}$
-0:  Selection ratio $p$, number of tasks $K$
-0:  Selected subset $S\subset D$ of size $m\ll N$
-1:  // Compute voting thresholds
-2:  for each task $T_{k}\in T$ do
-3:     $\tau_{k}\leftarrow(1-p)\text{-th percentile of }\{\bar{I}_{k}(z_{i})\}_{i=1}^{N}$
-4:  end for
-5:  // Voting process
-6:  for each training sample $z_{i}\in D$ do
-7:     $I_{\text{vote}}(z_{i})\leftarrow 0$
-8:     for each task $T_{k}\in T$ do
-9:        $\text{vote}_{k}(z_{i})\leftarrow\mathbbm{1}[\bar{I}_{k}(z_{i})\geq\tau_{k}]$
-10:        $I_{\text{vote}}(z_{i})\leftarrow I_{\text{vote}}(z_{i})+\text{vote}_{k}(z_{i})$
-11:     end for
-12:  end for
+1:Task-specific influence scores $\{\bar{I}_{k}\}_{k=1}^{K}$
+2:Selection ratio $p$, number of tasks $K$
+3:Selected subset $S\subset D$ of size $m\ll N$
+4:// Compute voting thresholds
+5:for each task $T_{k}\in T$ do
+6:   $\tau_{k}\leftarrow(1-p)\text{-th percentile of }\{\bar{I}_{k}(z_{i})\}_{i=1}^{N}$
+7:// Voting process
+8:for each training sample $z_{i}\in D$ do
+9:   $I_{\text{vote}}(z_{i})\leftarrow 0$
+10:   for each task $T_{k}\in T$ do
+11:      $\text{vote}_{k}(z_{i})\leftarrow\mathbbm{1}[\bar{I}_{k}(z_{i})\geq\tau_{k}]$
+12:      $I_{\text{vote}}(z_{i})\leftarrow I_{\text{vote}}(z_{i})+\text{vote}_{k}(z_{i})$
 13:// Select top samples based on total votes
@@
 15:return Selected subset $S$
-Appendix D Future Work
+Appendix F Limitations
+Our approach primarily faces one practical limitation: computing gradients for large training datasets is computationally expensive (Appendix §A).
+This computational overhead could potentially constrain the method’s applicability when working with extremely large-scale datasets.
+To support broader research community, we release LLaVA-ICONS-133K dataset to help research iteration and model development under resource-constrained settings.
+Appendix G Broader impact
+Our exploration focuses on scientific understanding and practical applications of vision-language data selection.
+While our work does not directly imply negative impacts, it may indirectly propagate existing biases present in the original datasets.
+Therefore, it is important to incorporate rigorous bias-mitigation measurements for data selection.
+On the positive side, our method enables more efficient and sustainable model development by reducing data redundancy, computational cost while maintaining or even improving performance.
+Discussion on these critical aspects should remain a priority as we further explore the potential of vision-language data selection.
+Appendix H Future work
 Our work opens several promising research directions for improving vision-language data selection.
@@
 Additionally, investigating more efficient gradient computation and storage methods would help scale these methods to larger datasets while maintaining strong performance across diverse vision-language tasks.
-Appendix E Visualizations
+Appendix I Visualizations
+I.1 Representation-based vs. Gradient-based data selection
+While ICONS leverages gradient-based influence signals, we explore how representation-based data selection (RDS) performs in the same setting (§3.2).
+We analyze the top-ranked training examples selected by RDS after the generalist stage in Fig.
+8 vs.
+samples selected by ICONS in Fig.
+19.
+Interestingly, we observe that the representation-based variants often favor training examples with repeated images or instructions, which may dominate the learned representations without contributing to better generalization.
+Some of the highest-scoring samples under representation-based similarity are duplicated image-question pairs with only the answer choices shuffled.
+We hypothesize that this is a side effect of the way multimodal representations are constructed—where visually dominant or textually redundant samples occupy high-density regions in embedding space.
+However, these samples do not necessarily translate into broader utility across tasks, as seen in the performance gap in Tab.
+1.
+This discrepancy raises broader questions about what it means for multimodal data to be diverse.
+While we leave these questions open for future exploration, our results suggest that gradient-based influence, though computationally more expensive, is better aligned with generalization and multi-task data mixture settings.
+Figure 8: Top three samples selected by RDS.
+These highly-ranked samples are selected due to representation-based similarity but do not necessarily contribute to better generalization, highlighting a key limitation of representation-based selection in multimodal settings.
+I.2 Specialists & Generalist
 We visualize the most influential top three examples across specialists (figs.
@@
 This suggests that the most effective training samples are those that combine multiple types of reasoning while maintaining clear, unambiguous ground truth that can be consistently learned across tasks.
-Figure 9: VQAv2. Top-left: A sample from VQAv2 [9]. Remaining panels show top three influential samples selected using the specialist influence ranking step.
-Figure 10: GQA. Top-left: A sample from GQA [13]. Remaining panels show top three influential samples selected using the specialist influence ranking step.
-Figure 11: VizWiz. Top-left: A sample from VizWiz [10]. Remaining panels show top three influential samples selected using the specialist influence ranking step.
-Figure 12: SQA. Top-left: A sample from SQA [25]. Remaining panels show top three influential samples selected using the specialist influence ranking step.
-Figure 13: TextVQA. Top-left: A sample from TextVQA [37]. Remaining panels show top three influential samples selected using the specialist influence ranking step.
-Figure 14: Pope. Top-left: A sample from Pope [21]. Remaining panels show top three influential samples selected using the specialist influence ranking step.
-Figure 15: MME. Top-left: A sample from MME [7]. Remaining panels show top three influential samples selected using the specialist influence ranking step.
-Figure 16: MMBench (en). Top-left: A sample from MMBench (en) [47]. Remaining panels show top three influential samples selected using the specialist influence ranking step.
-Figure 17: MMBench (cn). Top-left: A sample from MMBench (cn) [47]. Remaining panels show top three influential samples selected using the specialist influence ranking step.
-Figure 18: LLaVA-W Bench. Top-left: A sample from LLaVA-W Bench [23]. Remaining panels show top three influential samples selected using the specialist influence ranking step.
-Figure 19: Generalist. We show top three influential samples selected after the generalist step.
+Figure 9: VQAv2. Top-left: A sample from VQAv2 vqav2 . Remaining panels show top three influential samples selected using the specialist influence ranking step.
+Figure 10: GQA. Top-left: A sample from GQA gqa . Remaining panels show top three influential samples selected using the specialist influence ranking step.
+Figure 11: VizWiz. Top-left: A sample from VizWiz vizwiz . Remaining panels show top three influential samples selected using the specialist influence ranking step.
+Figure 12: SQA. Top-left: A sample from SQA sqa . Remaining panels show top three influential samples selected using the specialist influence ranking step.
+Figure 13: TextVQA. Top-left: A sample from TextVQA textvqa . Remaining panels show top three influential samples selected using the specialist influence ranking step.
+Figure 14: Pope. Top-left: A sample from Pope pope . Remaining panels show top three influential samples selected using the specialist influence ranking step.
+Figure 15: MME. Top-left: A sample from MME mme . Remaining panels show top three influential samples selected using the specialist influence ranking step.
+Figure 16: MMBench (en). Top-left: A sample from MMBench (en) mmbench . Remaining panels show top three influential samples selected using the specialist influence ranking step.
+Figure 17: MMBench (cn). Top-left: A sample from MMBench (cn) mmbench . Remaining panels show top three influential samples selected using the specialist influence ranking step.
+Figure 18: LLaVA-W Bench. Top-left: A sample from LLaVA-W Bench llava . Remaining panels show top three influential samples selected using the specialist influence ranking step.
+Figure 19: Generalist. We show top three influential samples selected after the generalist stage.

### v4

@@
 Abstract
-Training vision-language models via instruction tuning often relies on large mixtures of data spanning diverse tasks and domains.
-However, these mixtures frequently include redundant information, increasing computational costs without proportional performance gains—necessitating more effective data selection strategies.
-Existing methods typically rely on task-agnostic heuristics to estimate data importance or focus on optimizing single tasks in isolation, limiting their effectiveness in multitask settings.
-In this work, we introduce ICONS, a gradient-based Influence CONsensus approach for vision-language data Selection.
-Our method leverages first-order training dynamics to estimate the influence of individual training examples on validation performance and aggregates these estimates across tasks via majority voting over task-specific influences.
-This cross-task consensus identifies data points that are consistently valuable across tasks, enabling us to prioritize examples that drive overall performance.
-The voting-based design further mitigates issues such as score calibration and outlier sensitivity, resulting in robust and scalable data selection for diverse multitask mixtures.
-With only 20% of the data from LLaVA-665K and Cambrian-7M, our selected subsets retain 98.6% and 98.8% of the performance achieved with full datasets—and can even surpass full-data training at a 60% selection ratio on LLaVA-665K.
-Our approach also generalizes to unseen tasks and architectures, demonstrating strong transfer.
-We release two compact, high-utility subsets—LLaVA-ICONS-133K and Cambrian-ICONS-1.4M —preserving impactful training examples for efficient and scalable vision-language model development.
+Training vision-language models via instruction tuning relies on large data mixtures spanning diverse tasks and domains, yet these mixtures frequently include redundant information that increases computational costs without proportional gains.
+Existing methods typically rely on task-agnostic heuristics to estimate data importance, limiting their effectiveness across tasks.
+We introduce ICONS, a gradient-based Influence CONsensus approach for vision-language data Selection.
+Our method leverages first-order training dynamics to estimate each example’s influence on validation performance, then aggregates these estimates across tasks via majority voting.
+This cross-task consensus identifies consistently valuable data points while mitigating score calibration and outlier sensitivity, enabling robust and scalable data selection for diverse multitask mixtures.
+Models trained on our selected 20% data subset from LLaVA-665K (respectively: from Cambrian-7M, from Vision-Flan-186K) retain 98.6% (respectively: 98.8%, 99.8%) of full-dataset performance.
+We demonstrate that our selected data generalizes to unseen tasks and model architectures, and release three compact subsets LLaVA-ICONS-133K, Cambrian-ICONS-1.4M, and Vision-Flan-ICONS-37K for efficient vision-language model development.
 1 Introduction
-Visual instruction tuning is a crucial step in training multimodal language models liu2024improved ; llava , enabling them to follow language instructions grounded in visual content.
-Recent approaches rely on large-scale datasets such as LLaVA-665K liu2024improved and Cambrian-7M tong2024cambrian , which contain 665K and 7M examples, respectively.
-While effective, these datasets introduce significant barriers to iteration and deployment: prolonged training times brown2020language ; kaplan2020scaling , high storage demands rajbhandari2020zero ; gadre2024datacomp , and substantial compute costs maini2024rephrasing ; wu2022sustainable .
-Moreover, not all examples contribute equally to all tasks—naively scaling up diverse data mixtures can introduce redundancy and inefficiency.
+Figure 1: Influence consensus for vision-language data selection.
+(Left) Given a large-scale visual instruction tuning dataset (LLaVA-665K), our method uses majority voting across task-specific influence scores to identify training samples that are consistently influential across multiple tasks, forming a compact 20% subset (LLaVA-ICONS-133K) with data points achieving influence consensus.
+(Right) The radar plot compares performance between LLaVA-665K and our selected subset, showing the selected subset achieves comparable results to the full dataset.
+Visual instruction tuning (VIT) is a crucial step in training vision-language models (VLMs) [liu2024improved, llava], enabling them to follow language instructions grounded in visual content.
+Recent approaches rely on large-scale instruction following datasets such as LLaVA-665K [liu2024improved] and Cambrian-7M [tong2024cambrian].
+While effective, these datasets introduce significant barriers to model training and deployment: prolonged training times [brown2020language, kaplan2020scaling], high storage demands [rajbhandari2020zero, gadre2024datacomp], and substantial compute costs [maini2024rephrasing, wu2022sustainable].
+Moreover, not all training samples contribute equally to diverse downstream tasks and naively scaling up diverse data mixtures can introduce redundancy and inefficiency.
 This raises a fundamental question:
-{mdframed}
-[leftmargin=0pt, rightmargin=0pt, innerleftmargin=10pt, innerrightmargin=10pt, skipbelow=0pt]
-Can we identify a compact, multitask-effective subset of training data that preserves model capabilities arcoss tasks while enabling faster experimentation?
-Prior work has explored various data selection strategies, including gradient-based approaches xia2024less ; deng2024influential , influence functions yang2022dataset ; koh2017understanding , and diversity-based sampling yu2024diversify ; bukharin2023data .
+Can we identify a compact, multitask-effective subset of training data that preserves model capabilities across tasks while enabling faster experimentation?
+Prior work has explored various data selection strategies, including gradient-based approaches [xia2024less, deng2024influential], influence functions [yang2022dataset, koh2017understanding], and diversity-based sampling [yu2024diversify, bukharin2023data].
 However, many of these methods either optimize for single tasks in isolation or maximize source diversity without aligning to downstream needs.
-In multitask visual instruction tuning, this is particularly limiting: optimizing for one task may hurt generalization, and task-agnostic diversity may dilute impact.
-Rather than selecting
-Figure 1: Influence consensus for vision-language data selection.
-(Left) Given a large scale visual instruction tuning dataset (LLaVA-665K), our method uses majority voting across task-specific influence scores to identify training samples that are consistently influential across multiple tasks, forming a compact 20% subset (LLaVA-ICONS-133K) with data points achieving influence consensus.
-(Right) The radar plot compares performance between LLaVA-665K and our selected subset, showing the selected subset achieves comparable results to the full dataset.
-data based on per-task influence, we aim to identify samples that are broadly useful—training examples that consistently contribute across tasks.
-To do this, we aggregate gradient-based influence scores using a simple yet effective majority voting scheme.
-We introduce ICONS (Influence CONsensus vision-language data Selection), a method that builds upon the gradient-based selection approach LESS xia2024less .
-Given access to validation data for each target task, our method: (1) computes first-order gradient influence scores to measure how each training sample impacts task-specific performance, and (2) uses influence consensus through majority voting to identify training samples that show consistent positive value across multiple tasks.
+In the context of VIT data that supports diverse vision-language tasks, this is particularly limiting: optimizing for one task may hurt generalization, and task-agnostic diversity may dilute impact.
+Rather than selecting data based on per-task influence, we aim to identify samples that are broadly useful, with training examples that consistently contribute across tasks.
+To do this, we aggregate gradient-based influence scores using a simple yet effective majority voting approach to reach influence consensus.
+We introduce ICONS (Influence CONsensus vision-language data Selection), a method that builds upon LESS [xia2024less], a gradient-based selection approach.
+Given access to validation data for each target task, our method: (1) computes first-order gradient influence scores to measure how each training sample impacts task-specific performance, and (2) reaches influence consensus via majority voting to identify training samples that show consistent positive impact across multiple tasks.
 This consensus-based mechanism identifies universally valuable training examples: while some samples might be highly influential for individual tasks, we prioritize those that demonstrate broad utility across the task spectrum.
-While the computational cost of influence estimation is expensive, this front-loaded, one-time investment yields a standardized, compact dataset that can significantly accelerate development of multimodal models, and enables reusable gradient datastores that amortize costs across iterations and deliver long-term savings.
-Using ICONS, we create LLaVA-ICONS-133K and Cambrian-ICONS-1.4M, automatically curated 20% subsets of the LLaVA-665K dataset liu2024improved and Cambrian-7M tong2024cambrian dataset, respectively.
-These compact datasets maintain 98.6% and 98.8% of their original performance across multiple vision-language tasks, providing significant improvements over randomly selecting same-sized subsets (95.8% and 95.4%) and eliminating approximately two-thirds of the performance drop from shrinking the training data.
+While the computational cost of influence estimation is high, this front-loaded, one-time investment yields a standardized, compact dataset that can significantly accelerate development of multimodal models, and enables reusable gradient datastores that amortize costs across iterations and deliver long-term savings.
+Using ICONS, we create LLaVA-ICONS-133K, Cambrian-ICONS-1.4M, and Vision-Flan-ICONS-37K, automatically selected 20% subsets of LLaVA-665K [liu2024improved], Cambrian-7M [tong2024cambrian], and Vision-Flan-186K [visionflan] datasets, respectively.
+These compact datasets maintain 98.6%, 98.8%, and 99.8% of their original performance across multiple vision-language tasks, providing significant improvements over randomly selecting same-sized subsets (95.8%, 95.4%, and 91.6%) and eliminating approximately two-thirds of the performance drop from shrinking the training data.
 Moreover, our ICONS outperforms all baselines across different selection ratios, and remarkably achieves above-full-dataset performance, surpassing the original datasets at a 60% selection ratio for LLaVA-665K.
-Importantly, the selected subset shows strong transferability, e.g., LLaVA-ICONS-133K maintains 95.5-113.9% relative performance across unseen tasks, suggesting that ICONS identifies fundamentally valuable training data.
+Importantly, the selected subsets show strong transferability, e.g., LLaVA-ICONS-133K maintains 95.5-113.9% relative performance across unseen tasks, suggesting that ICONS identifies fundamentally valuable training data.
 We summarize our key contributions:
@@
     Our consensus-based selection outperforms all baselines (§4.2) and we ablate influence aggregation strategies and show the advantage of voting-based consensus (§4.3).
-We further show that ICONS exceeds 102% of full-dataset performance at a 60% selection ratio on LLaVA-665K3.5).
+ICONS exceeds 102% of full-dataset performance at a 60% selection ratio on LLaVA-665K4.5).
   3. 3.
-    We release LLaVA-ICONS-133K and Cambrian-ICONS-1.4M, compact 20% subsets of LLaVA-665K and Cambrian-7M respectively, achieving near-full performance (98.6% and 98.8%), transferring well to unseen tasks (§3.4), and serving as standardized training sets for resource-efficient development.
-2 Influence consensus for vision-language data selection
+    We release subsets LLaVA-ICONS-133K, Cambrian-ICONS-1.4M and Vision-Flan-ICONS-37K for resource-efficient development, demonstrating competitive performance including generalization to unseen tasks (§4.4) despite a 5x reduction in training data.
+2 Related work
+Visual instruction tuning.
+Vision-language models (VLMs), e.g., Flamingo [flamingo], LLaVA [llava], BLIP2 [blip-2], and Cambrian [tong2024cambrian], enable understanding and reasoning across visual and textual modalities for various multimodal tasks.
+A key component in advancing VLMs is visual instruction tuning (VIT) [llava], a training process that enables these models to interpret and follow instructions within a vision-language context, transforming them into versatile multimodal assistants.
+This tuning process not only improves the models’ instruction-following abilities but also aligns their outputs more closely with user expectations, thus enhancing their utility in practical applications [llava].
+To support this paradigm, recent VIT datasets have scaled to contain hundreds of thousands to millions of instruction-response pairs [liu2024improved, tong2024cambrian].
+Data selection.
+Data selection methods [survey] can be categorized based on the types of information they utilize for selection.
+Representation-based approaches [semdedup, coincide] leverage embeddings to capture data representations.
+Loss trajectory-based methods [mindermann2022prioritized] prioritize data points that contribute most significantly to reducing generalization error over training.
+Gradient-based techniques [el2n, xia2024less, deng2024influential] select data based on gradient information.
+Recent work has explored various approaches to select optimal visual instruction tuning datasets.
+Concurrent work TIVE [liu2024less] employs gradient-based selection to identify representative instances.
+TIVE assumes that the number of specialist data should be proportional to task difficulty and thus samples specialist data based on an estimation of task difficulty.
+Our method does not rely on this assumption – we directly select samples that benefit the greatest number of tasks.
+COINCIDE [coincide] clusters data based on representations associated with concept-skill compositions.
+Our work follows targeted instruction tuning selection approach LESS [xia2024less] to utilize gradient information to calculate the specialist influence (i.e., the influence on a specific task) and extends it to general scenarios by aggregating information across tasks and selecting data influential for multiple downstream tasks via majority voting.
+3 ICONS: Influence CONsensus for vision-language data Selection
 We propose a consensus-driven, gradient-based data selection framework (Fig.
@@
 Figure 2: ICONS.
-The Specialist stage (left) processes each task individually through three steps: (1) warmup training on a small subset of data, (2) gradient computation for both training and target task validation data, and (3) influence matrix computation to generate per-task influence scores.
-This process is repeated for each target task.
-The Generalist stage (right) performs Influence Consensus to aggregate information across tasks, where samples scoring above the $80^{\text{th}}$ percentile threshold for each task receive a vote.
-The final selection is made by summing votes across tasks and selecting the top 20% most influential samples, creating a compact yet highly effective training dataset that performs well across all tasks.
+The Specialist stage (left) processes each task through: (1) warmup training on a small subset, (2) gradient computation for training and validation data, and (3) influence matrix computation for per-task scores.
+The Generalist stage (right) performs Influence Consensus to aggregate information across tasks.
+For each task, samples scoring above a percentile threshold (e.g., $80^{\text{th}}$ for 20% selection) receive a vote.
+Votes are summed across tasks, and samples with the highest vote counts are selected, creating a compact yet effective training dataset that performs well across all tasks.
 3.1 Problem formulation
-Given a large-scale visual instruction tuning dataset $\mathcal{D}=\{({\bm{I}}_{i},{\bm{x}}_{i},{\bm{y}}_{i})\}_{i=1}^{N}$ containing $N$ samples, where each data point ${\bm{z}}_{i}=({\bm{I}}_{i},{\bm{x}}_{i},{\bm{y}}_{i})$ includes an image ${\bm{I}}_{i}$, natural language instruction ${\bm{x}}_{i}$, and corresponding target response ${\bm{y}}_{i}$ [^1] The framework supports multi-turn conversational data, yet we formalize the problem setup for single-turn instruction-tuning for clarity and simplicity., and given access to validation data $\mathcal{V}_{k}$ for each downstream task $T_{k}\in\mathcal{T}=\{T_{1},...,T_{K}\}$, our goal is to select a compact subset $\mathcal{S}\subset\mathcal{D}$ of size $M\ll N$ that maximizes model performance across multiple downstream tasks:
-$\mathcal{S}^{*}=\operatorname*{arg\,max}_{\mathcal{S}\subset\mathcal{D},|\mathcal{S}|=M}\sum_{k=1}^{K}\text{Rel}(f_{\mathcal{S}},T_{k}),\quad\text{Rel}(f_{\mathcal{S}},T_{k})=\frac{\text{Score}(f_{\mathcal{S}},T_{k})}{\text{Score}(f_{\mathcal{D}},T_{k})},$ (1)
-where $f_{\mathcal{S}}$ and $f_{\mathcal{D}}$ denote models trained on subset $\mathcal{S}$ and full dataset $\mathcal{D}$, respectively.
-$\text{Score}(f,T_{k})$ is the task-specific evaluation score achieved by model $f$ on task $T_{k}$.
+Given a large-scale visual instruction tuning dataset $\mathcal{D}=\{({\bm{I}}_{i},{\bm{x}}_{i},{\bm{y}}_{i})\}_{i=1}^{N}$ containing $N$ samples, where each data point ${\bm{z}}_{i}=({\bm{I}}_{i},{\bm{x}}_{i},{\bm{y}}_{i})$ includes an image ${\bm{I}}_{i}$, natural language instruction ${\bm{x}}_{i}$, and corresponding target response ${\bm{y}}_{i}$ [^1] This framework supports multi-turn conversational data, yet we formalize the setup for single-turn instruction-tuning for clarity and simplicity., and given access to validation data $\mathcal{V}_{k}$ for each downstream task $T_{k}\in\mathcal{T}=\{T_{1},...,T_{K}\}$, our goal is to select a compact subset $\mathcal{S}\subset\mathcal{D}$ of size $M\ll N$ that maximizes model performance across multiple downstream tasks.
+Concretely, we let $f_{\mathcal{S}}$ and $f_{\mathcal{D}}$ denote models trained on subset $\mathcal{S}$ and full dataset $\mathcal{D}$, respectively, and $\text{Score}(f,T_{k})$ be the task-specific evaluation score achieved by model $f$ on task $T_{k}$.
+$\text{Rel}(f_{\mathcal{S}},T_{k})=\frac{\text{Score}(f_{\mathcal{S}},T_{k})}{\text{Score}(f_{\mathcal{D}},T_{k})}$ quantifies the subset-trained model’s performance relative to that of the model trained on the entire dataset, with values close to 1 indicating that the subset maintains the performance of full training [coincide].
+Our goal is then to identify the subset:
+$\mathcal{S}^{*}=\operatorname*{arg\,max}_{\mathcal{S}\subset\mathcal{D},|\mathcal{S}|=M}\sum_{k=1}^{K}\text{Rel}(f_{\mathcal{S}},T_{k}).$ (1)
 We define the average relative performance across all tasks as Rel.
-= $\sum_{k=1}^{K}\text{Rel}(f_{\mathcal{S}},T_{k})/K$.
-Rel.
-quantifies the subset-trained model’s performance relative to that of the model trained on the entire dataset, with values close to 1 indicating that the subset maintains the performance of full training coincide .
-Our objective is to select a subset where Rel.
+= $\sum_{k=1}^{K}\text{Rel}(f_{\mathcal{S}},T_{k})/K$, and our objective is to select a subset where Rel.
 $\approx 1$, i.e., the model trained on the subset achieves comparable performance to the one trained with full dataset.
@@
 Denote a training data point as $\bm{z}$} and a validation data point as $\bm{z}$$^{\prime}$ from validation set $\mathcal{V}_{k}$ for task $T_{k}$.
-Following pruthi2020estimating ; xia2024less , we estimate how ${\bm{z}}\in\mathcal{D}$ affects validation loss by measuring its gradient alignment with reducing validation loss on $\mathcal{V}_{k}$, which directly impacts task-specific evaluation.
+Following [pruthi2020estimating, xia2024less], we estimate how ${\bm{z}}\in\mathcal{D}$ affects validation loss by measuring its gradient alignment with reducing validation loss on $\mathcal{V}_{k}$, which directly impacts task-specific evaluation.
 When training with SGD and batch size 1, using data point $\bm{z}$} at timestep $t$ leads to a model update ${\bm{\theta}}_{t+1}={\bm{\theta}}_{t}-\eta_{t}\nabla\ell({\bm{z}};{\bm{\theta}}_{t})$, where $\eta_{t}$ is the learning rate.
@@
 $\ell({\bm{z}}^{\prime};{\bm{\theta}}_{t+1})\approx\ell({\bm{z}}^{\prime};{\bm{\theta}}_{t})+\langle\nabla\ell({\bm{z}}^{\prime};{\bm{\theta}}_{t}),{\bm{\theta}}_{t+1}-{\bm{\theta}}_{t}\rangle.$
-The influence of a training data point ${\bm{z}}$ on a validation data point ${\bm{z}}^{\prime}$ is:
-$\mathcal{I}_{t}({\bm{z}}\rightarrow{\bm{z}}^{\prime})$ $=\ell({\bm{z}}^{\prime};{\bm{\theta}}_{t+1})-\ell({\bm{z}}^{\prime};{\bm{\theta}}_{t})\approx-\eta_{t}\langle\nabla\ell({\bm{z}}^{\prime};{\bm{\theta}}_{t}),\nabla\ell({\bm{z}};{\bm{\theta}}_{t})\rangle,$
+Training sample $\bm{z}$}’s influence on validation sample $\bm{z}$$^{\prime}$ is:
+$\mathcal{I}_{t}({\bm{z}}\rightarrow{\bm{z}}^{\prime})$ $=\ell({\bm{z}}^{\prime};{\bm{\theta}}_{t+1})-\ell({\bm{z}}^{\prime};{\bm{\theta}}_{t})$
+$\approx-\eta_{t}\langle\nabla\ell({\bm{z}}^{\prime};{\bm{\theta}}_{t}),\nabla\ell({\bm{z}};{\bm{\theta}}_{t})\rangle,$
 which we refer to as an influence score.
-The gradient-based selection approach selects training samples $\{{\bm{z}}\}$ that maximize the gradient inner product $\langle\nabla\ell({\bm{z}}^{\prime};{\bm{\theta}}_{t}),\nabla\ell({\bm{z}};{\bm{\theta}}_{t})\rangle$ [^2] In practice, we use cosine similarity instead of direct inner products to avoid biasing selection toward shorter sequences, since gradient norms tend to be inversely correlated with sequence length as noted in xia2024less .
+The gradient-based selection approach selects training samples $\{{\bm{z}}\}$ that maximize the gradient inner product $\langle\nabla\ell({\bm{z}}^{\prime};{\bm{\theta}}_{t}),\nabla\ell({\bm{z}};{\bm{\theta}}_{t})\rangle$ [^2] In practice, we use cosine similarity instead of direct inner products to avoid biasing selection toward shorter sequences, since gradient norms tend to be inversely correlated with sequence length as noted in [xia2024less].
 through a greedy, first-order approximation, which leads to larger reductions in validation loss for point $\bm{z}$$^{\prime}$.
-While it omits second-order terms compared to influence functions koh2017understanding , it provides an efficient approximation to rank the impact of training samples xia2024less .
+While it omits second-order terms compared to influence functions [koh2017understanding], it provides an efficient approximation to rank the impact of training samples [xia2024less].
 3.3 Specialist: individual task influence ranking
 To rank the influence of training data for each target task, we compute the influence score of each training data point on a validation set that represents the target task distribution.
-Following LESS xia2024less , the process involves three steps: (1) training the model on 5% randomly selected data as a lightweight warm-up to initialize visual instruction-following capabilities, (2) computing gradients for training and validation data and compressing the gradients via random projection, and (3) computing the influence score to quantify the impact of each training data on validation set.
+Following LESS [xia2024less], the process involves three steps: (1) training the model on 5% randomly selected data as a lightweight warm-up to initialize visual instruction-following capabilities, (2) computing gradients for training and validation data and compressing the gradients via random projection, and (3) computing the influence score to quantify the impact of each training data on validation set.
 Step 1: Warm-up Training.
-Following LESS xia2024less , we first apply LoRA hu2021lora on a small random subset $\mathcal{D}_{\mathrm{warmup}}\subset\mathcal{D}$ (5%) to obtain $f{\text{warmup}}=\text{LoRA}(f_{\text{base}},\mathcal{D}_{\text{warmup}})$.
-This allows the model to develop basic visual instruction-following capabilities.
+Following LESS [xia2024less], we apply LoRA [hu2021lora] on a small random subset $\mathcal{D}_{\mathrm{warmup}}\subset\mathcal{D}$ (5%) to obtain $f_{\text{warmup}}=\text{LoRA}(f_{\text{base}},\mathcal{D}_{\text{warmup}})$, which develops basic visual instruction-following capabilities.
 Step 2: Gradient computation.
@@
 where $y_{i}$ and $y^{\prime}_{j}$ are the targets for $z_{i}$ and $z^{\prime}_{j}$, respectively.
-In order to reduce computational and storage overhead, we apply random projection to the gradient feature: $\tilde{g}_{i}=Rg_{i}$ and $\tilde{g}^{\prime}_{j}=Rg^{\prime}_{j}$, where $R\in\mathbb{R}^{d^{\prime}\times d}$ is a random projection matrix with $d^{\prime}\ll d$ that preserves inner products with high probability (johnson1984extensions, ).
-We further normalize the projected gradients, $\tilde{g}_{i}=\frac{\tilde{g}_{i}}{\|\tilde{g}_{i}\|_{2}},\tilde{g}^{\prime}_{j}=\frac{\tilde{g}^{\prime}_{j}}{\|\tilde{g}^{\prime}_{j}\|_{2}}$ to prevent bias from sequence length differences xia2024less .
+In order to reduce computational and storage overhead, we apply random projection to the gradient feature: $\tilde{g}_{i}=Rg_{i}$ and $\tilde{g}^{\prime}_{j}=Rg^{\prime}_{j}$, where $R\in\mathbb{R}^{d^{\prime}\times d}$ is a random projection matrix with $d^{\prime}\ll d$ that preserves inner products with high probability [johnson1984extensions].
+We further normalize the projected gradients, $\tilde{g}_{i}=\frac{\tilde{g}_{i}}{\|\tilde{g}_{i}\|_{2}},\tilde{g}^{\prime}_{j}=\frac{\tilde{g}^{\prime}_{j}}{\|\tilde{g}^{\prime}_{j}\|_{2}}$ to prevent bias from sequence length differences [xia2024less].
 Step 3: Influence matrix computation.
-We compute the influence matrix $I\in\mathbb{R}^{|\mathcal{D}|\times|\mathcal{D}_{\text{val}}^{k}|}$ where each entry $I_{ij}=\langle\tilde{g}_{i},\tilde{g}^{\prime}_{j}\rangle$ is the influence of the training data ${\bm{z}}_{i}$ on the validation data ${\bm{z}}^{\prime}_{j}$, and then the average influence of training data ${\bm{z}}_{i}$ on the target task $k$ is calculated as:
+We compute the influence matrix $I\in\mathbb{R}^{|\mathcal{D}|\times|\mathcal{D}_{\text{val}}^{k}|}$ where each entry $I_{ij}=\langle\tilde{g}_{i},\tilde{g}^{\prime}_{j}\rangle$ is the influence of the training data ${\bm{z}}_{i}$ on the validation data ${\bm{z}}^{\prime}_{j}$, following the setting in LESS [xia2024less], we take the average influence of training data ${\bm{z}}_{i}$ across all validation examples on the target task $k$ to obtain a representative task-level influence score:
 $\bar{I}_{k}({\bm{z}}_{i})=\frac{1}{|\mathcal{D}_{\text{val}}^{k}|}\sum_{j=1}^{|\mathcal{D}_{\text{val}}^{k}|}I_{ij}.$ (2)
-This influence estimation process provides a task-specific ranking for the training set $\mathcal{D}$ with respect to task $\mathcal{T}_{k}$, where a higher influence score $\bar{I}_{i}$ suggests a higher influence $\mathcal{T}_{k}$.
+This influence estimation process provides a task-specific ranking for the training set $\mathcal{D}$ with respect to task $\mathcal{T}_{k}$, where a higher influence score $\bar{I}_{i}$ suggests a higher influence on $\mathcal{T}_{k}$.
 We can select a small training subset $\mathcal{S}_{k}$ for a given task $k$ by selecting the training examples ${\bm{z}}_{i}$ with the highest-influence values $\bar{I}_{k}({\bm{z}}_{i})$.
@@
 $I_{\text{Merge}}({\bm{z}}_{i})=\sum_{k=1}^{K}\bar{I}_{k}({\bm{z}}_{i}).$ (3)
-A similar aggregation approach is the one suggested in LESS xia2024less :
+LESS [xia2024less] suggests a similar approach:
 $I_{\text{Max}}({\bm{z}}_{i})=\max_{k=1,\dots K}\bar{I}_{k}({\bm{z}}_{i}),$ (4)
@@
 We can have a couple of options.
-First, we can select the training subset either using the Round Robin (RR) approach ivison2025large where we iterate over tasks and select the lowest-rank example which has not yet been selected to add to our set $\mathcal{S}_{\text{RR}}$.
+First, we can select the training subset either using the Round Robin (RR) approach [ivison2025large] where we iterate over tasks and select the lowest-rank example which has not yet been selected to add to our set $\mathcal{S}_{\text{RR}}$.
 Alternatively, we can select the training subset $\mathcal{S}_{\text{MinRank}}$ such that all the examples within it have a low rank for some task $k$.
@@
 i.e., all examples that are not included in $\mathcal{S}$ would have high relative ranks $\text{rank}_{k}({\bm{z}}_{i})$ for all tasks $k$.
-However, this approach does not consider the potential interplay between tasks.
-Recall that in Eqn.
-1 we aim to maximize sum of the relative performance across all tasks $k$; thus, if a training example is beneficial for multiple tasks, we may want to include it even at the expense of a lower-ranked example for a different task $k$.
-Thus, we introduce a simple consensus-based voting strategy that identifies training samples that consistently show a high influence score across various tasks.
-Concretely, we leverage the specialist training sets $\mathcal{S}_{k}$ as defined in §2.3 consisting of the $M$ highest-influence training examples for each task.
-We then select a combined training set as follows:
-$S_{\text{ICONS}}=\operatorname*{arg\,max}_{\mathcal{S}\subset\mathcal{D},|\mathcal{S}|=M}\sum_{k=1}^{K}\sum_{{\bm{z}}_{i}\in S}\mathbbm{1}[{\bm{z}}_{i}\in S_{k}]$ (6)
+However, this approach ignores the potential interplay between tasks.
+Since Eqn.
+1 maximizes the sum of relative performance across tasks $k$, if a training example is beneficial for multiple tasks, we may want to include it even at the expense of lower-ranked examples for individual tasks.
+Thus, we introduce a consensus-based voting strategy that identifies training samples consistently showing high influence across tasks.
+For each task $k$, we set a threshold $\tau_{k}$ at the $((1-p)\times 100)$-th percentile of the influence score distribution, where $p$ represents the selection ratio ($p=0.2$ in our main experiments).
+A sample ${\bm{z}}_{i}$ is considered important for task $k$ if $\bar{I}_{k}({\bm{z}}_{i})\geq\tau_{k}$.
+We define the voting-based influence score as:
+$I_{\text{vote}}({\bm{z}}_{i})=\sum_{k=1}^{K}\mathbbm{1}[\bar{I}_{k}({\bm{z}}_{i})\geq\tau_{k}],$ (6)
+where $\mathbbm{1}[\cdot]$ is the indicator function.
+The final score ranges from 0 (no importance for any task) to $K$ (consistent importance for all tasks).
+Equivalently, we can define the threshold-based specialist sets as: $\mathcal{S}_{k}=\{{\bm{z}}_{i}\in\mathcal{D}:\bar{I}_{k}({\bm{z}}_{i})\geq\tau_{k}\}$.
+We then select the combined training set as:
+$\mathcal{S}_{\text{ICONS}}$ $=\operatorname*{arg\,max}_{\mathcal{S}\subset\mathcal{D},|\mathcal{S}|=M}\sum_{k=1}^{K}\sum_{{\bm{z}}_{i}\in S}\mathbbm{1}[{\bm{z}}_{i}\in S_{k}]$ (7)
+$=\operatorname*{arg\,max}_{\mathcal{S}\subset\mathcal{D},|\mathcal{S}|=M}\sum_{{\bm{z}}_{i}\in S}I_{\text{vote}}({\bm{z}}_{i})$ (8)
 This simple approach offers a key advantage: it does not rely on calibration of influence scores across tasks, and does not make any a-priori assumptions about the relationship between tasks (e.g., that every task needs to have its highest-scoring examples included in the combined training set).
@@
 4 Experiments
-In this section, we first discuss our experiment setup and evaluation benchmarks (§3.1).
-We then present our main results by comparing ICONS with the state-of-the-art methods (§3.2), followed by analysis of different selection strategies (§3.3).
-We further evaluate the transferability of ICONS3.4).
-Lastly, we provide analyze performance trends under different selection ratios (§3.5).
+We first discuss our experiment setup (§4.1), then compare ICONS with state-of-the-art methods (§4.2) and analyze different selection strategies (§4.3).
+We evaluate transferability to unseen tasks (§4.4), performance across selection ratios (§4.5), vote distributions (§4.6), and characteristics of low-vote samples (§4.7).
 4.1 Evaluation test-bed
 Datasets & model.
-We apply ICONS on major visual instruction tuning (VIT) training datasets: LLaVA-665K liu2024improved , Cambrian-7M tong2024cambrian and Vision-Flan-186K visionflan .
+We apply ICONS on major VIT training datasets: LLaVA-665K [liu2024improved], Cambrian-7M [tong2024cambrian] and Vision-Flan-186K [visionflan].
 The majority of our analysis and ablation experiments are conducted on LLaVA-665K.
-For our experiments, we use the LLaVA-v1.5 model (liu2024improved, ) checkpoint after Stage 1 (pre-training for feature alignment) as defined in the original LLaVA training pipeline, with a default size of 7B parameters and LLaVA-665K unless otherwise specified.
+For our experiments, we use the LLaVA-v1.5 model [liu2024improved] checkpoint after Stage 1 (pre-training for feature alignment) as defined in the original LLaVA training pipeline, with a default size of 7B parameters and LLaVA-665K unless otherwise specified.
 This checkpoint333llava-v1.5-mlp2x-336px-pretrain-vicuna-7b-v1.5, which has no prior exposure to the visual instruction tuning data.
@@
 In all experiments, we train the models for one epoch following the official finetuning hyperparameters using LoRA.
-More details on computation, including hardware specifications and runtime are in Appendix A.
+More details on computation, including hardware specifications and runtime, are in Appendix §A.
 Target tasks.
-We evaluate ICONS across diverse multimodal benchmarks (Appendix C, Tab.
-5) that test different capabilities of vision-language models: 1) Multiple-choice understanding: MMBench (mmbench, ) and MME (mme, )  [^4] For MME, we focus on its perception section following coincide , which evaluates vision capablities.
-2) Visual question answering: VQAv2 (vqav2, ), GQA (gqa, ), and VizWiz (vizwiz, ); 3) Text understanding in images: TextVQA (textvqa, ); 4) Scientific reasoning: ScienceQA (sqa, ); 5) Open-ended generation: LLaVA-W Bench (llava, ); 6) Factual consistency: POPE (pope, ).
+We evaluate ICONS across diverse multimodal benchmarks (details are in Appendix §B, Tab.
+6) that test different capabilities of vision-language models: 1) Multiple-choice understanding: MMBench [mmbench] and MME [mme]  [^4] For MME, we focus on its perception section following [coincide], which evaluates vision capabilities.
+2) Visual question answering: VQAv2 [vqav2], GQA [gqa], and VizWiz [vizwiz]; 3) Text understanding in images: TextVQA [textvqa]; 4) Scientific reasoning: ScienceQA [sqa]; 5) Open-ended generation: LLaVA-W Bench [llava]; 6) Factual consistency: POPE [pope].
 Baselines.
-We compare our ICONS against several baselines, including random selection, CLIP-Score clip for measuring image-text alignment, EL2N el2n based on embedding L2 norms, and Perplexity perplexity using language model scores.
-We also compare against SemDeDup semdedup for semantic deduplication and D2-Pruning (d2, ) for distribution-aware pruning.
-Additional baselines include Self-Sup (selfsup, ) leveraging self-supervised signals, while Self-Filter (self-filter, ) and COINCIDE (coincide, ) are designed for vision-language data selection.
-We reference LLaVA-665K baseline results from COINCIDE coincide .
-Additionally, we compare with representation-based data selection baseline (RDS) ivison2025large ; xia2024less .
-Method VQAv2 GQA VizWiz SQA-I TextVQA POPE MME MMBench LLaVA-W Rel. (%)
-en cn Bench
-Full 79.1 63.0 47.8 68.4 58.2 86.4 1476.9 66.1 58.9 67.9 100
-Random 75.7 58.9 44.3 68.5 55.3 84.7 1483.0 62.2 54.8 65.0 95.8
-CLIP-Score clip 73.4 51.4 43.0 65.0 54.7 85.3 1331.6 55.2 52.0 66.2 91.2
-EL2N el2n 76.2 58.7 43.7 65.5 53.0 84.3 1439.5 53.2 47.4 64.9 92.0
-Perplexity perplexity 75.8 57.0 47.8 65.1 52.8 82.6 1341.4 52.0 45.8 68.3 91.6
-SemDeDup semdedup 74.2 54.5 46.9 65.8 55.5 84.7 1376.9 52.2 48.5 70.0 92.6
-D2-Pruning d2 73.0 58.4 41.9 69.3 51.8 85.7 1391.2 65.7 57.6 63.9 94.8
-Self-Sup selfsup 74.9 59.5 46.0 67.8 49.3 83.5 1335.9 61.4 53.8 63.3 93.4
-Self-Filter self-filter 73.7 58.3 53.2 61.4 52.9 83.8 1306.2 48.8 45.3 64.9 90.9
-COINCIDE coincide 76.5 59.8 46.8 69.2 55.6 86.1 1495.6 63.1 54.5 67.3 97.4
-RDS ivison2025large ; xia2024less 75.1 57.9 48.6 68.0 54.9 86.3 1393.8 61.2 52.7 63.7 95.2
-ICONS (ours) 76.3 60.7 50.1 70.8 55.6 87.5 1485.7 63.1 55.8 66.1 98.6
+We compare our ICONS against several baselines, including random selection, CLIP-Score [clip] for measuring image-text alignment, EL2N [el2n] based on embedding L2 norms, and Perplexity [perplexity] using language model scores.
+We also compare against SemDeDup [semdedup] for semantic deduplication and D2-Pruning [d2] for distribution-aware pruning.
+Additional baselines include Self-Sup [selfsup] leveraging self-supervised signals, while Self-Filter [self-filter] and COINCIDE [coincide] are designed for vision-language data selection.
+We reference LLaVA-665K baseline results from COINCIDE [coincide].
+Additionally, we compare with representation-based data selection approach (RDS) [ivison2025large, xia2024less].
 Table 1: Selection results on LLaVA-665K.
 Performance comparison of different data selection approaches when trained on 20% of the LLaVA-665K dataset.
+Tasks are ordered by random selection’s relative performance.
 The best and second best results for each benchmark are shown in bold and underlined, respectively.
 Our method ICONS achieves the highest overall Rel.
-(98.6%), consistently outperforming existing approaches including COINCIDE coincide (97.4%) and D2-Pruning d2 (94.8%).
+(98.6%), consistently outperforming existing approaches including COINCIDE [coincide] (97.4%) and D2-Pruning [d2] (94.8%).
+Method MME SQA-I POPE VQAv2 LLaVA-W TextVQA MMBench GQA VizWiz Rel. (%)
+Bench en cn
+Full 1476.9 68.4 86.4 79.1 67.9 58.2 66.1 58.9 63.0 47.8 100
+Random 1483.0 68.5 84.7 75.7 65.0 55.3 62.2 54.8 58.9 44.3 95.8
+CLIP-Score [clip] 1331.6 65.0 85.3 73.4 66.2 54.7 55.2 52.0 51.4 43.0 91.2
+EL2N [el2n] 1439.5 65.5 84.3 76.2 64.9 53.0 53.2 47.4 58.7 43.7 92.0
+Perplexity [perplexity] 1341.4 65.1 82.6 75.8 68.3 52.8 52.0 45.8 57.0 47.8 91.6
+SemDeDup [semdedup] 1376.9 65.8 84.7 74.2 70.0 55.5 52.2 48.5 54.5 46.9 92.6
+D2-Pruning [d2] 1391.2 69.3 85.7 73.0 63.9 51.8 65.7 57.6 58.4 41.9 94.8
+Self-Sup [selfsup] 1335.9 67.8 83.5 74.9 63.3 49.3 61.4 53.8 59.5 46.0 93.4
+Self-Filter [self-filter] 1306.2 61.4 83.8 73.7 64.9 52.9 48.8 45.3 58.3 53.2 90.9
+COINCIDE [coincide] 1495.6 69.2 86.1 76.5 67.3 55.6 63.1 54.5 59.8 46.8 97.4
+RDS [ivison2025large, xia2024less] 1393.8 68.0 86.3 75.1 63.7 54.9 61.2 52.7 57.9 48.6 95.2
+[…]
 4.2 Main results
@@
 Approaches like EL2N, Perplexity, SemDeDup achieve only 91-92% Rel., showing limitations in preserving performance.
+Table 2: Selection results on Vision-Flan-186K and Cambrian-7M.
+Performance comparison of different data selection approaches when trained on 20% of the Vision-Flan-186K [visionflan] and Cambrian-7M [tong2024cambrian] datasets.
+ICONS achieves strong performance (99.8% and 98.8% Rel.
+respectively) while using only 20% of the training data, significantly outperforming random selection which is one of the strongest baselines, and approaching full performance.
 Dataset #Data Method VQAv2 GQA VizWiz SQA-I TextVQA POPE MME MMBench LLaVA-W Rel. (%)
@@
 37k Random 64.1 45.8 37.5 58.7 45.3 82.9 1,079.8 46.5 39.6 58.7 91.6
-ICONS (ours) 67.4 48.8 44.1 60.2 49.9 83.0 1,252.5 51.9 45.4 63.1 99.8
+[…]
 Cambrian-7M 7,068k Full 80.2 62.9 58.4 75.3 60.9 86.5 1,524.6 69.1 58.9 67.6 100.0
 1,414k Random 74.2 57.5 61.9 71.0 57.1 86.4 1,465.7 63.3 49.6 70.4 95.4
-ICONS (ours) 79.6 62.1 60.7 73.9 59.8 86.2 1,503.1 67.8 55.8 67.0 98.8
-Table 2: Selection results on Vision-Flan-186K and Cambrian-7M.
-Performance comparison of different data selection approaches when trained on 20% of the Vision-Flan-186K visionflan and Cambrian-7M tong2024cambrian datasets.
-ICONS achieves strong performance (99.8% and 98.8% Rel.
-respectively) while using only 20% of the training data, significantly outperforming random selection which is one of the strongest baselines, and approaching full performance.
+[…]
 Cambrian-7M & Vision-Flan-186K selection.
-We further provide results on Vision-Flan-186K and Cambrian-7M in Tab.
+We provide results on Vision-Flan-186K [visionflan] and Cambrian-7M [tong2024cambrian] in Tab.
 2.
-On Vision-Flan-186K, our method achieves near-full performance (99.8% Rel.) using just 37k samples, significantly outperforming random selection (91.6%) across tasks.
-Similarly, on Cambrian-7M, ICONS maintains a strong performance (98.8% Rel.) with 1,414k samples, while random selection achieves 95.4%.
-These results demonstrate that our approach scales effectively to both small and large datasets, consistently preserving model capabilities while drastically reducing the training data required.
+On Vision-Flan-186K, our method achieves near-full performance (99.8% Rel.) using just 37k samples, significantly outperforming random selection (91.6%).
+Similarly, on Cambrian-7M, ICONS maintains strong performance (98.8% Rel.) with 1,414k samples, while random selection achieves 95.4%.
+These results demonstrate our approach works effectively on both small and large datasets, consistently preserving model capabilities while drastically reducing training data.
 Comparisons with representation-based data selection.
-We compare our method against RDS (Representation-based Data Selection)ivison2025large ; xia2024less , a strong baseline in language-only instruction tuning.
+We compare against RDS (Representation-based Data Selection) [ivison2025large, xia2024less], a strong baseline in language-only data selection.
 RDS computes training-validation similarity using final-layer representation of the last token in each sequence instead of gradients.
-For a fair comparison, we use the same influence matrix formulation (Eqn.2) and apply majority voting to reach influence consensus.
+For a fair comparison, we use the same influence matrix formulation (Eqn.
+2) and apply majority voting.
 Our method consistently outperforms RDS across all tasks, particularly those requiring perceptual grounding – e.g., higher scores on GQA (60.7 vs.
@@
 1393.8).
-While RDS is effective in selecting large-scale text-only data (e.g.
-TULU-2/3 ivison2023camels ; lambert2024t ), its evaluation has largely focused on language-only tasks, where semantic similarity alone is often sufficient.
-In contrast, vision-language tasks demand alignment between modalities, where representation similarity is limited as it only reflects current embedding proximity and gradient-based approaches directly estimate each sample’s contribution to the validation loss.
-Our gradient-based approach directly estimates each sample’s impact on validation loss, capturing cross-modal training dynamics and prioritizing impactful training points.
-We further provide qualitative comparisons in Appendix §I.
+While RDS is effective for text-only data (e.g.
+TULU-2/3 [ivison2023camels, lambert2024t]), vision-language tasks demand alignment between modalities, where representation similarity is limited as it only reflects current embedding proximity, whereas gradient-based approaches directly estimate each sample’s contribution to validation loss, better capturing cross-modal training dynamics.
+We further provide qualitative comparisons in Appendix §F.
 4.3 Analysis of aggregation strategies
@@
 As introduced in §3.4, we explore multiple strategies for aggregating task-specific influence rankings into a single compact subset.
-We compare our majority voting approach (Vote) with different aggregation approaches for combining task-specific influence scores: 1) score-based methods (Merge, Max) and their normalized variants (Merge-GausNorm, Merge-SumNorm), 2) rank-based methods (Round Robin, MinRank).
-Our voting-based strategy outperforms all alternatives (Tab.
-3), achieving the highest overall Rel.
-(98.6%).
-This shows that building a cross-task consensus via majority voting is a simple yet effective strategy for identifying consistently influential examples across tasks, without assuming calibration or comparability of scores.
-Aggregation VQAv2 GQA VizWiz SQA-I TextVQA POPE MME MMB (en) MMB (cn) LLaVA-W Rel. (%)
-Full 79.1 63.0 47.8 68.4 58.2 86.4 1476.9 66.1 58.9 67.9 100
-Merge 75.7 59.6 47.9 65.5 55.5 86.0 1422.1 59.0 51.0 66.2 96.4
-Max 75.2 59.8 48.1 66.2 55.5 85.5 1470.7 58.3 51.8 66.2 96.1
-Merge-GausNorm 75.1 60.1 46.4 69.8 54.5 85.6 1482.6 58.9 52.5 66.3 96.8
-Merge-SumNorm 75.5 59.1 51.7 68.7 43.5 87.1 1478.3 59.5 50.9 69.8 95.3
-Round Robin 75.4 59.1 48.3 70.6 55.2 86.6 1474.5 61.6 51.5 66.9 96.7
-MinRank 75.2 59.0 49.7 70.4 55.1 86.9 1456.3 61.1 52.4 68.4 97.1
-Vote (ours) 76.3 60.7 50.1 70.8 55.6 87.5 1485.7 63.1 55.8 66.1 98.6
+We compare our majority voting approach (Vote) with score-based methods (Merge, Max, and normalized variants (Merge-SumNorm, Merge-GausNorm)) and rank-based methods (Round Robin, MinRank).
+Our approach outperforms all alternatives (Tab.
+3; full results in Appendix Tab.
+5), achieving 98.6% Rel.
+Our majority voting effectively identifies consistently influential examples without requiring score calibration across tasks.
 Table 3: Comparison of aggregation approaches.
-Performance of different influence aggregation methods when selecting 20% of the LLAVA-665K dataset.
-Our proposed aggregation approach (Vote) consistently achieves the best overall performance (98.6% Rel.), outperforming both score-based (Merge, Max), their noramlized variants (Merge-GausNorm, Merge-SumNorm) and rank-based (Round Robin, MinRank) baselines.
+Relative performance of different influence aggregation methods when selecting 20% of the LLAVA-665K dataset on the ten tasks from Tables 1 and 2.
+Our Vote achieves the best performance (98.6% Rel.), outperforming score-based, their normalized variants, and rank-based baselines.
+See Appendix Tab.
+5 for per-task results.
+Type Score-based Normalized Score-based
+Method Merge Max Merge-GausNorm Merge-SumNorm
+[…]
+Type Rank-based Vote-based
+Method Round Robin MinRank ICONS (ours)
+[…]
 Limitations of score-based aggregation.
@@
 4) assume calibrated influence scores across tasks, which is rarely the case.
-We observe substantial variation in the distribution of influence scores across tasks with standard deviations spanning from $8.15\times 10^{-3}$ (MME) to $1.26\times 10^{-3}$ (VQAv2), indicating that influence scores for MME are much more spread out, while those for VQAv2 are tightly concentrated.
-Similarly, mean influence scores vary in both magnitude and sign: MME has a relatively high positive mean ($1.68\times 10^{-3}$), while tasks like POPE ($-2.83\times 10^{-4}$) and GQA ($-8.89\times 10^{-5}$) have negative means.
-These divergent patterns show that certain tasks have much wider influence scores distributions, making a sample helpful for one task but neutral or harmful for another.
-Aggregating raw scores thus biases selection toward tasks with higher variance or skewed means.
-To address this calibration issue, we experimented with normalization strategies: Merge-SumNorm, which rescales each task’s influence scores by dividing them by a task-specific normalization factor (i.e., sum), and Merge-GausNorm, which normalizes the scores using task-wise mean and standard deviation before averaging:
-$I_{\text{Merge-SumNorm}}({\bm{z}}_{i})=\sum_{k=1}^{K}\frac{I_{k}({\bm{z}}_{i})}{\sum_{j}I_{k}({\bm{z}}_{j})}\qquad I_{\text{Merge-GausNorm}}({\bm{z}}_{i})=\sum_{k=1}^{K}\frac{I_{k}({\bm{z}}_{i})-\mu_{k}}{\sigma_{k}}$
+We observe substantial variation in influence score distributions across tasks with standard deviations spanning from $8.15\times 10^{-3}$ (MME) to $1.26\times 10^{-3}$ (VQAv2), indicating that influence scores for MME are much more spread out, while those for VQAv2 are tightly concentrated.
+Similarly, mean influence scores vary in both magnitude and sign: MME has a relatively high positive mean ($1.68\times 10^{-3}$), while tasks like POPE ($2.83\times 10^{-4}$) and GQA ($8.89\times 10^{-5}$) have negative means.
+These patterns show certain tasks have wider influence scores distributions, making a sample helpful for one task but neutral or harmful for another.
+Aggregating raw scores biases selection toward tasks with higher variance or skewed means.
+To address this calibration issue, we experiment with normalization: Merge-SumNorm, which rescales each task’s influence scores by dividing them by a task-specific normalization factor (i.e., sum), and Merge-GausNorm, which normalizes the scores using task-wise mean and standard deviation before averaging:
+$I_{\text{Merge-SumNorm}}({\bm{z}}_{i})=\sum_{k=1}^{K}\frac{I_{k}({\bm{z}}_{i})}{\sum_{j}I_{k}({\bm{z}}_{j})}$
+$I_{\text{Merge-GausNorm}}({\bm{z}}_{i})=\sum_{k=1}^{K}\frac{I_{k}({\bm{z}}_{i})-\mu_{k}}{\sigma_{k}}$
 However, as shown in Tab. 3, both methods still underperform compared to our voting-based strategy, reinforcing the limitations of relying on score magnitudes directly.
 Limitations of rank-based aggregation.
-Rank-based methods sidesteps the challenge of comparing raw influence scores by focusing on within-task ranking.
+Rank-based methods sidestep the challenge of comparing raw influence scores by focusing on within-task ranking.
 Round Robin selects samples by cycling through each task and picking the highest-scoring remaining sample for that task, ensuring balanced coverage.
@@
 Although these methods ensure balanced coverage across tasks, they can overfit to outlier tasks.
-This is particularly evident with LLaVA-W Bench llava , which is an outlier in its influence ranking: both Round Robin and MinRank achieve relatively high scores on it (e.g.
-MinRank: 68.4), but this comes at the cost of lower performance on all other tasks (Tab.
+This is particularly evident with LLaVA-W Bench [llava], which is an outlier in its influence ranking: both Round Robin and MinRank achieve relatively high scores on it (e.g.
+MinRank: 68.4) but at the cost of lower performance on all other tasks (Tab.
 3).
-This suggests that purely rank-based selection can trade off some overall efficacy on the mainstream tasks and hurt multi-task balance.
-In contrast, our Vote approach is more robust and avoids this by focusing on multi-task consensus rather than forcing equal representation, yielding better balance and higher overall performance (98.6% Rel.), highlighting the importance of identifying broadly influential examples rather than narrowly optimizing per-task rankings.
-Figure 3: Pairwise overlap heatmap between specialists. The values show overlap percentages between benchmarks’ selected samples.
-Figure 4: Data overlap between specialists and generalist selection.
-Overlap varies significantly, from 3.27% (VQAv2) to 24.21% (LLAVA-W Bench), reflecting varying alignment between task-specific and consensus selections.
-Divergent multi-task influence patterns.
+Rank-based selection can trade off overall efficacy on the mainstream tasks.
+In contrast, our Vote approach focuses on cross-task consensus rather than forcing equal representation, yielding better balance and higher overall performance (98.6% Rel.), showing the importance of identifying broadly influential examples rather than optimizing per-task rankings.
+Figure 3: Unseen task generalization across methods.
+Performance on eight unseen benchmarks using 20% selected subsets, sorted by the average performance across 5 random selection runs (left: lower, right: higher).
+Shaded region shows mean and standard deviation across random runs.
+Our ICONS consistently outperforms all baselines.
+Full detailed results are in Appendix Tab.
+7.
+4.4 ICONS generalizes to unseen tasks
+Data selected via ICONS demonstrate strong generalization on unseen benchmarks not used during data selection.
+In Fig.
+3 and Appendix Tab.
+7 we test LLaVA-ICONS-133K on diverse tasks including MMVet [mmvet], NaturalBench [naturalbench], AI2D [ai2d], ChartQA [chartqa], DocVQA [docvqa], InfoVQA [infographicvqa], RealWorldQA [xai2024grok] and CMMMU [cmmmu].
+Notably, LLaVA-ICONS-133K achieves substantial improvements on challenging benchmarks (MMVet: 95.5%, AI2D: 97.3%) and even exceeds full dataset performance on harder tasks (CMMMU: 114.0%, RealWorldQA: 104.4%, InfoVQA: 103.8%, NaturalBench: 103.2%), demonstrating superior generalization to entirely unseen task formats and domains despite these tasks not being included in the selection process.
+Importantly, LLaVA-ICONS-133K outperforms random selection across all benchmarks.
+This suggests that our selection approach successfully captures fundamental visual-language understanding capabilities that transfer well across different task formats and domains.
+We further provide the cross-architecture generalization results in Appendix §C.
+4.5 ICONS outperforms baselines across ratios and exceeds full-data training at 60%
+To understand how ICONS scales, we evaluate it across different selection ratios, progressively scaling the subset size from 5% to 90% of LLaVA-665K.
 As shown in Fig.
-4, the pairwise overlap heatmap shows notable variation in training data influence across tasks.
-High overlap – e.g., VQAv2 and VizWiz (49.0%) or POPE and GQA (60.2%), suggests that certain samples are beneficial across similar tasks.
-However, low overlap, like the 3.3% between MMBench (en) and GQA, highlights that highly influential samples for one task may have limited impact on others.
-Even closely related tasks, such as MMBench in different languages (English and Chinese), share 67.4% of influential samples.
-To understand task-specific influence matrices from the specialist stage, we select the top 20% samples per task (Specialists).
-Overlap with our generalist subset (Fig.
-4) varies significantly, from minimal in tasks like VQAv2 (3.27%) and VizWiz (3.28%) to substantial agreement in tasks like LLAVA-W Bench llava (24.21%).
-These findings empirically demonstrate significant overlap in influential samples across tasks and validate our approach: by analyzing task-specific gradient-based influence patterns and building consensus across tasks, we can identify a compact subset that captures broadly useful samples across tasks, yielding strong performance with significantly less data.
-3.4 ICONS generalizes to unseen tasks
-LLaVA-ICONS-133K demonstrates exceptional generalization on entirely unseen benchmarks that were not used during data selection.
-As shown in Tab.
-4, we test across a diverse spectrum of tasks including MMVet mmvet , NaturalBench naturalbench , AI2D ai2d , ChartQA chartqa , DocVQA docvqa , InfoVQA infographicvqa , RealWorldQA xai2024grok and CMMMU cmmmu .
-LLaVA-ICONS-133K achieves 95.5-113.9% (Rel.) compared to full dataset training (InfoVQA: 103.8%, NaturalBench: 105.5%, RealWorldQA: 104.4%, and CMMMU: 113.9%), suggesting that, in some cases, training on LLaVA-ICONS-133K may even outperform training on the full dataset, despite these tasks not being included in the selection process.
-Importantly, LLaVA-ICONS-133K significantly outperforms random selection across all benchmarks.
-This suggests that our selection approach successfully captures fundamental visual-language understanding capabilities that transfer well across different task formats and domains.
-We further provide the cross-architecture generalization results in Appendix §D.
-AI2D ChartQA DocVQA InfoVQA MMVet Naturalbench RealworldQA CMMMU Rel. (%)
-Full 55.4 17.5 28.9 26.5 31.1 12.4 52.4 22.1 100.0
-Random 50.2 15.1 25.2 24.3 27.6 11.1 49.8 21.9 91.6
-LLaVA-ICONS-133K 53.9 17.1 27.9 27.5 29.7 12.8 55.0 25.2 98.7
-Per-task Rel. (%) 97.3 97.7 96.5 103.8 95.5 103.2 104.4 114.0 -
-Table 4: Unseen-task generalization.
-Performance comparison on unseen benchmarks when trained on selected subsets.
-Notably, we observe improvements on InfoVQA (103.8%), RealWorldQA (104.4%), and CMMMU (114.0%), highlighting strong generalization to unseen tasks.
-3.5 ICONS outperforms baselines across ratios and exceeds full-data training at 60%
-Figure 5: Different selection ratios. ICONS consistently outperforms all baselines across different selection ratios and remarkably exceeding 102% at 60% selection ratio.
-To understand how ICONS scales, we evaluate it across different selection ratios, progressively scaling the subset size from 5% to 60% of LLaVA-665K.
-As shown in Fig.
-5, our results reveal several key patterns: First, ICONS shows particularly strong performance in the low-selection regime (5-20%), where identifying the most influential samples is crucial.
+4, our results reveal several key patterns: First, ICONS shows particularly strong performance in the low-selection regime (5-20%), where identifying the most influential samples is crucial.
 Second, as the selection ratio increases, the performance gap between different methods gradually narrows.
@@
 One hypothesis is that ICONS can also effectively filter out potentially harmful or noisy training samples that might negatively impact model training, thereby surpassing the full training performance.
+Figure 4: Different selection ratios. ICONS consistently outperforms all baselines across different selection ratios and remarkably exceeds 102% at a 60% selection ratio.
+4.6 How does voting change with selection ratio?
+We analyze vote distribution across selection ratios to understand task consensus patterns (Tab.
+4, Appendix Fig.
+6).
+At 5% selection, 72.1% of samples receive zero votes, indicating highly task-specific selection.
+As the ratio increases to 50% and 95%, the distribution shifts rightward with median votes rising to 4 and 5-7 respectively, and zero-vote samples decreasing to 8.3% and 1.2%.
+This rightward shift demonstrates that task consensus emerges as selection criteria relax, revealing samples with broader multi-task utility, while persistent zero-vote samples indicate universally low-quality or redundant data.
+These patterns validate our voting-based aggregation approach, which prioritizes samples with higher multi-task agreement.
+We further visualize zero-vote and one-vote samples in §4.7.
+Table 4: Vote distribution across selection ratios. Vote distributions shift from task-specific selection at low ratios to broader consensus at high ratios.
+[…]
+5% 0.50 $\pm$0.98 0 7 72.1% 3 votes
+20% 2.00 $\pm$1.56 2 7 28.4% 5 votes
+50% 5.00 $\pm$2.21 5 7 8.3% 6 votes
+90% 9.00 $\pm$1.84 9 10 1.2% 9 votes
+95% 9.50 $\pm$1.42 10 10 0.8% 10 votes
+4.7 What do low-vote training samples look like?
+To understand what samples are filtered out, we examine zero-vote and 1-vote samples at a 90% selection ratio for LLaVA-665K, where the majority of samples receive votes, making the remaining low-vote samples particularly indicative of low-quality data.
+Zero-vote samples do not rank in the top percentile for any task, while 1-vote samples are selected by only a single task-specific specialist.
+As shown in Appendix Fig.
+10, these low-vote samples exhibit characteristics that explain their low selection frequency: (1) subjective questions lacking objective ground truth (e.g., “Is the woman pretty?”), (2) trivial questions providing minimal visual learning signal (e.g., “Is the door open?” with answer “Yes”), and (3) redundant questions about the same image content.
+By filtering out such potentially harmful or redundant examples, ICONS focuses training on more informative and generalizable samples, and thus exceeds full-dataset performance.
 5 Conclusion
-In this work, we introduce ICONS, a simple yet effective influence consensus-based approach for visual instruction tuning data selection.
-By leveraging gradient-based influence estimation and aggregating task-specific selections through majority voting, our two-stage specialist-to-generalist approach selects training examples that are broadly beneficial across multiple downstream tasks.
-ICONS addresses limitations of prior selection methods by avoiding assumptions about score comparability across tasks and reducing the sensitivity to outlier task rankings, which can bias selection in both score-based and rank-based approaches.
+We introduce ICONS, an influence consensus-based approach for visual instruction tuning data selection.
+It leverages gradient-based influence estimation and aggregates task-level influence scores through majority voting to select training samples broadly beneficial across tasks.
+ICONS addresses limitations of prior approaches by avoiding assumptions about score comparability across tasks and reducing sensitivity to outlier task rankings that bias selection.
+ICONS builds compact, high-impact datasets achieving 98.6% of full performance using only 20% of LLaVA-665K while generalizing well to unseen tasks and architectures.
 Beyond data selection, it provides a principled way to reason about data influence in multitask mixtures.
-Through extensive experiments, we show that ICONS builds compact, high-impact datasets without sacrificing performance or generalization, achieving 98.6% of full dataset performance using only 20% of LLaVA-665K and generalizing well to unseen tasks or architectures.
-We release LLaVA-ICONS-133K and Cambrian-ICONS-1.4M, 20% subsets of the LLaVA-665K and Cambrian-7M datasets, maintaining strong performance on diverse tasks and transferring well to unseen ones.
-We hope our work inspires further exploration into data-efficient methods for vision-language models across diverse applications.
+We release LLaVA-ICONS-133K, Cambrian-ICONS-1.4M, and Vision-Flan-ICONS-37K as compact 20% subsets maintaining strong performance, hoping to inspire further exploration into data-efficient vision-language models.
+Limitations.
+Our approach’s main limitation is computational expense: computing gradients for large datasets is costly (Appendix §A), potentially constraining applicability to extremely large-scale data.
+To support the research community under resource constraints, we release the LLaVA-ICONS-133K dataset for iteration and model development.
+Broader impact.
+While our work does not directly imply negative impacts, it may indirectly propagate biases present in the original datasets.
+Positively, it enables more efficient and sustainable model development by reducing data redundancy and computational cost while maintaining performance.
+Continued discussion on these aspects remains critical for data-efficient vision-language models.
+Acknowledgments.
+This material is based upon work supported by the National Science Foundation under Grant No.
+2107048 and No.2112562.
+Any opinions, findings, and conclusions, or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the National Science Foundation.
+This work is also supported by the Singapore National Research Foundation and the National AI Group in the Singapore Ministry of Digital Development and Information under the AI Visiting Professorship Programme (award number AIVP-2024-001).
+We thank many people for their helpful discussion and feedback, listed in alphabetical order by last name: Allison Chen, Hee Seung Hwang, Hamish Ivison, Carlos E.
+Jimenez, Polina Kirichenko, Jaewoo Lee, Zhiqiu Lin, Tiffany Ling, Shengbang Tong, Ethan Tseng, Esin Tureci, Justin Wang, Zirui Wang.
+[…]
 Appendix A Computational complexity
@@
 This makes our method extendable for new tasks, as the expensive training data gradient computations only need to be performed once.
+Table 5: Detailed aggregation approaches comparison for Tab.
+3.
+Our proposed aggregation approach (Vote) consistently achieves the best overall performance (98.6% Rel.), outperforming both score-based (Merge, Max), their normalized variants (Merge-GausNorm, Merge-SumNorm) and rank-based (Round Robin, MinRank) baselines when selecting 20% of the LLAVA-665K dataset.
+Aggregation VQAv2 GQA VizWiz SQA-I TextVQA POPE MME MMB (en) MMB (cn) LLaVA-W Rel. (%)
+Full 79.1 63.0 47.8 68.4 58.2 86.4 1476.9 66.1 58.9 67.9 100
+Merge 75.7 59.6 47.9 65.5 55.5 86.0 1422.1 59.0 51.0 66.2 96.4
+Max 75.2 59.8 48.1 66.2 55.5 85.5 1470.7 58.3 51.8 66.2 96.1
+Merge-GausNorm 75.1 60.1 46.4 69.8 54.5 85.6 1482.6 58.9 52.5 66.3 96.8
+Merge-SumNorm 75.5 59.1 51.7 68.7 43.5 87.1 1478.3 59.5 50.9 69.8 95.3
+Round Robin 75.4 59.1 48.3 70.6 55.2 86.6 1474.5 61.6 51.5 66.9 96.7
+MinRank 75.2 59.0 49.7 70.4 55.1 86.9 1456.3 61.1 52.4 68.4 97.1
+Vote (ours) 76.3 60.7 50.1 70.8 55.6 87.5 1485.7 63.1 55.8 66.1 98.6
 A.3 Discussion on cost-benefit justification
@@
 By leveraging the curated, higher-quality dataset, they can substantially reduce training costs.
-Appendix B Related work
-B.1 Visual instruction tuning
-Multimodal large language models (MLLMs), e.g., Flamingo flamingo , LLaVA llava , BLIP2 blip-2 , and Cambrian (tong2024cambrian, ), enhance the capabilities of large language models (LLMs) on various multimodal tasks.
-A key component in advancing MLLMs is visual instruction tuning (llava, ), a training process that enables these models to interpret and follow instructions within a vision-language context, transforming them into versatile multimodal assistants.
-This tuning process not only improves the models’ instruction-following abilities but also aligns their outputs more closely with user expectations, thus enhancing their utility in practical applications (llava, ).
-B.2 Data selection
-Data selection methods survey can be categorized based on the types of information they utilize for selection.
-Representation-based approaches (semdedup, ; coincide, ) leverage neural embeddings to capture data representations.
-Loss trajectory-based methods (mindermann2022prioritized, ) prioritize data points that contribute most significantly to reducing generalization error over training.
-Gradient-based techniques (el2n, ; xia2024less, ; deng2024influential, ) select data based on gradient information.
-Recent work has explored various approaches to select optimal visual instruction tuning datasets.
-Concurrent work TIVE (liu2024less, ) employs gradient-based selection to identify representative instances.
-TIVE assumes that the number of specialist data should be proportional to task difficulty and thus samples specialist data based on an estimation of task difficulty.
-Our method does not rely on this assumption – we directly select samples that benefit the greatest number of tasks.
-COINCIDE (coincide, ) clusters data based on representations associated with concept-skill compositions.
-Our work follows targeted instruction tuning selection approach LESS xia2024less to utilize gradient information to calculate the specialist influence (i.e., the influence on a specific task) and extends it to general scenarios by aggregating information from various tasks and selecting data for multiple downstream tasks via majority voting.
-Appendix C Additional experiment details & ablations
-Task MME POPE SQA-I MMBench VQAv2 GQA VizWiz TextVQA LLaVA-W
-en cn
-$|\mathcal{D}_{\text{val}}|$ 986 500 424 1,164 1,164 1,000 398 800 84 84
-$|\mathcal{D}_{\text{test}}|$ 2,374 8,910 4,241 1,784 1,784 36,807 12,578 8,000 5,000 84
-Task Type Y/N Y/N MCQ MCQ MCQ VQA VQA VQA VQA VQA
 Table 6: Statistics of Target Tasks.
@@
 Task types include Multiple-Choice Questions (MCQ), Visual Question Answering (VQA), and Yes/No Questions (Y/N).
-C.1 Additional task details
-Here, we provide further details on the target tasks, as summarized in Tab.
-5.
-These tasks cover a wide range of multimodal benchmarks commonly used, including Yes/No questions (Y/N), multiple-choice understanding questions (MCQ) and visual question answering (VQA).
-C.2 Projection dimension
-We primarily set the projection dimension to 5120, reducing features from 338.7M to 5120 dimensions.
-The choice of 5120 was empirically validated for its trade-off between effective capturing gradient representation and maintaining a manageable parameter space.
-Our LLaVA-v1.5-7b-lora architecture includes a total of 7.4B parameters, with 338.7M parameters being trainable after LoRA adaptation, accounting for approximately 4.58% of the total parameter count.
-We further ablate different projection dimensions (1024, 2560, 5120, and 10240), with results provided in Fig.
-6(a).
-C.3 Warm-up Ratio
-To initiate training, we use 5% of the total training data.
-We conducted ablation studies to evaluate the impact of varying warm-up ratios (5%, 10%, 20%, and 100%) on selection performance, as shown in Fig.
-6(b).
-Our experiments reveal that increasing the warm-up data size does not lead to performance improvements.
-Surprisingly, models trained with smaller warm-up ratios (5-20%) consistently outperform those trained with the full dataset (100%).
-Specifically, the 5% warm-up ratio achieves the best performance at 98.6%, while using the complete dataset results in a performance drop to 97.8%.
-This finding suggests that a small subset of training data is sufficient and even beneficial for model initialization, and potentially gives better signals in the early training stages.
+Task MME POPE SQA-I MMBench
+en cn
+$|\mathcal{D}_{\text{val}}|$ 986 500 424 1,164 1,164
+$|\mathcal{D}_{\text{test}}|$ 2,374 8,910 4,241 1,784 1,784
+[…]
+Task VQAv2 GQA VizWiz TextVQA LLaVA-W
+$|\mathcal{D}_{\text{val}}|$ 1,000 398 800 84 84
+$|\mathcal{D}_{\text{test}}|$ 36,807 12,578 8,000 5,000 84
+[…]
+Appendix B Additional experiment details & ablations
+We provide comprehensive experimental details and ablation studies to support our main findings.
+We first present detailed task descriptions and dataset statistics (Tab.
+6), followed by detailed per-task results for aggregation method comparisons (Tab.
+5) and full results for unseen benchmarks (Tab.
+7).
+We then ablate key hyperparameters including projection dimension (§B.2) and warm-up ratio (§B.3).
+Table 7: Detailed results of unseen task generalization.
+Detailed results for Fig.
+3.
+Performance comparison on unseen benchmarks when trained on selected 20% subsets.
+Notably, we observe improvements on InfoVQA (103.8%), RealWorldQA (104.4%), and CMMMU (114.0%), highlighting strong generalization to unseen tasks.
+AI2D ChartQA DocVQA InfoVQA MMVet Naturalbench RealworldQA CMMMU Rel. (%)
+Full 55.4 17.5 28.9 26.5 31.1 12.4 52.4 22.1 100.0
+Random 50.2 15.1 25.2 24.3 29.6 11.1 49.8 21.9 91.8
+CLIP-Score 52.0 16.4 27.1 24.9 29.2 11.6 49.3 20.8 93.9
+EL2N 52.1 17.9 26.9 26.4 29.6 12.5 51.9 21.7 97.8
+Perplexity 53.0 16.4 27.1 27.1 29.4 11.4 50.3 23.9 97.0
+SemDeDup 53.2 16.4 26.4 26.6 29.1 13.2 51.1 22.8 97.8
+D2-Pruning 52.1 17.3 26.5 24.7 33.1 11.3 52.0 21.9 96.7
+Self-Sup 52.7 16.6 27.4 25.2 29.6 11.8 50.1 21.0 95.0
+Self-Filter 52.1 16.0 27.1 25.0 29.3 11.5 49.5 20.6 94.0
+COINCIDE 53.7 17.0 27.9 26.0 30.2 12.1 50.9 21.4 97.2
+RDS 53.0 16.8 27.5 25.4 29.8 12.0 50.3 21.3 95.6
+LLaVA-ICONS-133K 53.9 17.1 27.9 27.5 29.7 12.8 55.0 25.2 101.6
+[…]
 (a) Projection Dimension Ablation.
@@
 Figure 5: Ablation studies on (left) projection dimension and (right) warm-up ratio.
-Appendix D Additional analysis
-In this section, we provide additional analyses to better understand different aspects of our ICONS.
-We begin by analyzing the effectiveness of task-specific selections and their aggregation into a generalist subset (§D.1).
-We further explore whether incorporating visual dependency information into the selection process affects performance across different types of vision-language tasks (§D.2).
-We also evaluate the transferability of our selected subset across different model scales (§D.3).
-Additionally, we evaluate the consistency of our method across multiple runs (§D.4), showing its robustness and reliability.
+B.1 Additional task details
+Here, we provide further details on the target tasks, as summarized in Tab.
+6.
+These tasks cover a wide range of multimodal benchmarks commonly used, including Yes/No questions (Y/N), multiple-choice understanding questions (MCQ) and visual question answering (VQA).
+B.2 Projection dimension
+We primarily set the projection dimension to 5120, reducing features from 338.7M to 5120 dimensions.
+The choice of 5120 was empirically validated for its trade-off between effectively capturing gradient representations and maintaining a manageable parameter space.
+Our LLaVA-v1.5-7b-lora architecture includes a total of 7.4B parameters, with 338.7M parameters being trainable after LoRA adaptation, accounting for approximately 4.58% of the total parameter count.
+We further ablate different projection dimensions (1024, 2560, 5120, and 10240), with results provided in Fig.
+5(a).
+B.3 Warm-up Ratio
+To initiate training, we use 5% of the total training data.
+We conducted ablation studies to evaluate the impact of varying warm-up ratios (5%, 10%, 20%, and 100%) on selection performance, as shown in Fig.
+5(b).
+Our experiments reveal that increasing the warm-up data size does not lead to performance improvements.
+Surprisingly, models trained with smaller warm-up ratios (5-20%) consistently outperform those trained with the full dataset (100%).
+Specifically, the 5% warm-up ratio achieves the best performance at 98.6%, while using the complete dataset results in a performance drop to 97.8%.
+This finding suggests that a small subset of training data is sufficient and even beneficial for model initialization, and potentially gives better signals in the early training stages.
+Table 8: Single-task Selection (Specialist) vs.
+Consensus-aware Multi-task Selection (Generalist).
+The single-task data selection approach selects 20% of LLaVA-665K per task, while our consensus-aware multi-task data selection approach selects a total of 20% data across all tasks.
 Method VQAv2 GQA VizWiz SQA-I TextVQA POPE MME MMBench (en) MMBench (cn) LLaVA-W Bench
@@
 Generalist 76.3 60.7 50.1 70.8 55.6 87.5 1485.7 63.1 55.8 66.1
-Delta (%) 1.04 0.65 5.65 -1.43 0.18 -1.04 1.35 4.34 1.06 1.49
-Table 6: Single-task Selection (Specialist) vs.
-Consensus-aware Multi-task Selection (Generalist).
-The single-task data selection approach selects 20% of LLaVA-665K per task, while our consensus-aware multi-task data selection approach selects a total of 20% data across all tasks.
-D.1 From specialist to generalist
+[…]
+Appendix C Additional analysis
+In this section, we provide additional analyses to better understand different aspects of our ICONS.
+We include detailed vote distribution visualizations (Fig.
+6) extending the analysis from §4.6.
+We analyze the effectiveness of task-specific selections and their aggregation into a generalist subset (§C.1).
+We further explore whether incorporating visual dependency information into the selection process affects performance across different types of vision-language tasks (§C.2).
+We also evaluate the transferability of our selected subset across different model scales (§C.3).
+Additionally, we evaluate the consistency of our method across multiple runs (§C.4), showing its robustness and reliability.
+Figure 6: Vote distributions across selection ratios.
+Each line shows the vote distribution for a different selection ratio (5% to 95%).
+At lower ratios (e.g., 5%), most samples receive zero votes, resulting in a left-skewed distribution.
+As ratios increase (e.g., 50%, 70%, 90%), the distribution shifts rightward and flattens, with more samples receiving votes from multiple tasks.
+Stricter criteria (lower ratios) identify task-specific samples, while relaxed criteria (higher ratios) reveal broader multi-task agreement.
+C.1 From specialist to generalist
 To understand the intermediate task-specific influence matrices we obtained from the specialist stage, we select 20% of data for each individual task.
 The task-specific data (Specialists) achieves comparable or superior performance than full data training (Tab.
-6).
+8).
 With influence consensus at the generalist stage, we select a 20% subset with only a 1.33% average drop across tasks compared to specialist baselines.
 This validates our approach: by understanding task-specific influence patterns and building consensus across tasks, we can identify a compact, universal training set that maintains strong performance with significantly less data.
-D.2 Visual dependency influence ranking
-Recent work tong2024cambrian has shown that vision-language tasks vary in their reliance on visual information: tasks like MMBench mmbench depend heavily on visual grounding, while others like SQA-I sqa can be handled primarily through language, showing only a 5% drop in performance when visual input is removed tong2024cambrian .
+Figure 7: Pairwise overlap heatmap between specialists. The values show overlap percentages between benchmarks’ selected samples.
+Divergent multi-task influence patterns.
+As shown in Fig.
+7, the pairwise overlap heatmap shows notable variation in training data influence across tasks.
+High overlap – e.g., VQAv2 and VizWiz (49.0%) or POPE and GQA (60.2%), suggests that certain samples are beneficial across similar tasks.
+However, low overlap, like the 3.3% between MMBench (en) and GQA, highlights that highly influential samples for one task may have limited impact on others.
+Even closely related tasks, such as MMBench in different languages (English and Chinese), share 67.4% of influential samples.
+To understand task-specific influence matrices from the specialist stage, we select the top 20% samples per task (Specialists).
+Overlap with our generalist subset (Fig.
+8) varies significantly, from minimal in tasks like VQAv2 (3.27%) and VizWiz (3.28%) to substantial agreement in tasks like LLAVA-W Bench [llava] (24.21%).
+These findings empirically demonstrate significant overlap in influential samples across tasks and validate our approach: by analyzing task-specific gradient-based influence patterns and building consensus across tasks, we can identify a compact subset that captures broadly useful samples across tasks, yielding strong performance with significantly less data.
+Figure 8: Data overlap between specialists and generalist selection.
+Overlap varies significantly, from 3.27% (VQAv2) to 24.21% (LLAVA-W Bench), reflecting varying alignment between task-specific and consensus selections.
+VQAv2 GQA VizWiz SQA-I TextVQA POPE MME MMBench (en) MMBench (cn) LLaVA-W Bench
+Full 79.1 63.0 47.8 68.4 58.2 86.4 1476.9 66.1 58.9 67.9
+w/o VDS 76.3 60.7 50.1 70.8 55.6 87.5 1485.7 63.1 55.8 66.1
+w/ VDS 75.8 60.9 50.3 69.5 54.8 86.8 1489.3 64.3 56.3 67.9
+Delta (%) -0.66 +0.33 +0.40 -1.84 -1.44 -0.80 +0.24 +1.90 +0.90 +2.72
+Table 9: Impact of Visual Dependency Score (VDS) on Selection Performance.
+Rows show performance without VDS, with VDS, and the performance change (Delta).
+VDS improves performance on LLaVA-W Bench (+2.72%), MMBench (en) (+1.90%), and MMBench (cn) (+0.90%), but decreases performance on SQA-I (-1.84%), TextVQA (-1.44%), and POPE (-0.80%).
+C.2 Visual dependency influence ranking
+Recent work [tong2024cambrian] has shown that vision-language tasks vary in their reliance on visual information: tasks like MMBench [mmbench] depend heavily on visual grounding, while others like SQA-I [sqa] can be handled primarily through language, showing only a 5% drop in performance when visual input is removed [tong2024cambrian].
 To take visual dependency of training data into account, we further explored gradient-based Visual Dependency Score (VDS).
@@
 This matrix can be used to further rank and select training data that are most impactful for tasks requiring strong visual grounding, ensuring that the selected subset effectively supports vision-dependent performance.
-Our empirical results demonstrate that VDS-based data selection has varying effectiveness across different vision-language tasks (Tab.
-7).
-The approach shows substantial improvements on tasks requiring strong visual understanding, such as open-ended generation (LLaVA-W Bench: +2.72%) and multiple-choice understanding (MMBench-EN: +1.90%, MMBench-CN: +0.90%).
+Our empirical results demonstrate that VDS-based data selection has varying effectiveness across different tasks (Tab.
+9).
+It shows substantial improvements on tasks requiring strong visual understanding, such as open-ended generation (LLaVA-W Bench: +2.72%) and multiple-choice understanding (MMBench-EN: +1.90%, MMBench-CN: +0.90%).
 However, tasks that primarily rely on textual reasoning show decreased performance, including SQA-I (-1.84%) and TextVQA (-1.44%).
-These results align with and extend the findings in Cambrian tong2024cambrian , demonstrating that the effectiveness of VDS corresponds to a task’s visual dependency - tasks that maintain performance without visual inputs show limited or negative impact from VDS-based selection, while visually-dependent tasks benefit significantly.
+These results align with and extend the findings in Cambrian [tong2024cambrian], demonstrating that the effectiveness of VDS corresponds to a task’s visual dependency - tasks that maintain performance without visual inputs show limited or negative impact from VDS-based selection, while visually-dependent tasks benefit significantly.
 This pattern suggests that VDS effectively identifies training samples where visual information plays an important role in training.
-VQAv2 GQA VizWiz SQA-I TextVQA POPE MME MMBench (en) MMBench (cn) LLaVA-W Bench
-Full 79.1 63.0 47.8 68.4 58.2 86.4 1476.9 66.1 58.9 67.9
-w/o VDS 76.3 60.7 50.1 70.8 55.6 87.5 1485.7 63.1 55.8 66.1
-w/ VDS 75.8 60.9 50.3 69.5 54.8 86.8 1489.3 64.3 56.3 67.9
-Delta (%) -0.66 +0.33 +0.40 -1.84 -1.44 -0.80 +0.24 +1.90 +0.90 +2.72
-Table 7: Impact of Visual Dependency Score (VDS) on Selection Performance.
-Rows show performance without VDS, with VDS, and the performance change (Delta).
-VDS improves performance on LLaVA-W Bench (+2.72%), MMBench (en) (+1.90%), and MMBench (cn) (+0.90%), but decreases performance on SQA-I (-1.84%), TextVQA (-1.44%), and POPE (-0.80%).
+Table 10: Cross-Architecture Generalization.
+Our LLaVA-ICONS-133K selected via LLaVA-v1.5-7B model (7B-selected) shows strong cross-architecture transferability, achieving 97.3% Rel., while the data selected via LLaVA-v1.5-13B model (13B-selected) reaches 98.1%, showing that our selected subset generalizes well to different architectures.
 VQAv2 GQA Vizwiz SQA-I TextVQA POPE MME MMBench (en) MMBench (cn) LLAVA-W Rel. (%)
@@
 13B-selected 78.9 61.2 57.5 71.3 58.4 85.9 1535.2 66.1 59.8 68.8 98.1
-Table 8: Cross-Architecture Generalization.
-Our LLaVA-ICONS-133K selected via LLaVA-v1.5-7B model (7B-selected) shows strong cross-architecture transferability, achieving 97.3% Rel., while the data selected via LLaVA-v1.5-13B model (13B-selected) reaches 98.1%, showing that our selected subset generalizes well to different architectures.
-D.3 Cross-architecture generalization
-We further conduct experiments on cross architecture generalization to evaluate the transferability of our selected data across different model scales.
+C.3 Cross-architecture generalization
+We conduct experiments on cross architecture generalization to evaluate the transferability of our selected data across different model scales.
 While our subset was initially selected using LLaVA-v1.5-7B as the base model, we investigate whether these same examples remain effective for training larger architectures like LLaVA-v1.5-13B.
@@
 Both 7B-selected and 13B-selected subsets outperform random selection (95.7%), with the 13B-selected option showing particular strength in reasoning tasks like MMBench and POPE.
-This suggests that our selected subset captures fundamental visual-language understanding patterns that generalize well across different model architectures.
+This suggests our selected subset captures fundamental visual-language understanding patterns that generalize well across different model architectures.
 Figure 9: Rel. (%) Across Runs. We show the performance across three different runs for random selection and our ICONS.
-D.4 Consistency analysis
-To evaluate the consistency of ICONS, we conduct three independent runs of the experiments.
-As shown in Fig.
-7, our method demonstrates high consistency across different runs, achieving 98.6$\pm$1.2Rel., which shows a notable improvement over random baseline, which achieves 95.8$\pm$2.7%.
-The lower standard deviation (1.2% vs 2.7%) further indicates that our approach produces more stable and reliable outcomes compared to the random baseline.
-Appendix E Algorithm details
+C.4 Consistency analysis
+[…]
+Appendix D Algorithm details
 We provide detailed pseudocode for our two-stage ICONS framework.
@@
 13:// Select top samples based on total votes
-14:$S\leftarrow\text{top-}p\text{ samples by }I_{\text{vote}}$
+[…]
 15:return Selected subset $S$
-Appendix F Limitations
-Our approach primarily faces one practical limitation: computing gradients for large training datasets is computationally expensive (Appendix §A).
-This computational overhead could potentially constrain the method’s applicability when working with extremely large-scale datasets.
-To support broader research community, we release LLaVA-ICONS-133K dataset to help research iteration and model development under resource-constrained settings.
-Appendix G Broader impact
-Our exploration focuses on scientific understanding and practical applications of vision-language data selection.
-While our work does not directly imply negative impacts, it may indirectly propagate existing biases present in the original datasets.
-Therefore, it is important to incorporate rigorous bias-mitigation measurements for data selection.
-On the positive side, our method enables more efficient and sustainable model development by reducing data redundancy, computational cost while maintaining or even improving performance.
-Discussion on these critical aspects should remain a priority as we further explore the potential of vision-language data selection.
-Appendix H Future work
+Appendix E Future work
 Our work opens several promising research directions for improving vision-language data selection.
@@
 Additionally, investigating more efficient gradient computation and storage methods would help scale these methods to larger datasets while maintaining strong performance across diverse vision-language tasks.
-Appendix I Visualizations
-I.1 Representation-based vs. Gradient-based data selection
+Appendix F Visualizations
+In this section, we provide detailed visualizations.
+We compare representation-based versus gradient-based data selection approaches (§F.1), showing qualitative differences in selected samples, and visualize the most influential examples identified by both task-specific specialists and the generalist model across all target tasks (§F.2).
+Additionally, we present examples of low-vote training samples (Fig.
+10), which are referenced in §4.7 of the main paper, to illustrate the characteristics of data filtered out by our method.
+F.1 Representation-based vs. Gradient-based data selection
 While ICONS leverages gradient-based influence signals, we explore how representation-based data selection (RDS) performs in the same setting (§4.2).
-We analyze the top-ranked training examples selected by RDS after the generalist stage in Fig.
-8 vs.
-samples selected by ICONS in Fig.
-19.
+We compare the top-ranked training samples selected after the generalist stage by RDS with those selected by ICONS in Fig.
+22.
 Interestingly, we observe that the representation-based variants often favor training examples with repeated images or instructions, which may dominate the learned representations without contributing to better generalization.
 Some of the highest-scoring samples under representation-based similarity are duplicated image-question pairs with only the answer choices shuffled.
-We hypothesize that this is a side effect of the way multimodal representations are constructed—where visually dominant or textually redundant samples occupy high-density regions in embedding space.
+We hypothesize that this is a side effect of the way multimodal representations are constructed, where visually dominant or textually redundant samples occupy high-density regions in embedding space.
 However, these samples do not necessarily translate into broader utility across tasks, as seen in the performance gap in Tab.
@@
 While we leave these questions open for future exploration, our results suggest that gradient-based influence, though computationally more expensive, is better aligned with generalization and multi-task data mixture settings.
+Figure 10: Examples of 0 & 1-vote training samples.
+These low-vote samples exhibit characteristics that explain their low selection frequency: (1) subjective questions lacking objective ground truth (e.g., “Is the woman pretty?”), (2) trivial questions providing minimal visual learning signal (e.g., “Is the door open?”), (3) redundant questions about the same image content.
+By filtering out such potentially harmful or redundant examples, ICONS focuses training on more informative and generalizable samples.
 Figure 11: Top three samples selected by RDS.
 These highly-ranked samples are selected due to representation-based similarity but do not necessarily contribute to better generalization, highlighting a key limitation of representation-based selection in multimodal settings.
-I.2 Specialists & Generalist
-We visualize the most influential top three examples across specialists (figs.
-9, 10, 11, 12, 13, 14, 15, 16, 17 and 18) and the generalist selection (Fig.
-19), along with samples from their corresponding tasks.
+F.2 Specialists & Generalist
+We visualize the most influential top three examples across specialists (Figs.˜12, 13, 14, 15, 16, 17, 18, 19, 20 and 21) and the generalist selection (Fig.
+22), along with samples from their corresponding tasks.
 Notably, the selected high-influence examples by specialists show strong task-specific characteristics both structurally and contextually - they mirror the key attributes of their target tasks in terms of question structure, reasoning patterns, and required visual-language understanding capabilities.
@@
 This suggests that the most effective training samples are those that combine multiple types of reasoning while maintaining clear, unambiguous ground truth that can be consistently learned across tasks.
-Figure 9: VQAv2. Top-left: A sample from VQAv2 vqav2 . Remaining panels show top three influential samples selected using the specialist influence ranking step.
-Figure 10: GQA. Top-left: A sample from GQA gqa . Remaining panels show top three influential samples selected using the specialist influence ranking step.
-Figure 11: VizWiz. Top-left: A sample from VizWiz vizwiz . Remaining panels show top three influential samples selected using the specialist influence ranking step.
-Figure 12: SQA. Top-left: A sample from SQA sqa . Remaining panels show top three influential samples selected using the specialist influence ranking step.
-Figure 13: TextVQA. Top-left: A sample from TextVQA textvqa . Remaining panels show top three influential samples selected using the specialist influence ranking step.
-Figure 14: Pope. Top-left: A sample from Pope pope . Remaining panels show top three influential samples selected using the specialist influence ranking step.
-Figure 15: MME. Top-left: A sample from MME mme . Remaining panels show top three influential samples selected using the specialist influence ranking step.
-Figure 16: MMBench (en). Top-left: A sample from MMBench (en) mmbench . Remaining panels show top three influential samples selected using the specialist influence ranking step.
-Figure 17: MMBench (cn). Top-left: A sample from MMBench (cn) mmbench . Remaining panels show top three influential samples selected using the specialist influence ranking step.
-Figure 18: LLaVA-W Bench. Top-left: A sample from LLaVA-W Bench llava . Remaining panels show top three influential samples selected using the specialist influence ranking step.
+Figure 12: VQAv2. Top-left: A sample from VQAv2 [vqav2]. Remaining panels show top three influential samples selected using the specialist influence ranking step.
+Figure 13: GQA. Top-left: A sample from GQA [gqa]. Remaining panels show top three influential samples selected using the specialist influence ranking step.
+Figure 14: VizWiz. Top-left: A sample from VizWiz [vizwiz]. Remaining panels show top three influential samples selected using the specialist influence ranking step.
+Figure 15: SQA. Top-left: A sample from SQA [sqa]. Remaining panels show top three influential samples selected using the specialist influence ranking step.
+Figure 16: TextVQA. Top-left: A sample from TextVQA [textvqa]. Remaining panels show top three influential samples selected using the specialist influence ranking step.
+Figure 17: Pope. Top-left: A sample from Pope [pope]. Remaining panels show top three influential samples selected using the specialist influence ranking step.
+Figure 18: MME. Top-left: A sample from MME [mme]. Remaining panels show top three influential samples selected using the specialist influence ranking step.
+Figure 19: MMBench (en). Top-left: A sample from MMBench (en) [mmbench]. Remaining panels show top three influential samples selected using the specialist influence ranking step.
+Figure 20: MMBench (cn). Top-left: A sample from MMBench (cn) [mmbench]. Remaining panels show top three influential samples selected using the specialist influence ranking step.
+Figure 21: LLaVA-W Bench. Top-left: A sample from LLaVA-W Bench [llava]. Remaining panels show top three influential samples selected using the specialist influence ranking step.
 Figure 22: Generalist. We show top three influential samples selected after the generalist stage.